A Complete Bibliography of Publications in Technometrics for the decade 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

04 March 2019 Version 1.03

Title word cross-reference

0.5 [Lad77]. 18 [Joh73a]. 2 [BN72]. 2^{m} [SC71]. 2^{n} [Joh79]. 2^{n−k} [Joh79]. 2 × 2 [MDM77]. 3 [Lin79]. 3^{n} [Hok74]. ^{4} [Dic74]. ^{5} [Dic74]. ^{n} [TW70]. A [Bad70, Bor72]. α [AS73, McC78a]. A ⊂ B [Bad70, Bor72]. B [Bad70, Bor72]. β [AS73, GPU76]. C_{P} [Mal73, Ano73i, Ken71]. D [GK77a, Luc78, Mit74a, Mit74b, MB78a, MB78b, NGM70, SD75]. F [DW72, DW73, HH77, J.79u, Joh73b, Obe77, Pap72, PW72, HB71]. k [HP73, Ode71, WK71, Zei73]. log Y [Scl72]. m [FM77, HP73]. m ≤ 6 [SC71]. N [Shu71, Gre70b, Sap73]. n = 92 [Wat70a]. O [MHL74]. p [Dun78a, FM77, For72, Lad77, MC72]. P (Y < X) [KKS76]. P [Y < X] [WK77]. Pr(Y < X) [Dow73]. Pr{Y < X} [JTS75, Ton74, Ton75]. R^{2} [DH74]. s [Bha73, Lik74]. s^{n} [AT79]. σ^{2} [JMM78]. t [AS73, KP79, Ros75c, SI70, TC72, YM74]. T^{2} [BJN71]. T_{M}^{2} [JH79]. V [SC71]. W [GH79, HH73, HL76, Ste78, Tho75]. X [Scl72]. Y [Scl72]. |X′X| [Ano71a, Ano72e, Ano73h, BD71, Dyk71, Eva79].

-acceptability [McC78a]. -Blue [MHL74]. -Charts [Dun78a]. -content [GPU76]. -distribution [TC72]. -Distributions [HB71]. -Optimal [Luc78, MB78b, Mit74a, Mit74b, MB78a]. -Optimality [NGM70, SD75]. -optimum [GK77a]. -Sample [Ode71]. -sided [BN72]. -Solution [Ros75c]. -spare [Bha73]. -Spline [Lin79]. -Statistic [DW72, HH77, PW72]. -Statistics [Pap72]. -Test [AS73, SI70]. -Tests [DW73, Obe77]. -Transform [HL76].

1 [Bla77, Lac71b]. 105D [Koy78, Koy79]. 129th [Ano70a]. 14 [GH75]. 16 [GH75, JTS75, Ton75]. 17 [Smi76]. 18 [HK77]. 1972 [J.75c].

2 [Che71, J.79f, Sne74]. 21 [Kit81]. 25 [K.73a].

3 [BCK78, J.76h].

4 [J.77k].

5 [J.78o].

77 [Hea78]. ’78 [J.79n]. 781 [BS79].

80f [Kit81].

A. [GW71, HH73]. Abadie [Sha71]. Abridged [Ano71f]. Absolute [JH79, NW79]. Absorbing [Bri74]. abstract [Tuk70]. Abstracted [J.75c]. Abstracts [Ano70a]. Accelerated [BS71, MH78, Nel72a, NK76, NM78, Sha78, Sin71, Sin73]. acceptability [McC78a]. Acceptable [OB76]. Acceptance [Bin74b, BLB73, MCM75]. Accountability [Lev79]. Accounting [PB77]. Accuracy [Gru73b]. Adaptive [HR75, MRR78, RRH73]. Addendum [Ano71a, Ano77a, CL71, D’A71a]. Addition [Bor72]. Additive [Man71a]. Adequacy [HJF78, McK79]. Adequate [SD77, Ell78]. Adjacent [AJ77]. Adjusted [MHL77b]. Adjustment [CH74]. Advances [Law75b, J.77a]. Advisory [K.74e]. Affairs [Bon77]. after [Kow70, Lem74]. against [EB75, Ode71]. Age [CH74, IS76, RR76, Sch71b]. Agreement [Leo70]. Agterberg [Jon76b]. AIC [Kit81, Kit79]. Aid [CO75, Sne75a]. Aided [Cha70a]. Aims [Bon76b]. Air [Ano73j, CG76, Mar73, MS73, Sin72]. Aircraft [Bur73, Pas79]. Alan [Fie73, Hei76]. Albert [Web73]. Alder [Woo74b]. Alfred [Rog70c]. Algorithm [BD72a, GH79, HG73, HL76, Jon76a, LF79, Mit74a, RJ78]. Algorithmic [J.79b]. Algorithms [Ens76, JS76, YC71, Gor76, J.79u]. Allan [Har75, Jac80]. Allin [Kup71]. Allocation [CHH^{+}71, Gro72, Har74, BN72]. Allocations [Spr79]. Alternate [GM79, ZNL79]. Alternating [HD70]. Alternative [BA70, HM77, Igl73, Man71d, Sac72]. Alternatives [EB75, KARR75, Ode71]. Amendment [Goo76]. American [Ano70a]. Among [J.76d, JH73a, Par78b, SU79, Ury76]. Among-Groups [J.76d]. Amstadter [Eas72]. Amstutz [Nay72]. Analyses [Odo77, Hub78]. Analysing [Gen77]. Analysis [Add72, Aro76, Bai74b, Bak74, BH70, BD78, Bis79, BHME73, BM74, Bur73, Bus71, Car77b, DM77, Dra79, Dun71, Ens76, Fed74b, Fed77, Fre73c, Ger75, Goo71, Gri70, HH77, Hah79, Har75, HM76b, Hoo79, J.75c, J.75f, J.77m, J.77h, J.78f, JM79, JK71, Joi77, KO77, KM79, Kin78, Kle77, Krz79, Lac74b, LM72, Man71a, Mar75, Mar76, McC75a, McD75, McK76, MBC79, McL72, MW75b, Mul70, Nai74, Nel73, Obe75, Odo78, PW77, Pre77, Ran70c, Sac76, Sau73, SH79, See73, SW72, Sim79, SOE75, Sne71, Sne72b, Sne73, Sto76, WGM74, Weg76, WB75, WG79, Win71, Wol74, vNS76, vN79, Cro78, Dav77, Dun72, Dun77a, Fre74b, Gle78, Gou73, Hil74, J.76b, J.77j, J.78d, J.79e, Ket75, Kha74, Lan75, Nel74, O’F73, Ram73, Ste75, Gri73, K.73a, Sne76]. Analysts [Sto76]. Analytic [FH79, LSYF79, SCSB76]. Analyzing [HJ76, HM77, Whi70, J.79c]. Anant [Ing78b, Kha74]. Anderson [Lan75, Ano77f]. Anderssen [Lac71a]. Andrews [Law75b]. Annotated [Ano73a, J.77b]. Annual [Ano70a]. Anthropology [Ber73]. Anwendungen [Ens77]. Any [Cha79, GZ79]. AoV [HJ76]. APL [Mar76]. Application [BB79, Ber72b, Bur73, CG76, Eva71, Eva72, FM71, FH79, Gre70a, Gro71a, Hah77, HS71, Law71, LSYF79, Low70, MRR78, ML75, NH72, Pas79, Ros75c, Spr78, SU79, TW72a, Zel72, ZNL79, TC72, Sid75]. Applications [Aro76, CB75, Cle72b, DV78, Dun77b, Hei76, HB71, HK70a, HR75, J.77n, J.78i, J.79v, Nel72b, Par72, Sin72, Ten72, Urq73, Zys70, Ban73, Bri73a, Car77b, Fre73c, Ham75, J.76i, J.77g, Joh75, Joh71b, Jon76b, Lev79, Man71c, Ran70a, Ran70c, Rog70d, Sam73, Sie75, Sim76, Spr72, Zac74, J.78a, Lyn78]. Applied [Ano80c, Cro78, J.76a, J.78c, McD79, Mye74, J.75b, J.79e, J.79d]. Approach [AB74, BA70, Gut73a, HG78, Leo75, Man71d, MS79, Spr78, GDF78, LH77, McC72, Mer77]. Appropriate [Mie74]. Approximate [BK70b, DW72, DW73, DV78, Gru71b, Gru71c, HM76b, Kam73, Kus79, Lan72, Lun75, MC72, Pre75, Ros75c, WB75]. Approximately [MG74]. Approximating [Bin74a, DH74, Joh73b, Cra74]. Approximation [BG74, GZ79, Gue72, Jer70, GH75, KP79]. Approximations [Bod70, Dun78c, IM70, MCL76, Rey75, Hel71a]. Aram [Rog70d]. Arbitrary [Man75, Ury76]. Arguments [Cha70b]. Arithmetic [Ste72b]. Armitage [Fle76, Lac72b]. Arnold [Goo75, Nay72, Goo76, Sac75]. Arrhenius [Sin73]. Arrow [K.73c]. Artificial [BL74a, Hel71b]. Ash [Moe72]. ASN [HM76a, Kow71, SP72]. Aspects [Boy74, Dol70, Gua71, Jac80]. Assay [Sal75b]. Assessment [Gro71a, J.78j]. Assessments [FB75]. Associated [BHME73, JM79, LC70]. Association [Ano70a]. Assumptions [BSH71, Sne71]. Assurance [Ziz78]. Asymmetrical [CDR79]. Asymptotic [Joh70a, KB75, McL72, McL74a, McL74b, Kra76]. Asymptotically [Her70, Law77b, MH77]. Atchley [J.76d, J.76e]. Atom [JK74]. Attribute [Ano72g, Chi75, Coh70, Gue71a, Gue71b, Hal77a, Lad77]. Attributes [BL74a, BLB73, Chi74, HM76a, SP77]. Auditor [Ale72, Ano74g, Com71, Com70]. Audrey [Qua71]. Augmentation [All74, Ano71a, Dyk71, Eva79]. Augmented [Haw75]. Augmenting [See73]. August [Bor72]. Auswertungsmethoden [Bed74]. Author [K.71a]. Authors [J.77e, J.77p]. Autocorrelations [Cle72b, Par72]. Autoregression [McC75b, Weg74]. Autoregressive [SWD70, SWD74]. Autoregressive-Moving [SWD74]. Availabilities [CH73a]. Availability [GC79, ST72]. Available [Weg72]. Avenhaus [Lev79]. Average [DS79, Eld74, Gue71a, Hel74, Rey75, RH78, Ste70, SWD74, Wic72]. Averages [Hin77a, Ste72b]. Averaging [Yue78]. Axiomatic [Sim76]. Aykac [J.78i].

B [Boa79a, Dil71, Dun77b, Ehr77, Hah79, Har72, Hea78, J.75g, J.76h, J.77j, J.77k, J.78l, J.78k, J.78g, J.78m, J.78j, J.78o, J.79n, K.73a, Moe72, Odo78, Sen75, Spr79, Woo74b]. Back [Ano70b, Ano70c, Ano70d, Ano70e, Ano71b, Ano71c, Ano71d, Ano71e, Ano72a, Ano72b, Ano72c, Ano72d, Ano73b, Ano73c, Ano73d, Ano73e, Ano74a, Ano74b, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano76a, Ano76b, Ano76c, Ano76d, Ano77b, Ano77c, Ano77d, Ano77e, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d]. Background [Jon76b, YB76]. Badia [Qua71]. Baines [J.78p]. Bajpai [Boa79a]. Balanced [Joh78, SC71, Joh73a]. Balkema [Sen74]. Bancroft [K.73d]. Band [Mal74, Mar74, BT74]. Band-Spectroscopic [Mal74, Mar74, BT74]. Bands [Gaf78, Sch74a, SW75, TC73, TS79, TC72]. Baneriee [Cam77]. Banerjee [Ben76]. Bard [Cha75b]. Bardhan [Kee75]. Barlow [J.77j, Kal75, Smi77, Rea71, WG78]. Barnett [Bon76a]. Barra [J.78m]. Barrie [Tsi78]. Bartholomew [Kal75]. Bartlett [J.76f]. Based [Ano72g, AJ77, Bai72, BR79, CC71, Coh65, Coh72, Dye73, FKRO71, Gue71b, HM77, Hok74, Kam74, LaB77, Lem75, Man70a, MF73, MG74, McD76, Mer79, Nád73, OlHS77, PS70, SR75, ST73, SS74, Sor73, Tak72, LR78]. Basic [Joh75, Jon73, Moe72, Sim76]. Batches [Leo70]. Bayes [Gro72, LaM78, TC75a]. Bayesian [Goo75, Goo76, Hol73, J.75h, J.78i, BB79, Bar72, BSH71, DG71, Dun74, Gut73a, GDF78, HG78, Joh70a, LH77, Leo75, MW78, Pie73, SOE75, WG78, Zac70, Hil74]. Be [Hoc72]. Bedient [Fin75, Str75]. before [Kow70]. Behavior [BGH73, BF74, GW75, Lut73]. Behavioral [Leh71]. Behrens [LH74, Ros75c]. Behrens-Fisher [LH74]. Bell [J.78b]. Bendat [Har75]. Benjamin [Kup71]. Bennett [K.73b]. Bensoussan [Dun77b, J.79d]. Bergstrom [K.74b]. Berkeley [Fre73b]. Bernard [Gle77, Hel71b]. Bernoulli [Sap73]. Bert [K.71a]. Bertram [Eas72]. Best [DH74, GH75, Luc76, Man71b, MF73, Nád73, NH73, Ng76, SG71]. between [All74, AR73, BA70, DA73, DAH73, Hah77, Har73a, Haw75, McL74b, Ng76]. Bias [AR79, Nas78, Tho73]. Biased [HSL76, HK70b, Low74, Mar70, MW73, Sch72, BC71]. Bibby [Boa79b].

Bibliography [Ano73a, Cor79, J.75c, J.77b]. Bickel [Law75b]. Bill [J.79a]. Billingsley [Sne70]. Binary [WRLI75]. Binomial [BG74, Bin74a, CL70, CL71, DG71, Fau70, Gue72, Gue74, Hel71a, MC72, Nel75, Pat73, GH75]. Biochemical [K.73a]. Biological [vdV75]. Biomedical [Hea78, Hei76]. Biometrika [Sne74]. Biometry [J.75f]. Bionomial [Bin74a]. Biplot [BG78]. Birnbaum [J.78l]. Bishir [Leh71]. Bishop [Odo77, Ano79e]. Bivariate [BJN71, GKS75, Ham72a, Kow70, LO79, Par70, Smi76]. Blakeslee [Mil72a]. Block [Jon76a, TC75b]. Blocks [HS71]. BLUE [SG71, MHL74]. BLUEs [NS79]. Blum [Hog74]. BMDP [Hea78]. BMDP-77 [Hea78]. Bock [Wil79a]. Bofinger [J.76b]. Boltyanskii [Bis72b]. Book [Add72, Add79, And70, And75, Ano71f, Ano74e, Bak74, Ban73, Bed74, Bel70, Ben76, Ber72b, Ber73, Bha72, Bis72a, Bis72b, Boa79a, Boa79b, Bon76a, Bon76b, Bon77, Bow77, Bow78, Boy74, Bri73a, Buc73, Cam77, Car77a, Car77b, Car74, Cas76, Cha75a, Cha75b, Che71, Cle72a, Con74, Coo72, Cor71, Cra74, Cro78, Cru78a, Cru78b, Dal77, Dav70, Dav77, Dil71, Doo72, Dra79, Dud79, Dun72, Dun77a, Dun77b, Dye74, Dyk76, Eas72, Ehr77, Ens76, Ens77, Ens78, Eri72, Fed74a, Fed77, Fie73, Fin75, Fle76, For71, For73, Fre73a, Fre73b, Fre73c, Fre74b, Fre74a, Fre75, Gen77, Geo72, Ger75, Gib74, Gie74, Gil74, Gle78, Gle77, Gol74, Goo75, Gor76, Gou73, Gow72, Gra72, Gri73, Gro77, Gru71a, Gua71, Hac78, Hac79, Hah79, Hal77b, Ham75, Han73]. Book [Har77, Har70, Har75, Har72, Hay73, Hea78, Hei76, Hei78, Hei79, Hel71a, Hel71b, Hil74, Hog74, Hol73, How77, Hub78, Hul73, Hun70, Hun74, Ing75, Ing78a, Ing78b, J.75d, J.75c, J.75g, J.75b, J.75h, J75e, J.75a, J.75f, J.76c, J.76b, J.76g, J.76d, J.76e, J.76f, J.76a, J.76i, J.76h, J.77b, J.77d, J.77f, J.77e, J.77l, J.77c, J.77n, J.77j, J.77g, J.77m, J.77i, J.77k, J.77h, J.77a, J.78n, J.78h, J.78l, J.78k, J.78e, J.78f, J.78g, J.78d, J.78p, J.78q, J.78i, J.78m, J.78j, J.78c, J.78b, J.78a, J.78o, J.79q, J.79i, J.79p, J.79g, J.79e, J.79j, J.79c, J.79l, J.79r, J.79k, J.79o, J.79t, J.79b, J.79v, J.79d, J.79n, J.79a, J.79s, J.79u, J.79f, J.79h, J.79m, Jac80, Jer70, Joh75, Joh71b, JH73b, Joh73c, Jon73, Jon76b, K.71a, K.71d, K.71b, K.71c, K.71e, K.73d, K.73b, K.73c, K.73e, K.73a, K.74c, K.74e, K.74a, K.74b, K.74d, Kab74, Kal75, Kee75]. Book [Ket75, Kha74, Kin78, Kin79a, Kin79b, Kle77, Kno71, Kra76, Kup71, Lac71b, Lac71a, Lac72b, Lac72a, Lac74a, Lan73, Lan75, Law75b, Leh71, Lev79, Lin75, Luk73, Lut73, Lyn78, Man71c, Man73, Mar75, McC72, McC78b, Mer77, Mer79, Mer72, Mih75, Mil72a, Moe72, Mye74, Nay72, Nel74, O’F73, Odo77, Odo78, Pre77, Qua71, Qua74b, Qua74a, Qua75, Ram73, Ran70a, Ran70c, Ran70b, Rog70d, Rog70b, Rog70a, Rog70c, Ron72, Ros75a, Rus73, Sac75, Sal75a, Sam73, Sax72, Sch74b, Sch75, Sen70, Sen74, Sen75, Ser73, Sha71, Sid75, Sie75, Sim76, Sim79, Sim73, Smi70a, Smi77, Sne72a, Sne74, Sne76, Sne70, Spr72, Spr78, Spr79, Sta70, Ste75, Sto76, Str75, Sym71, Sys75, Sys76, Sys78, Tan70, Ton72, Tsi78, Urq73, Wat73, Web73, Weg76, Wei71, Wil79a]. Book [Wil70, Win71, Woo75, Woo74b, Woo74a, Zac74, Zie78, Ziz78, Zys70, vdV75, Boa79a, Mih77]. Books [Ano70k, Ano70l, Ano70m, Ano70n, Boa79a]. Bose [Ano78e]. Both [Cha72, RSO72]. Boullion [Gow72]. Bound [AF76, Her70, Sar71]. Boundary [And75]. Bounds [Fin76, FW74, Gru71b, Gru71c, JL79, KARR75, Law75a, Man70a, MF73, MG74, MF75a, MF77, Ury72]. Bowker [Web73]. Box [Add79, And70, Hil74, Kle77, GK77a, J.79p]. Boyce [Gor76]. Bradley [Bel70]. Bradu [Ano78e]. Branch [AF76]. Branch-and-Bound [AF76]. Brandt [Dav77]. Bremner [Kal75]. Breslow [J.79h]. Bressler [Kno71]. Brian [J.79i, K.71a, Ket75]. Bridge [ZNL79]. Brillinger [Weg76]. Brodeau [J.78m]. Bronchitis [Ens78]. Bross [Bon77]. Brown [K.71a, Hea78, Joh75]. Brownian [CF76, Nád73, PK76, SS74, Ser73]. Bruce [Rog70b]. Brumat [J.78i]. Brunk [Hal77b, Kal75]. Brunner [Jac80]. Bryant [J.76e, Hay73, J.76d]. Buddy [Rus73]. Building [Goo71]. Bulk [BR75]. Burdick [J.78j]. Burman [Wat70a]. Burr [Aus73]. Business [Ham75, J.78p, Joh75, Jon73, Kin79b, Hun70].

C [Bak74, Ban73, Boa79a, Cru78b, Dil71, Dun77a, For73, Gen77, Gil74, Gle78, GH75, Hay73, Hei79, Hil74, J.77b, J.77f, J.77e, J.77n, J.77m, J.77q, J.77p, J.78i, J.79d, Joh75, K.71e, K.73a, Kup71, Lyn78, Pre77, Sax72, Sen75, Sne72a, Ton72, Woo75, Woo74a]. Calculation [MT72, WB75]. Calculations [NL75]. Calibration [Kal71, MK70, MW75b, Nas78, Pep73, Shu72]. Calus [Boa79a]. Cam [Fre73b]. Candidates [Ano71f]. Cardiac [BKRS78]. Care [Gut73a, GDF78]. Carl [Gro77]. Carlo [Rel70]. Case [BR73, Dow73, Her70, KKS76, LO79, OL77, Sor73, Ton74, Ton75, WK77, JTS75]. Categorical [Dra79, Fed77, J.79c]. Category [Ala71]. Cauchy [CC71]. Causes [Moe74]. Cecily [Ens78]. Censored [Ano73k, Bai72, Bai73, BAB72, Coh65, Coh75, Coh76, CN77, Dan70, DM78, EB73, EB79, Hoo79, KB75, Lem75, MN77, MS79, MHL74, NH72, Nel72b, Nel73, NH73, NK76, NM78, NS79, Pat76a, SH79, Sin71, TW72b, Zel72, BD73]. Censoring [Man71b]. Centering [Bro77]. Central [HP78, KP79]. Centrosymmetric [And73, Goo70]. Certain [BK70b, Dav79a, FB75, GH70, Gri73, Sch76]. Chain [Bri74, JK74]. Chains [WB71, Cra74, Doo72, Sys76]. Change [SS75]. Changes [Dic74]. Character [Dol70]. Characteristic [SLK79]. Characteristics [KKM79, Koy79, SR75]. Characterization [Ham72b, WRLI75, Sen75]. Charges [CH74]. Charles [Car74, Ens78, Gil74]. Chart [Aus73]. Charts [Chi75, Dun78a, GW71, J.76g, Rey75, RH78, SWD70, SWD74]. Chatfield [Dal77, J.79s, Pre77]. Chatterjee [Hah79]. Checking [Fra77]. Chemical [BA70, Sch75, SLM74, TB75]. Chemistry [FH79, LSYF79]. Chi [Dye73, SR75]. Chi-Square [SR75]. Chinn [Mil72a]. Chistyakov [Spr79]. Choice [Ait74, Ano77h, Bli72, GZ79, J.76g, Mac76, MK70, YC71]. Choosing [BS72, GHW79, NV73a]. Chow [Sim73]. Chris [Zac74]. Christopher [Dal77, J.79s]. Chronic [Ens78]. Citation [J.77d, J.77o]. Civil [K.73e, Kup71]. Claim [Hah77]. Clark [Hei76]. Clarke [Rog70b]. Class [AvdL70, BLB73, Bur73, GPS73, HSL76, SP77, Sha78, Tak72, TK72, VM79]. Classical [Obe77]. Classification [CR79, Dun71, Ens76, HW75, Kus79, Low76, MA78]. Classifications [Goo71, HG73]. Classified [Dra79]. Claudia [Gro77]. Clelland [Joh75]. Clinical [FH79, LSYF79]. Close [MC72]. Closed [BM74, BM76, Ell76, PW77]. Closed-Loop [BM74, BM76, Ell76, PW77]. Cloud [DV78]. Cloud-Seeding [DV78]. Cluster [Ens76, Ket75]. Clustering [JK74]. Clusters [CBI79, YB76]. Cochran [J.78n, LH74]. Code [MBC79]. Codes [ZNL79]. Coefficient [Ano75e, Fil75, IM70, Kow71, Zei73, McC78a]. Coefficients [AR73, For72, JH73a, Wol77a]. Coherence [CP75]. Collected [K.73b]. Collection [J.76f]. Combination [Pap72]. Combinations [Coh72, Joh71a]. Combining [EP71, Lev70]. Comment [BC71, Cro71]. Commentaries [CGO07]. Comments [Cor75, HK77, Mal73, DV79, MK70]. Commission [Car77b]. Comparative [BJN71, Bon76a]. Comparing [AB71, Ber78, Jae70, MDM77, Nel72a, Rub76, Sal75b, SS76, BN72]. Comparison [BM78, CM71, DW73, Dub73, Fed75, FH79, GK77a, GK77b, GH70, Gri70, Gru73b, GWM76, HN70, HW75, Hin77a, Kam73, Kam77, KP71, LSYF79, Low70, Luc76, MBC79, MH78, MPB70, MA78, San73, TA75, Ury76, Vin76, WC78, HK77]. Comparisons [IM70, SdV73, SU79, Ury76, BN72]. Compartmental [Gri79]. Competing [Moe74]. Competitive [Nay72]. Complete [Ano73k, Bis78, Coh65, Coh72, EB73, EB79, MF77, MHL74, SW72]. Complex [Bri74, Hun74, Sne76, Cas76]. Component [EP71, GC79, GH79, Gru71b, HL76, JM79, JS76, Krz79, Low76, MG74, Sar71, Var70, Bha73]. Components [CS76, DS77a, GG70b, JH73a, JH79, MA78, Par78b, See70, SS78, TA71, Wol78, YC78]. Composite [HP78, Luc74]. Composited [Ede74]. Compound [BK70a, Cho79, Fal70, Boa73]. Compressed [Kam74]. Comprised [WK71]. Computation [CGO07, Fur71, Sch76, Som70, Yue78, K.71e]. Computational [Dav77, LH70]. Computations [J.77b, McC75a]. Computer [BH70, BCK61, Cha70a, Eva79, Gri72, Gro77, Hea78, Hub78, JB76, MBC79, Mil72b, Mit74b, Sie75, Wig76, BCK78, Mer72, Nay72]. Computer-Aided [Cha70a]. Computers [Cha70a, J.79q, Mul70]. Computing [HH73, KKM79, Mar76]. Concentric [CK79]. Concept [NGM70, Buc73]. Concepts [Bis72a, Car74, Sim76, SM76, Har77, Mih77, Mih75, Woo74a]. concerning [AK71, Law71]. Condition [Law77b, MH77]. Conditional [Ham72b, Jae74, Kap75, Sor71]. Conditions [Krz77, See73]. Conference [K.74e]. Confidence [AJ77, BS79, BWL78, Cha79, CM71, DV78, Dun78b, Eas78, EB78c, Fin76, FB75, Gaf78, Her70, Jae70, Jon71, Kam72, Kam73, Kam74, Kap75, Kus79, Lan72, Law78, LR78, Man70a, MF73, MG74, MF75a, MF77, MN77, Obe77, Pat76a, Ros75c, Sar71, Sch74a, SW75, Tar76, TBA70, TC75a, TC73, TS79, Ury72, TC72]. Conflicting [J.78b]. Conjecture [AK71]. Conover [Bha72]. Consideration [PDB77]. Considerations [GM79]. Consistent [IS76]. Constant [CK79, FR72]. Constants [Aus73]. Constrained [Ano76e, HB78, Sne75b, TW70]. Constraints [HS73].

Constructed [Dun78b]. Constructing [BCK61, BCK78]. Construction [CM71, EB78a, Law75a, Mit74a, Mit74b]. Consultants [Wat70b]. contain [HP73]. content [GPU76]. Contigency [MDM77]. Contingency [Gar71, Goo71, JK71, Odo78]. Continuity [Ham74]. Continuous [Che71, Fre74a, Hol75, RR76, Sac72, Sha79]. Continuously [Kao72]. Contour [GW71]. Contrasts [Par78b, Ury76]. Contributions [Ing78b, J.76a, J.78c]. Control [AB79, Ano77h, Aus73, BR79, BH78, Bin74b, Chi75, Dun78a, Dun77b, GW71, Gru73a, JM79, K.74a, Kab74, Luc73, Mac76, Pat73, RSO72, Rey75, ST73, SdV73, TW72a, BN72, Bis72b, J.79d, Kle77, Tsi78, Woo75, Ziz78]. Convergence [Ber72b, Sne70]. Convolutions [Bai74a]. Cooley [O’F73]. Coordinate [Kow70, Nád70a]. Cornell [Kup71]. Correct [Dun71]. Correcting [BH74]. Correction [AR79, Ham74, MM73]. Correlated [Shu71]. Correlation [Ano75e, AR73, BJ75, Fil75, Haw75, JB74, Kow71, Par70, Wol77a]. Corrigenda [Ano73f, Ano79e, Ano79f, Ano80a, Ano80b, Kit81, Mih77]. Corrigendum [Ano72g, Ano72e, Ano72f, Ano73g, Ano73h, Ano74f, Ano75e, Ano77a, Ano79g, Ano80c, Ano84]. Cosines [Cha70b]. Cost [Ano72g, Gue71b, RR76, Sch71b, ST73, TC75b, Ben76]. Costs [CH74, HS73, PS70]. Countable [Cra74]. Counting [CM71]. Counts [K.71d]. Course [Boa79a, Cro78, Geo72, Har70, Nas78, Sch74b, Str75]. Covariance [GC74, Ron77, Sne71, vN79]. Covariances [Cha70b]. Covariate [Har74]. Covariates [Kus79]. Covariation [J.76d, J.76e]. Cox [Ing75, J.79o, Cor75]. Criteria [Hoc72, J.78h, MPB70, Tho73]. Criterion [All71, Ano72e, Ano73h, Box70b, BD71, Bro79, Hel74, McL74a, ML75, NW79, SD77]. Critical [Zah75b]. Cross [Dra79, GHW79, J.79m, Mie74, Sim79, Wol78]. Cross-Classified [Dra79]. Cross-Over [Mie74]. Cross-Tabulated [Sim79]. Cross-Validation [GHW79]. Cross-Validatory [Wol78]. Crossed [HG73]. Cryogenic [Joi77]. CSP [Bla77]. CSP-1 [Bla77]. Cubic [LW71]. Cum [Qua74b, Qua74a]. Cumulative [Bec74, Rey75]. Current [Dun78a]. Curtailed [Coh70, HM76a, SP72, SP74, SP77]. Curve [LS71b, Sne72b, SLM74]. Curves [SLK79, Dav70]. CUSUM [BJ75, JB74, MW75a, GW71]. Cut [BJ70b]. Cutsem [J.78m]. Cyclic [Ano74e]. Cyrus [Sid75].

D [Bon77, Bow77, Cas76, Dud79, Ens77, Fin75, Gib74, Gor76, Gou73, GH75, Hal77b, Har72, HK77, Ing75, J.75g, J.76h, J.77j, J.77k, J.78l, J.78k, J.78e, J.78d, J.78b, J.78o, J.79e, J.79o, J.79m, Joh73c, K.74c, Kal75, Law75b, Str75, Sys76, Wil70]. Daniel [Sne72a, Wei71, Lyn78]. Data [All74, Ano71a, BH70, Bli72, BHME73, BM76, Cha79, CR79, CG76, Dav77, Dyk71, Ell76, EB79, Ens76, Fed74b, Har75, HJ76, Hoo79, J.78f, J.78d, Jac80, JK71, KN74, KB75, Mal74, Man71a, MG74, Mar74, MN77, Mul70, Nea70, NH72, Nel72b, Nel73, NH73, O’F73, PW77, Pat76a, PS77, RTDM79, SH79, See73, Sne72b, Sne73, Ten72, WB74, Weg76, Wol74, Woo73, Zel72, BT74, Boa73, Dra79, Fed77, Gen77, J.77m, J.79i, J.79c, Sim79, Sne72a, Wat73, Gen77, Lac71a]. David [Cra74, Doo72, For71, Kin79a, Mil72a, Ser73, Sim73, Weg76, Ano74e, Geo72, Gru71a, J.78c]. Davis [Ano80a]. Dealing [GS71, Gut73b]. deCani [Joh75]. Decays [SCSB76]. Decision [J.78h, Kin78, Kup71, Lac71b, Sym71, Win71, Hol73]. Decisions [J.78b, Ran70b]. Decomposition [Bod70, Fre74b, SCSB76]. Defective [BR73, SP74]. DeFrancesco [Sto76]. DeGroot [How77]. Delay [CP75]. Deliveries [BS72]. Delphi [J.76i]. Densities [BMP77, Par70]. Density [Gro71b, Spe71, Weg72, WRLI75]. Departure [Gre71]. Depend [Wio70]. Dependent [AvdL70, Bri74, Kus79, Moe74]. Deposits [CBI79]. Depreciation [CH74]. Derivation [Lad77]. Derivative [RJ78]. Derivative-Free [RJ78]. Derivatives [JB76, Wig76]. Deriving [Jon76a]. Derman [Sid75]. Derrick [K.74d]. Design [Add72, Ant77, Box70b, Box71, Buc73, Car74, Chi75, Con70, Cru78b, Dun78a, FM71, Gri73, HH74, Hil78, HS71, J.75a, J.76b, Joh78, Joi77, Luc76, Man72, Mar75, Mar76, MHC77, MH78, Mie74, OL77, PB77, SN77, See73, Ste75, Wat70a, Zac70, ZNL79, Joh73a, Kev75]. Designing [Joh71a, JC76, Tak72]. Designs [AT79, Ano72e, Ano73h, Ano73f, Ano76e, Ber72a, BD71, BD74, BH77, Cor75, DS77a, Ede74, Eva79, GK77b, GG70b, HP78, HW75, Har79, HM72, Hok74, IR77, Joh79, Jon76a, Kle75, Luc74, Luc76, Luc78, Mit74a, Mit74b, MB78a, MB78b, MA78, OM72, Par78a, Roq76, See70, SM74, Sne75b, SM76, SC71, SD75, Ste72a, Tho73, TA75, Web71, GK77a, Ano74e, Cam77, Sne76]. detect [GDF78]. Detecting [CBI79, GW75, Gut73a, Shu71]. Detection [Ano79g, Coo77, Dav79b, Kit79, LE73, Ros75b, TM72, Kit81]. Determination [Ano72g, GW71, Gru73a, Gue71b, Hal77a, Jae74]. Determines [Nád71]. Determining [AB71, MW78]. Development [Kee75]. Developments [J.78i, J.78m, J.76b, J.79g, J.79f]. Deviation [BMH76]. Deviations [For72]. Devices [HN70]. Diagnostic [BG78, DM77]. Difference [Hah77]. Differences [Sal75b]. Different [Par70]. Differential [Bri73a, Cha75a, Fin75, JB76, Sac75, Str75, Wig76]. Differentiation [AB74]. Diffusion [Ser73]. Digamma [Lin79]. Digital [J.79q, Mih75, Mih77, Nel74]. Dimension [Cha79, vNS76, vN79]. Dimensional [GLV74]. Dimensions [DL70]. Dimitri [Ziz78]. Direct [Aro76, Goo71]. Directional [Wat73]. Disciplines [J.79v]. Discounting [RR76]. Discrete [AvdL70, Ano77h, Bai74a, DY70, Gri72, Mac76, Mih75, Mih77, PS77, Spr78, Odo77, Smi70a]. Discretization [Hol75]. Discriminant [Dun71, HH77, Krz77, Lac74b, Lac79, McC75a, McK76, McL72, McL74b, Sor73, vNS76, vN79]. Discrimination [AR79, BA70, DA73, DAH73, Har73a, Hil78, MPB70, Ng76]. Discussion [Ant77, Eas78, Ell76, FH79, Har73a, HM71, Hoo79, Law77b, LvNN73, Luc78, Mal74, Mar74, McD75, Par72, Sne75a, Wig76, Zel72]. Disease [Ens78]. Disney [Rog70b]. Displays [DM77, Tuk70]. Distributed [Bli72]. Distribution [AA74, Ang74, Ano72g, AJ77, Aus73, Bai72, BE73, Bai73, Bai74b, BB79, BL71, Beh72, BMH76, BH73, CC71, CF76, CF77, Chi74, Coh65, CH73b, Coh73, Coh75, Coh76, CN77, CJ73, CL70, CL71, Dan70, DH74, DM78, Dun74, Dun78c, Dye73, EB73, EB74, Eng75, EB77a, EB77b, EB78a, EB78c, Fau70, FD70, GKS75, Gru71b, Gru71c, Gue71b, GS71, HBA71, Hah75, Ham72a, Har71, Has72, Haw72, Hel71a, Her70, HR75, Jos72, KN74, Kap75, KARR75, Law71, Law73, Law77a, Lem74, Lem75, LNN61, LC70, Lik74, Man72, MF73, Mat78, McD76, MW75a, MCL76, Nád70a, Nád71, O’B73, Par70, PK79, Pie73, RS71, RTDM79, RSO72, Riz71, SG71, Sar71, SA79, SLK79, SW72, SB72a, SW75, SU79, TBA70, Ury72, Wol77a, YM74, GPU76]. distribution [KP79, Smi76, TC72, Bel70]. Distribution-Free [HR75, MW75a, Ury72, Wol77a, Bel70]. Distributional [FP77]. Distributions [BL71, Beh70, Bry74, Bur73, CB75, Cha79, Che71, Cho79, DG72, Dub73, DA73, Eas78, Eva72, Fal70, GG70a, Gri70, Gro72, Hei76, HB71, Hol75, J.79u, KB75, KP71, Law75a, Law78, LF79, MF77, Nel75, NK76, NM78, NS79, Pat73, Pat76b, PS70, Rel70, Ron77, Sha78, Tak72, TW72b, WDFH73, Zac74, Fre74a, Smi70a, Che71, Fre74a]. Divisibility [Luk73]. Dixon [Hea78]. Dodge [KO77]. Does [Bad70, Bor72]. Dolby [Qua74b, Qua74a]. Don [Hac78]. Donald [Boa79a, Hay73, Leh71, Ran70c]. Donna [Boa79a]. Dorfman [GR72]. Double [ABA72, BE73, Bin74b, BR75, BH73, Eld74, Gue71a, Joh70a, Kap75, RS71, Ten72, Zac70]. Double-Stage [ABA72]. Doubly [HS71]. Draper [And70, GK77a]. Drewes [Leh71]. Drift [CF76, Nád73, PK76, SS74]. Drogin [Car77a]. Dual [MC73, ST73]. Duckworth [Wes71]. Dud [RJ78]. Dudewicz [Zie78]. Dunkelberg [Gle78]. Duplicate [TB75]. Dynamic [BM74, TW72a, Kev75, Sys75]. Dynamic-Stochastic [BM74].

Each [Gru71b, Joh70c, Joh70b, McL74b, Sar71]. Earl [Fin75, Str75]. Early [Ens78]. Easy [Gue77]. Echo [HT74]. Econometric [For71]. Econometrica [Fre73a, K.73c]. Econometrics [Fre73a, J.75h, Goo75, Goo76, Han73, Wil79a, J.77c]. Economic [BA70, Chi75, Dun78a, K.73c, Kao72, Mer72, Kee75]. Economical [Bli72, Gua71, Hok74]. Economics [Jon73]. Économiques [Gua71]. Editor [HK77, JTS75, And76, And77, Ano77g, BJ70a, BC79a, BM72, Fle71, Gar74, GMD^{+}75, GH75, HOC^{+}77, Ham70, Haw78, HJ73, JEBE74, Kru70, KBW72, Law76, LSP76, Lee79, MP72, Saw70, SBV79, Smi70b, Smi76, SM79, VK76, Wil79b, Wu73, ZA75]. Educational [Ger75]. Edward [Hay73, Woo74b, Zie78]. Edwin [J.77l]. Effect [BJ75, Bla77, CDR79, DR77, Dic74, JB74, RSO72, SG71, SOE75, TA71, Wic72]. Effects [HW75, Lac79, Nai74, Woo73, vNS76, vN79]. Efficiencies [BR73]. Efficiency [Igl73, LH70, MHL77a, MHL77b, PS70]. Efficient [BL74a, Her70, Jos72, KKS76, MF75b, MF77]. Ehrenberg [Gen77]. Ehrenfeld [Ano77a]. Eigenanalysis [Haw75]. Eight [Ens78]. Eighteen [Joh78]. Einfuhrung [Ens77]. Elderton [Dav70]. Eleanor [K.71a]. Elementary [Fin75, Hac78, Har77, J.76c, Mye74]. Elements [WK71].

Elimination [LS71a, Nas78]. Elizabeth [Fre73b]. Elliot [Hei78]. Ellipsoids [Tar76]. Emergency [BS72]. Emphysema [Ens78]. Empirical [Dun78b, Joh73b, Lev70, Zah75a]. Encyclopedia [J.79k]. Enduring [You72a]. Energy [J.79h]. Engineering [Cru78b, Gra72, J.78p, J.78j, Ziz78, K.73e, Sch75, K.71b, Web73]. Engineers [Boa79a, Fed74a, Rog70b, Cas76, Kup71]. Enhancing [CK75]. Enochson [Nel74]. Enrick [Kab74, Rus73]. Enslein [J.79q]. Entropy [Gri72]. Enumeration [BJ70b]. Envelope [CP75]. Equal [AR73]. Equality [BF74, CH73a, GC74, WDFH73]. Equation [SD77, Ell78]. Equations [And73, Bri73a, Cha75a, GM79, JB76, Joh73b, Sne72a, Wig76, Woo73, Fin75, Sac75, Str75]. Equipment [JR72]. Errata [Ano70f, Ano71g, Ano75f, Ano77f, Ano78e, BCK78, Ton75]. Erratum [Ano73k, Ano73i, Ano73j, Ano76e, Ano76f, Ano77g, Ano77h]. Erroneous [Gru73a]. Error [All71, Bus71, Cha72, Cha70b, Cor77, DW72, Eva71, Far75, GR72, HH77, Low74, McL74a, McL74b, ML75, SA79, Vin76, Wic72, HK77]. Errors [Bec74, Gru73b, McL74a, NW77, NW79, Pep73, SB72b]. Erwin [Cor71]. Essay [Bon76b]. Essays [K.74b, J.76f]. Estimate [Bri73b, Law77b, MH77, MH78, MA78, SP74]. Estimated [Hel74]. Estimates [BMP77, Cor77, GH70, GC79, HH73, Law75b, Low74, MF73, MS73, MH72, Nea70, Pas79, SG71, Spr76, McC78a]. Estimating [BL74b, Cha72, DS79, GG70b, HW75, HT74, Han75, JMM78, McL74b, Nád70a, OM72, Sal75b, SG71, See70, Sor71, Sor72a, Sor73, Ten72, Zac70]. Estimation [ABA72, AT73, AA74, AR79, Ano73k, AR73, BE73, BC71, BD72a, BS71, Bec74, BMH76, BK70a, Boa73, Boa79b, Box70a, BM76, BL73, BR75, Bro79, Bus71, Cha79, CH70, CP75, Coh65, Coh72, CH73b, CS76, D’A71a, D’A71b, Dan70, Dow73, DG71, Dye73, Eas76, Eas78, Ell76, EB74, Eng75, Fal70, For72, Fre75, Goo71, GH79, GS71, GPS73, HBA71, Hal70, Ham72a, HG73, Has72, HL76, HB78, HP71, HK70b, HG78, IS76, J.75d, JS76, Jos72, KS71, KKS76, KA71, Lan72, Law73, Law78, Lem75, Man71b, Man75, Mar70, Mat78, MW73, McL74a, NH72, NH73, O’B73, Pas79, Pat76b, RS71, Rah76, Rea71, Ron72, San72, SA79, SP77, Spe71, SWD70, SWD74, SS78, TBA70, TW72b, Ton74, Ton75, TS79, Var70, WK71, Weg72, Wic72]. Estimation [Wio70, Wol78, WK77, Zel72, DV79, JTS75, Cha75b, Rog70a]. Estimations [WRLI75]. Estimator [Far75, MC72, MRR78, Vin76, Whe70, HK77]. Estimators [BR73, GWM76, HSL76, Jae76, LaM78, Lev70, Low76, Low74, Man70a, MF75b, MF77, MHL74, MHL77a, MHL77b, Sch76, SB72a, Ste72b, TA75, WC78, Wil79a, Zei73]. Eugene [Lan73]. Euler [Lin79]. Evaluate [Ang74]. Evaluating [Lin74, Lin79]. Evaluation [GM79, Hay73, Joi77, Lan72, McC78a, Mil72b]. Event [Mih75, Mih77]. Everitt [J.79i, Ket75, Odo78]. Evidence [Ing78a]. Evolutionary [And70]. Exact [BD78, BK70b, Her70, JB76, Krz79, Man70a, Sam78, Sar71, SS74, TBA70, TJ79, Wig76]. Examination [Ano71f]. Examining [Sne71]. Example [Joi77, Hah79, Lac74a]. Examples [Sne77]. Exceedance [Kap72].

Expectations [Ham72b, Sim73]. Expected [BK70b, Dun71, KS71, Law77b, MH77, Sor72a, Sor73]. Experiences [Box70b]. Experiment [AT79, Ano73f, DV78, HS71, J.76b, Web71]. Experimental [Ano71a, Ano76e, Ant77, Box70b, Box71, Cru78b, Dyk71, Eva79, FM71, Har74, HS73, Hil78, Joi77, MHC77, Mit74a, Par78a, PB77, See73, Sne75b, Sne76]. Experimentation [Kle75, You72b, Lyn78]. Experimenters [Add79]. Experiments [Cor73, Cor79, Dav79a, Har79, HH74, HS71, J.76g, Joh71a, JC76, Man72, Mar76, Mer72, Pre71, San72, San73, SM76, Sta79, GK77a, Kev75, Add72, Car74, Gri73]. Explained [Wio70]. Explicit [Hem75, HB78]. Exploration [SN77]. Explorations [Ber73]. Exponential [BE73, BK70a, BH73, CH73b, DG72, Dub73, Dun78c, EB78c, Gro71a, Gru71b, Gru71c, Gru73a, Hah75, IR77, Jos72, Kam72, Kam73, Kam74, KN74, Kam77, KKM79, Kap75, KKS76, KARR75, KP71, Law71, Law77a, Lik74, Mat78, NL75, Pie73, RS71, SG71, Sar71, SLK79, SW72, SCSB76, Ton74, Ton75, Var70, WDFH73, Bha73, Boa73, GPU76, JTS75]. Exponentiality [EB75, Ng76, Ste78]. Exponentially [Bli72]. Extending [Nea70]. Extension [Ano73g, GR72, GB72, Vin76, HK77]. Exton [J.77g, J.79j]. Extracting [Fed74b]. Extraneous [Pas79]. Extreme [Bai72, Bai73, Eas78, EB73, EB74, Eng75, EB77a, Has72, Law75a, Law78, MF73, MF77, NM78, Sin72, SM74, TW72b]. Extreme-Value [Bai72, Bai73, EB73, EB74, Eng75, EB77a, Law75a, MF73, MF77, TW72b]. Extremes [SG71, ZNL79, Ber72b].

F [Ano74e, Geo72, Gro77, J.75f, J.76c, J.78g, J.78m, J.79e, J.79r, J.79b, Jon76b, K.71a, Kra76, Lac74a, Law75b, Luk73, Man71c, MCL76, Sne72a, Spr79, Sto76]. Factor [Dic74, Kle75, Sch71b, Wol78]. Factorial [AT79, Ano72e, Ano73h, Ano73f, Ber72a, BD71, Hok74, Joh79, Luc78, Mar76, MB78a, MB78b, SC71, Ste72a, Web71]. Factorials [CDR79, TW70]. Factors [Ano73f, Dun78c, HS73, Man75, Web71, GPU76]. Fail [Gru71b]. Failure [AJ77, Bai74b, BK70a, Bry71, CB75, FR72, JR72, Kam77, Law77b, LE73, Man72, MH77, Moe74, Nel72b, Sar71, Boa73]. Failure-Rate [Bai74b]. Failures [Bri74, Gru73a]. Fairley [Boa79a]. Faliva [J.75d]. Farhi [Gor76]. Fatality [Ano80c, McD79]. Fatigue [Sch74a]. Fault [J.77j]. Faultless [Wan71]. Features [Gua71]. Feedback [AB79, Bak74]. Feller [Sam73]. Few [Man70a]. Fiducial [Gru71b, Gru71c, Pie73]. Field [J.76b]. Fienberg [J.75h, Odo77, Dra79]. Films [Ziz78]. Financial [Ano74g]. Findley [J.79e]. Finite [Gro72, Ray70, WB71, Zac70, Ham75, Spr72]. Finney [Gou73]. First [CF76, Geo72, Gre70b, Mit74b, Nád73, PK76, Sch74b, SS74, SWD74, Weg74]. First-Order [Mit74b]. Fisher [J.79p, K.73b, J.78l, K.73b, Krz77, LH74, Ros75c]. Fishman [Mih75, Mih77]. Fit [Ano72f, DG72, DH71, Eas76, LaB77, MF75a, MS79, PS77, WB74, LR78]. Fitting [BDH70, Cle71, Har73b, HM73, JB76, Mal74, Mar74, Par78a, PDB77, RTDM79, Wig76, Woo73, BT74, Sne72a]. Flaw [Lem74]. Fleiss [Con74]. Fletcher [Ens78]. Flight [Bur73]. Flow [Joi77]. Folded [LNN61, Riz71]. Following [McD75, Obe75]. Foods [Car77b, Car77b]. Forecasting [Kle77, LR76, Dun77a]. Foreign [Kee75]. Forest [K.74e]. Form [SP72]. Formulas [Gue74]. Formulation [LvNN73, NV73b]. Foundations [Sim76, Rog70c]. Four [DS77a, HW75, Ury76, Boa79a]. Fox [Boa79a, J.76g]. Fraction [BR73, SP74]. Fractional [Ber72a, SC71]. Fractions [Hok74, Luc78, MB78a, MB78b]. Francis [Joh75]. Frank [Ens78, Smi77]. Franklin [Zys70]. Franks [Sch75]. Fraser [Wil70]. Fred [Cru78b]. Frederick [Odo77, Qua75]. Free [HR75, MW75a, RJ78, Ury72, Wol77a, Bel70]. Freedman [Kin79a, Cra74, Doo72, Ser73]. Freeman [Sne76]. Frequency [CP75, Dav70, Nea70, Pie73]. Fresh [McC72]. Freund [Lac72a]. Fried [Cru78a]. Front [Ano70g, Ano70h, Ano70i, Ano70j, Ano71h, Ano71i, Ano71j, Ano71k, Ano72h, Ano72i, Ano72j, Ano72k, Ano73l, Ano73m, Ano73n, Ano73o, Ano74h, Ano74i, Ano74j, Ano74k, Ano75g, Ano75h, Ano75i, Ano75j, Ano76g, Ano76h, Ano76i, Ano76j, Ano77i, Ano77j, Ano77k, Ano77l, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano79h, Ano79i, Ano79j, Ano79k]. Fuller [McC78b]. Function [Gre70a, HM76a, J.79v, Krz77, Lac79, Lin79, McL74b, ON75, Rou73, SA79, Sor73, Spe71]. Functions [Cha72, DV78, Gro71b, J.77g, K.74d, Kow71, Man71c, Nád70b, Wol74, Hel71a, Jer70, Rus73]. Fundamental [Car74]. Fundamentals [Eas72]. Further [PDB77, SB72a]. Fussell [J.77j, J.78j]. Future [AM79, FM77, Hah75, Hah77, HP73].

G [Add79, Bis72b, Boa79a, Ehr77, Eri72, Fie73, Gil74, GH75, Ham75, Har70, Har75, Hun74, J.76c, J.77l, J.78n, J.78e, J.78m, J.78j, Jon73, K.71d, K.74c, Kle77, Kno71, Lac74a, Man71c, MK70, McC72, Mil72a, Ram73, Ran70c, Sax72, Sch75, Sim79, Ton72, Tsi78, Wil79a]. Gabriel [Ano80a]. Gambling [Bon76b]. Game [Tan70]. Gamma [Ano73k, BG74, CN77, DG72, EB77b, EB78a, Gro71b, HBA71, Joh70b, O’B73, Ron77, SB72a, WK71, GH75, LF79]. Gani [J75e]. Gap [HD70]. Gary [Mar75]. Gass [Ran70a]. Gaussian [BB79, CF76, CF77, PK79, WY78]. Gene [CGO07]. General [AvdL70, AS73, Eva72, Hel74, SB72b, Tar76, TC75b, Bha73, GDF78]. Generalization [CJ73, HK75, ML75, Ell78]. Generalizations [Nel75]. Generalized [Aus73, BL71, GG70a, GHW79, HBA71, Hem75, HB78, Mar70, Sal75a, Ban73, Gow72, Urq73]. Generated [JB76, Wig76]. Generating [Ron77]. Generation [BJ70b, Jae74]. Gentle [Smi76]. Geo [Jon76b]. Geo-Science [Jon76b]. Geomathematics [Jon76b]. Geometric [AA74, RH78, Ste72b, Wic72]. George [And70, Fie73, Hil74, K.73d, Mih75, Mih77, Wil79a]. Gerald [Buc73, Dun77b, Web73]. GERT [KO77, Whi70]. Gestion [Gua71]. Gibbons [Dud79, Joh73c]. Gibbs [Bry71]. Gibbs-Type [Bry71]. Gibson [J.79m]. Gipps [GH75]. Giri [Hei79]. Given [Cha79, Hal77a]. Glenn [Ing78a]. Gleser [Sid75]. Gnanadesikan [Gri73, J.78f]. Golub [CGO07].

Good [GHW79, Ano70f]. Goodman [Spr72, J.79c]. Goodness [DG72, Eas76, LaB77, MF75a, MS79, PS77, LR78]. Goodness-of-Fit [LaB77, MF75a, PS77]. Gottfried [Mer77]. GPSS [Gie74]. Gradients [CO75, Sne75a]. Graduate [Har70]. Graham [Sim79]. Graph [Fed75]. Graphic [BH70, Tuk70]. Graphical [Ang74, CK75, Fed74b, Fed75, J.79i]. Gray [Sal75a]. Graybill [Zys70]. Great [Sim73]. Greatest [Kap72]. Green [Man71c]. Greenberg [Gie74]. Gregory [Kin79b]. Groot [Ran70b]. Gross [Hei76]. Grossman [vdV75]. Group [BR79, Ehr72, GR72, K.74e, PBT76, Sam78]. Group-Sequential [PBT76]. Group-Testing [Sam78]. Grouped [Fre75, PS77, Boa73]. Grouping [MCM75]. Groups [J.76d, J.76e, McK76]. Growth [Kee75, LR76, Rea71, WG78]. Grubbs [Ano79g, Jae76, TM72]. Grubbs-Type [Ano79g, TM72]. Guenther [GH75, Woo74a]. Guide [Sid75, Boa79a]. Guillermo [Tan70]. Gumbel [Her70]. Gupta [J.75b, K.74c]. Guttman [Gra72]. Gwilym [Ran70c].

H [Ano74e, Bow78, CGO07, Fed74a, Hal77b, Har72, HK77, How77, J.75g, J.76d, J.76e, J.76i, J.76h, J.78c, J.78b, J.79q, J.79k, J.79v, Jac80, JTS75, K.73b, K.73a, Kal75, Kin79b, Kra76, Lac74a, Law75b, Man73, Mer72, Ran70b, Ros75a, Sne74, Sne76, Ste75, Sto76, Sym71, Web73]. H-Function [J.79v]. Haan [Ber72b]. Haber [Qua71]. Haitovsky [Fre75]. Haley [J.79n]. Half [Zah75a, Zah75b]. Half-Normal [Zah75a, Zah75b]. Hamburg [Kin78, Win71]. Hamdan [Ano75f]. Hampel [Law75b]. Handbook [J.78g, K.71b, J.79j]. Handling [Gut73a, GDF78]. Hannan [Cle72a]. Harald [K.74b]. Harold [Dyk76, J.77g, J.79j, Sne76, Sta70]. Harter [Har72, J.75g, J.76h, Ano70f]. Hartley [Sne74]. Having [AR73, BL71]. Hays [Lac71b]. Hazard [Dan70, Jon71, Nád70b, Nel72b, Sha78]. Hazurbazar [Ing78b]. Health [J.79h]. Heavy [Bry74]. Heavy-Tailed [Bry74]. Helge [Boa79b]. Helmuth [Ens76]. Hemker [K.73a]. Henri [Fre74b, Han73]. Henrick [Sch74b]. Henry [Sal75a, Woo74b]. Henschke [Gro77]. Herbert [Gru71a, Sim73]. Hess [K.73a]. Heteroscedasticity [PB77, PDB77]. Hickman [Ram73]. Hicks [Car74]. Higher [Nai74]. Higher-Way [Nai74]. Highway [Ano73j, Mar73]. Hilary [Hun70]. Hill [J.78q, J.78q]. Hilton [Ram73]. Hine [Boa79a]. Hinkley [Ing75, J.79o]. Hirsch [J.79r]. History [Ens78, J.78k]. Hoel [Gil74, Jon73, Sax72, Ton72, Eri72]. Hoerl [Ano70f]. Hogg [Hei78]. Holland [Odo77]. Hollander [Gib74]. Homogeneity [Cho79, Sch71a]. Honor [J.75h, K.73d]. Honour [J.77l, K.74b]. Hooper [Fre73a]. Hordijk [Sys75]. Hotelling [JH79]. Howard [Hac78, Har70, Lut73]. Huang [For71]. Huber [Law75b]. Human [Bon77]. Hunter [Add79, Add79, Gra72]. Hurst [Dun77b]. Hybrid [Roq76]. Hyperbinomial [Gro72]. Hyperbola [WB74]. Hypergeometric [Hel71a, J.77g, J.79j]. Hypotheses [BH73]. Hypothesis [Leo70, McD75, Obe75, SLK79]. ibid [GH75]. Identifiability [Gri79]. Identification [Ray72, Kev75]. Identified [Nád71]. Identifying [Joh70c, Joh70b]. II [BCK78, Ens77, BJ75, BN72, BCK61, Eva71, GS71, Lac74b, NH73, TW72b, J.75g]. III [Cra74, Eva72, Har71, J.79q]. Illustrated [Mal74, Mar74, BT74]. Implications [BH70, Wil79a]. Implicit [BJ70b, BL73, Sac76]. Implies [Bad70, Bor72]. Important [JC76]. Imprecisions [Han75]. Improved [BG74, Boa79b, Box70a, HL76, SS71b, Sta79, GH75]. Improving [BWL78]. Including [J.77b]. Incomplete [Ano73f, HS72, OL77, Web71]. Incorrect [JR72]. Increases [LE73]. Increasing [And70, OB76, Sch71b]. Increments [AvdL70]. Independence [Cho73, Kra76]. Independent [Bai73, Lev70, McD76, SB72b, Bha73, LH77]. Index [Ben76, K.71a, J.77d, J.77o, Qua74b, Qua74a, J.77d, J.77f, J.77e, J.77o, J.77q, J.77p]. Individual [TS79, Woo73]. Induction [Bon76b]. Industrial [And70, Lyn78]. Inefficient [Ste72b]. Inequalities [HB71]. Inequality [Wan71]. Inference [Ait74, BAB72, Bis79, EB78a, FW72, Goo75, Goo76, Hil74, Hol73, Kal75, Lac71b, LL78, NS79, Pie73, Sin71, Sin73, Bon76a, Gol74, Hei78, Joh73c, K.71c, K.74c, Sta70, Sym71, Wil70, Woo74a]. Inferences [Bai72, TA71]. Infinite [Luk73]. Influential [Coo77]. Information [Ano70o, Ano71l, Ano72l, Ano73p, Ano74l, Ano75l, Ano76o, Ano77q, Ano78o, Ano78p, Ano79p, EP71, Fed74b, Hay73, Lin74, Sim76]. Infrared [WRLI75]. Ingram [Sid75]. Inhomogeneity [BH74]. Initial [Lac74b, Lac79, McL72]. Innovation [Dav79b, Sys78]. Input [MBC79]. Insect [HG78]. Insensitive [Bri73b]. Inspection [KO77, Lem74, Ten72, Tsi78]. Instabilities [MS73]. Instantaneous [BS72]. Institute [Ano71f, J.75g, J.76h, J.77k, J.78o]. Institutions [Jac80]. Instrument [Jae76, Mul70]. Instruments [Gru73b, Han75, Jae76, Pep73]. Integer [Lin75, Sha71]. Integral [Mil72b]. Integrals [J.79j]. Integration [Eva71, Eva72]. Intelligence [Hel71b]. Interaction [HJ76, YC78]. Interactions [Nai74]. Interactive [BH70, BH77, Mar76, J.78d]. Interdependence [Gor76]. Interest [CHH^{+}71, CH74]. International [Car77b, J.79k]. Interpretation [CO75, FB75, Sne75a, SWD70, SWD74]. Interpreting [Gen77]. Interregional [Kno71]. Interval [AB79, Ano77h, BE73, Cha79, Eas78, Hah77, Jae70, Jon71, Kus79, Lan72, Law78, Mac76, MF75b, O’B73, TS79]. Intervals [AM79, AJ77, BWL78, CM71, DV78, EB78b, FM77, Hah72, HM76b, IS76, Kam72, Kam73, Kam74, KN74, Kam77, Kap75, Law77a, MF77, Pat76a, Ros75c, SS71a, TBA70, HP73]. Intra [Ven70]. Intra-Year [Ven70]. Introduction [Goo75, Goo76, Hac79, Hal77b, Jer70, Ros75a, Sam73, Sym71, Zac74, Dyk76, Lan73, Bri73a, Cru78a, Dye74, Eri72, Gil74, Gro77, Hei79, Hol73, Hub78, McC72, McC78b, Mer77, Sax72, Sta70, Ton72, Woo74b, Zie78, Zys70]. Introductory [Man71c, Bow78, Cor71, Gra72, Joh71b, Lan75, Mil72a]. Invariant [KARR75, Man71b, MF73]. Inventory [K.74a]. Inverse [Ban73, BB79, Bro79, CF76, CF77, Cle72b, DS77a, DS77b, Goo70, Gow72, Hal70, Par72, PK79, Ray70, TS79, Urq73, WY78]. Inverses [Mar70].

Inversion [Fra77]. Investigation [GS71]. Invited [HM71]. Involving [And73, DW72, Har79, McK76, Riz71]. Ionospheric [HM73]. Irregular [Hok74]. Irving [Sym71]. Irwin [Bon77, Gra72]. Isaac [Bon76b]. Iterative [Cha72, HG73, Pre71]. IV [AT79, J.77h].

J [Add79, Ano74e, Bha72, Bis72a, Boa79a, Bon77, Bow78, Cle72a, Cro71, Dud79, Dyk76, Gil74, Gol74, Gou73, Gra72, Gri73, GH75, Ham75, Hea78, Hei76, J.75h, J75e, J.75a, J.75f, J.76b, J.76a, J.77f, J.77e, J.77l, J.77j, J.77q, J.77p, J.78p, J.78m, J.78j, J.79u, J.79m, Jer70, Jon73, K.71b, K.73b, K.73c, Kal75, Kle77, Law75b, Odo77, Ram73, Rog70d, Sax72, Sch74b, Sha71, Sid75, Sim79, Smi76, Spr72, Sta70, Ton72, Web73, Zie78]. Jack [Kup71]. Jackknife [Dun78b, GC79, Sal75a]. Jackknife-Constructed [Dun78b]. Jackknifing [Hin77b]. Jackson [Ano80a]. Jagers [K.74b]. James [Bel70, Qua74b, Qua74a, vdV75, DV79]. James-Stein [DV79]. Jan [Bak74, Hun74, K.71b]. Jean [Joh73c]. Jenkins [Ran70c, Kle77]. Jerome [And75]. Jerzy [Fre73b]. Jessen [Jon73]. Joan [J.79p]. John [Ano74e, Boa79b, Boy74, Cas76, Fre73a, Gle78, J.77b, J.77d, J.77o, Joh75, K.71e, Lac72a, Leh71, Qua74b, Qua74a, Add72, Cro71]. Johnson [Che71, Cru78b, Fre74a, Smi70a, Spr78, Dav70, Dun77a, Kin79b]. Joiner [J.78g, K.71a]. Jonckheere [Ode71]. Joseph [Con74]. Journals [K.71a]. Judah [Hog74]. Judge [Wil79a]. Judgments [Gru73a]. Judith [J.79k]. Julius [Har75, Hog74].

Kabe [K.74c]. Kagan [Sen75]. Kali [Ben76]. Karl [Jac80]. Kathleen [J.75c]. Kay [Ber73]. Kececioglu [Ziz78]. Keeney [J.78b]. Kemeny [Ham75]. Kendall [Fie73]. Kenneth [K.73c, Sch74b]. Kernel [BMP77]. Kimble [Kin79b]. Kinetic [PDB77]. King [Hay73, Kno71]. Kleindorfer [J.79d]. Knowledge [ABA72, Jac80, Rah76]. Known [MD71, Ste78]. Kocherlakota [J.75c]. Kodell [Smi76]. Kolchin [Spr79]. Kolmogorov [BD73, DM78, LR78, PS77]. Kolmogorov-Smirnov [BD73, LR78]. Konijn [Ste75]. Koosis [Boa79a]. Kossack [Gro77]. Kotz [Che71, Fre74a, Smi70a, Spr78]. Kreyszig [Cor71]. Krishnaiah [J.77h, J.78a, J.79g, J.79f]. Kruskal [J.79k, Lac74a]. Krutchkoff [MK70]. Krzyz [Hun74]. Kshirsagar [Ing78b, Kha74]. Kumar [Ban73, Ano71g]. Kurtosis [JL79, Rou73].

L [Ber72b, Che71, Con74, Cro78, Cru78b, Dav70, Dun72, Eas72, Fed77, Fre73b, Fre74a, Gow72, Gua71, Ham75, Har72, Hol73, Hun70, J.75g, J.76b, J.76h, J.78l, J.78g, J.78b, Joh71b, K.71a, Kab74, Lac71b, Lan75, Mer79, Qua74b, Qua74a, Rog70b, Ron72, Rus73, Sal75a, Smi70a, Smi76, Spr78, Wei71, Woo74b]. L. [GW71]. Laboratory [TB75]. Lack [Ano72f, DH71]. Ladany [J.77i, J.78e]. Lagrangian [Gre70a]. Lai [K.71c]. Lancaster [Ros75a]. Large [AM79, Cha79, CR79, Gut73b, HD70, Jae76, WG79]. Larger [Lad77].

Largest [Aus73, RRH73]. Larrieu [Gua71]. Larsen [Boa79a]. Larson [Dyk76, Sta70]. Latent [GWM76, WGM74, WG79]. Latin [BCK78, BCK61, Hed71]. Laubscher [K.71a]. LaValle [Sym71]. Law [Sin72]. Laws [Sen74]. Layout [Dan78]. Layouts [Nai74]. Leak [PBT76]. Leak-Testing [PBT76]. Leaps [FW74]. Least [AF76, BD72a, FM77, GZ79, GWM76, GPS73, Har73b, HM77, J.77b, JK71, MHL77a, RJ78, See73, Wat74, HP73]. Left [Man75]. Lehmann [Mer79]. Length [Lem74, Rey75]. Lengths [RH78]. Lennart [K.74b]. Lentner [Mye74]. Leo [J.79c]. Leon [Sid75]. Leonard [J.75h]. Leone [Ano80b, Cru78b]. Less [Fur71]. Lester [Bow77]. Letter [And76, And77, Ano77g, BC79a, BM72, Fle71, Ham70, Haw78, HK77, JTS75, Lee79, Saw70, Smi76, SM79, Wu73, ZA75]. Letters [BJ70a, GMD^{+}75, GH75, HOC^{+}77, HJ73, JEBE74, Kru70, KBW72, LSP76, MP72, SBV79, VK76]. Level [Ano73f, Dic74, Luc78, MB78a, MB78b, Sac72, SS75, Web71, LH77]. Levi [Bon76b]. Levy [K.71a]. Lewis [Ron72, Fre73c, J.79l]. Lieberman [Web73]. Life [Bai74b, BS71, Bar72, BL71, Bli72, BK70b, Dan70, Dun74, FP77, Gru73a, IR77, J.79p, KS71, Kam77, KA71, Law71, Law73, Man72, Moe74, NL75, Nel72a, NK76, NM78, NS79, Sch74a, SLK79, Sha78, Sin71, Sin73, Smi77, Bha73]. Life-Test [Man72]. Life-Testing [Bai74b, KA71]. Lifelength [J.75f]. Lifetime [CF77]. Lifetimes [KKM79]. Light [Odo77]. Likelihood [AA74, AR79, Ano73k, AR73, BR73, BWL78, Coh65, CS76, DAH73, Har73a, HH73, HP71, JS76, Lem75, Mat78, MN77, MW75b, SP74, SP77, SB72a, TBA70]. Limit [Hah75, Sen74]. Limits [BS79, EB78c, EB79, Fau70, FD70, FKRO71, MN77, RSO72, TBA70, TC75a, TJ79]. Lindman [Sne76]. Line [For72, WB74]. Linear [ABA72, And74, Ano72g, Ano74f, Bai74b, BD72a, BL74b, Boa79b, Bro79, Coo77, Coo72, D’A71a, D’A71b, Far75, Fuc78, Gue71b, GPS73, Hal70, Har73b, Hel74, HM77, HSL76, J.75d, J.75a, J.76g, J.77b, JB76, Joh71a, Kal71, Krz77, LaM78, LS71a, Leo75, LW71, Lun75, Man71b, MF73, Mar70, MK70, MW73, McL74b, MW75b, MHL74, MHL77a, MHL77b, NW77, NW79, NH72, NH73, OM72, Pas79, Pre75, Rah76, Ron72, SG71, SS71a, Sha78, Sie75, SM74, Sor73, Spr76, SB72b, Tar76, TW72b, TMB73, Wig76, Zel72, GDF78, Kev75, For73, Hul73, Lin75, Ran70a]. Link [Lac74a]. Linnik [Sen75]. Linstone [J.76i]. Living [Ben76]. Location [DAH73, Har73a, J.77p, Law75b, RRH73, LR78]. Locations [J.77e]. Locks [Sie75]. Loftsgaarden [Hac78]. Log [Coh76, DA73, NS79, SS71a]. Log-Linear [SS71a]. Log-Normal [Coh76, DA73]. Logarithmic [WY78]. Logistic [AR79, BMH76, CC71]. Lognormal [Jon71, Lan72, NK76, Sin72]. Lohnes [O’F73]. London [Ens78]. Loop [BM74, BM76, Ell76, PW77]. Loren [Nel74]. Lot [BS72, Eld74]. Lothar [Bed74]. Low [Law77b, MH77]. Lower [GZ79, Sar71]. Lucien [Fre73b]. Lukacs [Lan73]. Lumb [K.73e]. Lung [Ens78]. Luther [JH73b]. Lynwood [Dun77a].

M [Add72, And75, Boa79a, Coo72, Dud79, Dun72, Ger75, Gib74, Hea78, How77, Ing78b, J.76g, J.76f, J.76i, J.78h, J.79k, J.79v, J.79m, Kal75, Kha74, Lac71a, MRR78, Odo77, Ran70c, Rog70a, Sen75, Sim76, Urq73, Wil79a]. M-Estimator [MRR78]. Machine [WB75]. Machol [J.78e, J.77i]. Madan [K.71c]. Main [CDR79, DR77, Dic74, Nai74]. Main-Effect [CDR79, DR77]. Maintain [Dun78a]. Maintainability [Ziz78]. Maintenance [RR76]. Majorization [Low70]. Making [J.78h, Kin78, Win71]. Malik [Sch74b]. Management [Gua71, J.78p, Ziz78, Dun77b, J.78e]. Mantel [AK71]. Manual [Ziz78]. Manufactured [Kao72]. Many [Ros75b, Ros77]. Maps [J.79l]. Marc [Fre73a, J.77a]. Marcel [J.79b]. Mardia [J.79u, Wat73]. Marginal [Par70]. Margolin [Ano70f]. Mario [J.75d]. Market [Nay72]. Markets [Kno71]. Markov [Cra74, Sys75, AvdL70, Bri74, JK74, WB71, Doo72, Sys76]. Marquardt [SS71b]. Marvin [Mye74]. Mask [Luc73]. Massey [Wei71]. Material [BF72, Lev79]. Mathai [J.79v, Sim76]. Mathematical [Ber73, Bow77, Boy74, Cha75a, Cor71, Eri72, Fre73b, Hac79, Hal77b, Har72, Ing78a, J.75g, J.76h, J.77k, J.78o, Jon76b, Lan73, Sen75, Bis72b, Lac72a]. Mathematics [Cha70a, Eas72, Ham75, K.71b, Spr72, K.74b, Leh71, vdV75]. Mathematische [Ens77]. Matrices [And73, Ban73, GZ79, GC74, Urq73, Zys70, Gow72]. Matrix [CGO07, Fra77, Goo70, Haw75, Ray70, Ron77]. Matter [Ano70b, Ano70c, Ano70d, Ano70e, Ano70g, Ano70h, Ano70i, Ano70j, Ano71b, Ano71c, Ano71d, Ano71e, Ano71h, Ano71i, Ano71j, Ano71k, Ano72a, Ano72b, Ano72c, Ano72d, Ano72h, Ano72i, Ano72j, Ano72k, Ano73b, Ano73c, Ano73d, Ano73e, Ano73l, Ano73m, Ano73n, Ano73o, Ano74a, Ano74b, Ano74c, Ano74d, Ano74h, Ano74i, Ano74j, Ano74k, Ano75a, Ano75b, Ano75c, Ano75d, Ano75g, Ano75h, Ano75i, Ano75j, Ano76a, Ano76b, Ano76c, Ano76d, Ano76g, Ano76h, Ano76i, Ano76j, Ano77b, Ano77c, Ano77d, Ano77e, Ano77i, Ano77j, Ano77k, Ano77l, Ano78a, Ano78b, Ano78c, Ano78d, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano79a, Ano79b, Ano79c, Ano79d, Ano79h, Ano79i, Ano79j, Ano79k]. Matters [Ano72e, Ano73h, BD71]. Maurice [Fie73, Ger75]. Maximize [Ano71a, Dyk71, Eva79]. Maximum [AA74, AR79, Ano73k, AR73, BR73, BWL78, Coh65, CS76, Gri72, HH73, HP71, JS76, Lem74, Lem75, Mat78, MN77, SP74, SP77, SB72a, SU79, TBA70]. Maximum-Likelihood [Ano73k]. May [Wio70]. McFadden [Bis72a]. McNeil [J.78d]. Mead [ON75]. Mean [All71, Bad70, Bor72, BH73, CF76, Far75, FKRO71, GW71, Jos72, Kus79, Low74, McL74a, McL74b, ML75, PK79, Vin76, Wan71, HK77, LH77]. Meaning [Mal74, Mar74, BT74]. Meaningful [WW78]. Means [BF74, CHH^{+}71, Fed75, Hah75, Hah77, Lan72, SU79, Ury76, WC70]. Measure [Igl73]. Measurement [BR75, Gru73b, Har75, J.79t]. Measurements [TB75]. Measures [WW78, Sne70]. Measuring [Gru73b, HN70, Han75]. Median [Wol77b]. Medical [Fle76, Lac72b, Ros75a]. Meeting [Ano70a]. Meltzer [Hel71b, Jac80]. Men [Ens78]. Mensing [Gle77]. Metagames [Lut73]. Meters [Joi77]. Method [All74, And70, And74, Ang74, Aro76, CR79, CBI79, GHW79, Gue77, Hoo79, J.76i, Lin79, OB76, SH79, See70, TW70, Yue78, ZNL79]. Methoden [JH73b]. Methodology [Dud79, Man73]. Methods [Bis72b, Boa79a, BCK61, Con74, Dav77, Dav79b, Dun72, Ehr77, FM71, Goo71, Gri70, HR75, J.76d, J.76e, J.77b, J.79q, J.79b, Jae70, Lac72b, Lan72, Leo70, MBC79, Mer79, Mih75, Mih77, NH72, Pat73, Ran70a, RH78, Sal75b, San72, Sne77, Sto76, Weg72, Zel72, BCK78, And75, Cor71, For71, Gib74, J.78i, J.78f]. Meyer [Joh71b]. Michael [Car77a, Sie75]. Micro [Car77b]. Micro-Organisms [Car77b]. Microbiological [Car77b]. Microscale [Ano73j, Mar73]. MIL [BS79, Koy78, Koy79]. MIL-STD [BS79]. MIL-STD-105D [Koy78, Koy79]. Milestones [CGO07]. Milton [K.71e]. Minimal [AT79]. Minimization [ON75]. Minimizing [For72]. Minimum [BD74, BR73, Dic74, Far75, Hal79, Nád71, NW77, NW79, Sch74a, SA79, Sch76, SLK79, SS78, Tho73, Vin76, WK77, HK77]. Minimum-Point [BD74]. MINITAB [J.78g]. Misclassification [Lac74b, Lac79, McL74a, Sor71, Sor72a, Sor72b, Sor73]. Misclassified [Lac74b, McL72, Ten72]. Missing [BDH70, HG73, Joh79, Pre71, Rub76, Wat70a]. Misuse [Kin79b]. Mitra [Ban73]. MIVQUE [SS78]. Mixed [Beh72, CR79, CS76, DS79, Haw72, HH73, SWD74]. Mixing [BR75, Cho79, Luk73]. Mixture [Ano76e, Beh70, BF72, Cha79, Cor75, CO75, Cor77, CK79, GK77b, Har79, Joh70c, Joh70b, MS74, Par78b, Ray72, Sch71a, Sne73, SM74, Sne75a, Sne75b]. Mixtures [Cor73, Cor79, DS77a, DS77b, Gri70, LC70, SN77, SM76, GK77a]. Model [Ano72g, Ano73j, Ant77, BDH70, Bri74, CF77, CS76, Cor77, DS77b, GLV74, Gue71b, HBA71, Hah72, HH73, HP71, Hil78, HJF78, HM73, Leo75, Man71a, Mar73, MW78, MHC77, MPB70, MA78, Nai74, OM72, Par78b, Pas79, Rah76, RR76, SWD70, SB72b, Tar76, TA75, TC73, WK71, WB74, WG78, Wio70, GDF78, J.75d]. Model-Discrimination [MPB70]. Modeling [LS71b, LSM72, PW77, Sch75]. Modelling [MPBD79, J.79m]. Models [Ano76e, BK70a, BG78, BL73, Cle71, DS77a, DAH73, Dun74, Fuc78, GK77b, Goo71, Gri79, GPS73, Har73a, Har73b, Hel74, HM77, K.71d, Lac74b, LR76, Lun75, MS74, MW73, Mer72, Par78a, Par78b, Pre75, PDB77, San72, San73, See73, SM74, Sne75b, Sne77, Spr76, Spr78, SWD74, TW72a, Wol78, Boa73, Boa79b, Hul73, J.75a, J.76g, Ron72, Smi77]. Moderate [Gut73b, MF77]. Moderate-Size [MF77]. Modern [Dun77b, Spr78]. Modes [Beh70]. Modification [Dav79a, Mie74, SOE75]. Modifications [Zah75b]. Modified [DM78, KM79, Koy78, Lin79, Luc73]. Modular [Bod70]. Modulus [SU79]. Molenaar [Hel71a]. Moments [Cha70b, Gro71b, Jae74]. Monotone [Sen74]. Monotonicity [Gro71b]. Monte [Rel70]. Montgomery [Dun77a]. Moore [Hub78]. Morris [Kin78, Ran70b, Win71]. Morrison [J.78e]. Mortality [MS73]. Most [Ala71, Bli72, EB77b]. Mosteller [Lac74a, Qua75, Odo77]. MOSUMS [BH78]. Motion [CF76, Nád73, PK76, Ser73, SS74]. Motions [HM73]. Moving [CK75, RH78, Ste72b, SWD74, Wic72]. MR [Kit81].

MTBF [BS79]. Mullen [Sch74b]. Multi [Coh75, HP71, Sac72]. Multi-Censored [Coh75]. Multi-Level [Sac72]. Multi-Risk [HP71]. Multicollinearity [WW78]. Multicomponent [SCSB76]. Multidimensional [CR79, Goo71]. Multifunctional [Sac76]. Multinomial [AA74, Ang74, BK70b, OL77, Ten72]. Multinormal [HS72]. Multiple [Ait74, And74, Bea70, FH79, Goo71, Gor76, HG73, Haw72, LSYF79, SP74, SU79, Ury76, VM79, WB71, Cle72a, J.77g, J.78h]. Multiplicative [SOE75, YC78]. Multiresponse [BDH70, BHME73, Gri73, MPBD79]. Multivariate [BMP77, Cha79, Dun72, Fre74a, Gaf78, GG70a, Gut73a, HB71, J.78f, J.79i, McK79, Mil72b, Odo77, Pat73, Reg76, Ron77, Ger75, J.75c, J.76d, J.76e, J.77h, K.74c, Kha74, O’F73]. Murdoch [J.78p]. Muse [Ano79e]. Mutually [BCK61, Hed71, BCK78]. Myers [Rus73, Fed74a, Man73]. Myint [Cha75a]. Myint-U [Cha75a].

N [Bis72b, Che71, Dav70, Geo72, Gri73, J.75a, J.77n, J.77j, J.79t, J.79h, K.71a, Sim76, Smi70a]. Nadler [Buc73]. Namboodiri [J.79t]. Narayan [Hei79]. Narula [Ano80b]. Nash [J.77b]. Naslund [Dun77b]. Natural [Ens78]. Naylor [Mer72]. Near [AT79]. Nearly [Har74]. Negative [BG74, Bin74a, Gru71c, Gue72, MC72, Nel75, Pie73, Ron77, GH75]. Nelder [K.71e, ON75]. Nerlove [Fre73a]. Nested [Ede74, GG70b, OL77, TA75]. Net [Bin74b]. Network [Gor76, Mar75]. Neutral [Sta79]. Neuts [J.79b]. Newton [JS76]. Neyman [Fre73b]. Nigel [Lut73]. No [Bor72, Har73a, LvNN73, Mal74, Mar74, Par72, Zel72, GH75, HK77, Kit81, Smi76]. Noether [McC72, Mer77]. Noise [Shu71]. Nomogram [GW71]. Non [BC71, BH77, GPS73, HG73, J.77b, Joi77, Krz77, Lac74b, Man71a, Pep73, RSO72, Ron77, Tak72, TA71, Weg74, WB75]. Non-Additive [Man71a]. Non-Interactive [BH77]. Non-Iterative [HG73]. Non-Linear [GPS73, J.77b]. Non-Negative [Ron77]. Non-Normal [Tak72]. Non-Normality [RSO72, TA71, WB75]. Non-Optimal [Krz77]. Non-Orthogonal [BC71]. Non-Random [Lac74b, Pep73]. Non-Standard [Joi77]. Noncentral [MCL76, KP79]. Noncentral-F [MCL76]. Nonlinear [BD72a, BL74b, BDH70, Box70b, BL73, DV78, Dun78b, LS71a, LSM72, LvNN73, Mar70, NV73a, NV73b, RJ78, Sac76, Sha71, SS71b, Cha75b]. Nonlinearity [LaB77]. Nonorthogonal [HK70a, HK70b]. Nonparametric [Ano80c, BR79, Bha72, Cho73, McD79, Mer77, Weg72, Dun72, Gib74, Joh73c, K.71c]. Nonparametrics [Qua75, Mer79]. Nonrandomized [MDM77]. Nonstationary [HM73]. Norbert [Kab74, Rus73]. Normal [Aus73, Beh70, BR73, CC71, Cha79, Coh76, Dow73, DA73, FD70, FM77, GG70a, GW71, Gri70, GS71, Joh70c, LNN61, MD71, McM71, Mil72b, MM73, Nád71, NK76, NS79, OlHS77, RSO72, Ray72, Reg76, Riz71, Sor73, Tak72, WK77, Zah75a, Zah75b, LH77]. Normality [Ano75e, Fil75, Kow70, RSO72, TK72, TA71, WB75]. Normalizing [WY78]. Norman [And70, Cru78b, Fre74a, Spr78]. Note [AM79, Ano73i, AK71, Bai74a, Bea70, Bis78, Con70, Cro71, DR77, DL70, DY70, Eva71, Fau70, Fra77, Hal77a, Haw72, HK75, JK71, Kap72, Ken71, Lac79, MH72, MW75a, Ng76, Sch70, SS71a, Sne71, Tho75, Ton74, Ton75, JTS75]. Notes [Coo72]. Nuclear [BKRS78, HS71, J.78j]. Nuisance [EB77b]. Null [LF79]. Number [Dic74, Gru73a, Gue71a, Wio70, Wol78]. Numbers [Ben76]. Numerical [AB74, CR79, Dil71, Dun78a, Eva71, Eva72, Lin74, RH78].

O [Hac78, HH73, Ron72, Ros75a, Sie75, Sne74]. Objectives [J.78b]. Objects [Ens76]. Observable [Nád70a]. Observation [Coo77, Joh70c, Joh70b, KS71, Lik74]. Observations [Ano73g, FM77, GB72, Joh70c, Joh70b, Man70a, Ray72, Sin71, TS79, BN72, HP73, Fre75, J.78f]. Observed [FB75, Har74]. Obstructive [Ens78]. Obtain [Gre70a]. Obtained [Sin71]. Obtaining [CK79, Gue77, Jae70, MF73]. OC [Kow71]. O’Carroll [J.79m]. Occurrence [Gre70b, Sap73]. Odeh [J.76g, J.78l, J.77k, J.78o]. Odell [Gow72, Ron72]. Odoroff [Ano77a, Ano79f]. Olkin [Dud79, Sid75]. O’Muircheartaigh [J.77m]. One [Bai73, FM77, GLV74, GPU76, HP73, Hoc72, JD78, KB75, Kus79, Lad77, MD71, McM71, MM73, Rah76, SS75, SD77, Wes71, Boa79a]. One-Four [Boa79a]. One-Sided [FM77, SS75, Wes71, GPU76, HP73]. One-Way [Kus79]. Only [CHH^{+}71, Nád70a]. Operates [DY70]. Operating [BM76, Ell76, PW77, SLK79]. Operation [And70, Wan71]. Operational [J.79n]. Operations [J.77a]. Opinion [Gle78]. Optimal [Ano77h, BN72, Bis72b, BH77, EB78a, Gor76, Gre70a, Har74, J.79d, Jon76a, Krz77, Luc78, Mac76, MW78, MCM75, Mit74a, Mit74b, MB78a, MB78b, Sim73, Sor72a, SC71, Gor76, J.77i, Ran70b]. Optimality [NGM70, SD75]. Optimally [Kam72]. Optimization [Bro79, HS73]. Optimum [BS72, CF76, HS72, Law77b, Luc74, MG74, MH77, NK76, NM78, PK79, Sch71b, GK77a]. Orbital [Bus71]. Order [Ber72a, BMH76, BD74, Gue77, Hed71, Hok74, JC76, Kal75, Kam72, Kam74, Mit74b, OM72, Qua75, SWD70, SWD74, TW72b, Weg74, Gru71a]. Ordered [Lik74, Man70a, Ode71]. Ordering [Dud79, Gar71]. Orders [Dic74]. Ordinary [MHL77a, Sac75]. Organisms [Car77b]. Origin [Joh70c, Joh70b, Ste78]. Orkin [Car77a]. Orthogonal [BC71, BCK61, CDR79, DR77, Hed71, BCK78]. Osborne [Lac71a]. Ostle [Gle77]. Other [J.79v]. Otnes [Nel74]. Outlier [Ano74f, JD78, Jos72, KS71, Ros77, TMB73]. Outliers [Ano79g, GW75, GS71, Gut73b, JD78, JMM78, Kit79, LF79, Lun75, MD71, McM71, MM73, Pre75, Ros75b, Ste72a, TM72, GDF78, Kit81]. Outliners [Gut73a]. Outlying [Ano73g, GB72]. Output [MBC79]. Over-Stress [Law77b, Man72, MH77]. Owen [J.77k, J.78l, J.78o, Har72, J.75g, J.76h, J.78k, Tan70].

P [Add79, And70, Dav70, Dun72, Fle76, Gow72, GH75, Hea78, Hil74, J.75b, J.77n, J.77i, J.77h, J.78e, J.78q, J.78a, J.79g, J.79d, J.79u, J.79f, Jon76b, K.71d, K.74c, Kle77, Lac72b, Law75b, O’F73, Qua71, Ram73, Ron72, Sax72, Sim76, Sne70, Spr79, Ton72, Zac74]. Pacemakers [BKRS78]. Packages [Hub78]. Packing [GLV74]. Pair [Nád71]. Pairwise [Ury76]. Paliouras [Cas76]. Panne [For73]. Paper [Fed75, MK70]. Papermaking [TW72a]. Papers [Ano70a, Fie73, HM71, J.77l, K.73d, K.73b]. Paradoxes [Lut73]. Parallel [Bry71, Wan71]. Parameter [AvdL70, AJ77, BE73, Bec74, Box70a, BM76, Cha75b, Coh75, Coh76, CN77, CM71, DG71, Dun78c, Dye73, Eas76, Ell76, EB74, EB75, EB77b, EB78c, GHW79, Gru71c, Kam72, Kam73, Law77a, Lem75, Lik74, Low74, Man72, MF75a, Man75, Mat78, MC72, O’B73, Pas79, Pie73, RRH73, Rou73, TW72b, GPU76]. Parameterizations [Sin73]. Parameters [Ano73k, BD72a, BL74b, BL73, D’A71a, D’A71b, DAH73, Eng75, EB78a, Fal70, Fin76, Har73a, Has72, HH74, Jae70, Kam74, Kap75, LS71a, Man70a, Man71b, MF73, MF75b, OB76, Par78b, Pat76b, RS71, Rah76, SG71, SR75, Wio70, LR78]. Parametric [Rog70a]. Parent [YM74]. Parking [GLV74]. Part [Zel72, NH72, NH73]. Partial [LM72, Cha75a]. Partially [Har73b]. Particles [GLV74]. Partitioned [Far78]. Partitioning [CR79]. Passage [CF76, Nád73, PK76, SS74]. Path [Bur73]. Patil [K.71d]. Patterns [BD72b, Cor77, Dan78]. Paul [Ber73, Eri72, Gil74, Joh71b, Jon73, Kin79b, Odo77]. Payne [J.77m]. Pearson [Sne74, Har71]. Penalty [Gre70a]. Percentage [Ano73g, GB72, Gue77, Har71, IM70, KB75, Krz79, Ros77, Som70, YM74, TC72]. Percentages [RSO72]. Percentiles [CC71]. Performance [BJ75, Dub73, Jae74, JB74, KKM79, Kow70, Krz77, Luc76]. Permuted [J.77f, J.77q, K.71a]. Peter [Add72, Fre73c, J.77c, J.79l, K.73e, K.74b]. Peto [Ens78]. Phase [BC79b]. Phenomenon [Bry71]. Phillip [Fin75, Str75]. Physical [Cru78b, Bis72a]. Physics [Cha75a]. Piersol [Har75]. Pieters [Lac74a]. Pietro [Qua71]. Pisani [Kin79a]. Pitfalls [BK74]. Pitman [MC72, J.77l]. Pitman-Type [MC72]. Plackett [Fed77, Wat70a]. Plan [Joh70a, KO77, SP72]. Planning [Dan70]. Plans [Ano72g, BLB73, BS79, CDR79, DR77, Dic74, Gue71a, Gue71b, Hal77a, Hal79, Lad77, MCM75, MH78, PS70, Sac72, SP74, SP77, Sha79, Tak72, VM79, WB71]. Plot [Ano75e, Fil75, Zah75a, Zah75b]. Plots [LaB77, LM72]. Plotting [Nel72b]. Pocket [J.78l]. Point [BD74, EB74, Fre73c, MF75b, WK71, And75, J.75d]. Points [Ano73g, GB72, Gue77, Har71, IM70, Joh79, KB75, Krz79, Man75, Ros77, Som70, YM74, TC72]. Poisson [Hel71a, AT73, Coh72, CM71, CJ73, GH70, Ham72a, Jae70, Man75, Nel75, Pat73, Pat76b]. Policies [Sch71b, TC75b]. Policy [BS72, Jac80]. Policymaking [Jac80]. Political [Lut73]. Pollution [Ano73j, Boy74, CG76, Mar73, MS73, Sin72]. Poly [Kle75]. Poly-Factor [Kle75]. Polynomial [Cor75, Gaf78, LW71, Par78a]. Polynomials [Mal74, Mar74, BT74]. Population [Beh72, BSH71, FM77, HG78, Joh70c, Joh70b, McD76, RRH73, SR75, Zac70]. Populations [Ano73k, Dud79, Gro72, Joh70c, Joh70b, Ray72, Reg76, SS76, vN79]. Port [Gil74, Sax72, Ton72]. Portable [BK74, Whe74, Whe75]. Positive [Bin74a, CF76]. Possible [CBI79, Fur71, MT72, Sta79]. Postulated [SLM74]. Potential [BK74, Sys75]. Power [BJN71, BK74, BH74, Bri73b, DW72, Dub73, FR72, For72, GC74, HK75, KARR75, LH74, Mal74, Mar74, PK76, PK79, Whe74, Whe75, Yue78, BT74]. Powered [BKRS78]. Powerful [EB77b]. Practical [NGM70, Bha72, Sie75]. Practice [Ben76, J.79m, Odo77, Pre77]. Practices [Eas72]. Pranab [Kee75]. Pratt [Boy74]. Pre [Wil79a, Rah76]. Pre-Test [Wil79a, Rah76]. Precision [Ben76, Dan70, Gru73b, Jae76]. Prediction [All71, All74, AM79, BSH71, Cle71, CK79, EB78b, EB79, FM77, GM79, Hah72, Hah75, Hah77, HM76b, KN74, Kam77, Law71, Law77a, Lik74, MF77, NW77, SS71a, HP73, Boa79b]. Predictive [Dun74]. Predictor [LvNN73, NV73a, NV73b, Rub76]. Prekopa [K.74a]. Preliminary [McD75, Obe75]. Premium [Gut73b]. Prescribed [Dan70]. Presence [GR72, JMM78, KS71]. Present [Jos72]. Presented [Ano70a]. Preservation [Luk73]. Previous [Nel72a]. Price [Hah79]. Prices [Kno71]. Primer [Gie74]. Principal [JH73a, JM79, JH79, Krz79, Wol78]. Principles [Car77b, Cor71, Han73]. Pringle [Urq73]. Prior [ABA72, Ano72g, Chi74, Gro72, Gue71b, PS70, Rah76]. Probabilities [Dun71, Sor72a, Sor72b]. Probability [Ano75e, Bad70, Bor72, Bur73, Fil75, Fre73b, GH70, Hac78, Hac79, Hei79, Hol75, J.77d, J.77f, J.77e, J.77l, J.77o, J.77q, J.77p, JR72, Joh71b, K.73e, Kap72, LaB77, Lac71b, Law77b, MH77, MH78, Moe72, RTDM79, Rog70d, Sam73, ST72, Sch74b, Sid75, Smi77, Sne70, Sor71, Sor73, Spe71, Spr78, Sta70, Sym71, Ton72, Weg72, Woo74b, Zac74, Bis72a, Har70, J.78k, J.79b, Mil72a, Rog70c, Zie78, Bow77, Fed74a, Gol74, Hei78, Hog74, How77, J.76f, Kup71, Lan73, Ram73, Rog70b, Wei71, Zac74]. Probable [Ala71]. Probe [JK74]. Probit [Gou73]. Problem [AT73, AF76, Gri72, HN70, Haw72, Hel74, Kal71, KA71, Law71, Lem74, MW75b, MRR78, Ray72, Ros75c, Sam78, Shu72, Wat74, LH74]. Problems [And75, BC71, BHME73, EB79, Gue74, HK70a, HK70b, McL74b, Riz71, Sen75, Sin72, Whi70, Boy74, J.77b, Hun74, J.79o]. Procedure [BR79, BKRS78, Bin74b, Bis78, Bli72, Cha72, GG70a, GR72, KKM79, MB74, O’B73, ON75, RRH73, Ros77, SR75, SS71b]. Procedures [Ano73a, Ano80c, Bea70, Ber78, BD78, CF76, CG76, EB77a, EB78a, FH79, Gle78, Goo71, Gut73b, Hil78, Igl73, JM79, LSYF79, LO79, Man70b, McD79, McK76, PW72, Pre71, Reg76, SS76, Ury76, Eas72, Har75]. Process [Ano77h, BR79, Coh72, CM71, DS79, Dun78a, EB78b, Fin76, Har79, HD70, LL78, Mac76, TW72a]. Processes [AvdL70, AT73, BGH73, Fre73c, Rog70b, Gil74, Sys78]. Processing [Gle78]. Product [DS79, Hah77, Kao72, OB76, YC71, Ziz78]. Production [Bla77, DS79, Sha79]. Productivity [And70]. Products [Hay73]. Progamming [AA74]. Program [Wat70b]. Programmed [Boa79a]. Programming [AK71, For73, Ran70a, Sie75, Sys75, Coo72, Gro77, Lin75, Sha71]. Programs [Hea78]. Progress [Ano76f, J75e]. Progressive [Man71b]. Progressively [Coh76, CN77, TW72b]. Proper [BJ70b]. Properties [Bry74, Gro71b, Joh70a, Low76, Spr76]. Proportion [Wio70, Zac70]. Proportional [Nád70b]. Proportionality [Man75]. Proportions [AB71, BR75, BM78, Cho79, MDM77, Con74]. Proschan [J.75f, Smi77]. Protection [Gut73b]. PrY [Ury72]. Psychological [Ger75]. Publications [Ano75k, Ano76k, Ano76l, Ano76m, Ano76n, Ano77m, Ano77n, Ano77o, Ano77p, Ano78k, Ano78l, Ano78m, Ano78n, Ano79l, Ano79m, Ano79n, Ano79o]. Pure [Kee75]. Puri [Dun72, K.71c]. Purpose [ST73]. Purves [Kin79a].

Quadratic [Ano76e, AK71, GK77b, Lac79, Luc76, Roq76, Sne75b, SS78, TC73, Yue78]. Quadruple [Pat76b]. Qualitative [J.79c]. Qualitative/Categorical [J.79c]. Qualité [Gua71]. Quality [BH78, DS79, Gua71, Pat73, ST73, Tsi78, Ziz78, Kab74, Woo75]. Quantile [MF77]. Quantiles [Dye73, Has72, Kam73, SA79]. Quantitative [Gri73, Sto76]. Quartic [Mat78]. Quasilinearized [KME71]. Quasimidranges [LC70]. Query [Bad70, Bor72, Leo70, Wat70a].

R [And70, Bow78, Bri73a, Fed74a, Fed77, GW71, Gor76, Gri73, Hog74, Hul73, Ing75, J.75b, J.75f, J.76g, J.76d, J.76e, J.77j, J.77k, J.77h, J.78l, J.78e, J.78f, J.78d, J.78m, J.78j, J.78b, J.78a, J.78o, J.79g, J.79r, J.79o, J.79v, J.79d, J.79f, J.79m, K.73b, K.74c, Kal75, Kin79b, Kup71, Lac71b, Lac71a, Lac74a, Law75b, O’F73, Sal75a, Sen75, Smi76, Sne76, Urq73, Weg76]. R. [MK70, J.79p]. Rade [K.74b]. Radhakrishna [Ban73]. Radium [PBT76]. Raiffa [J.78b]. Rainville [Fin75, Str75]. Raj [Sen70]. Ralph [Rog70b]. Ralston [J.79q]. Ramachandran [Sen75]. Random [Bag70, BL71, BK70b, Bri73a, Cha70b, GLV74, HW75, J.78q, Kus79, Lac74b, Nád70a, Nai74, Pep73, Rog70b, TA75, WY78, LH77, Har75, K.71d, Spr79]. Randomization [TW70, You72b]. Range [Fed75, FKRO71, GKS75, Nea70, Smi76]. Ranges [McD76]. Rank [Ano84, BJN71, GZ79, IC79, Kra76, Mie74]. Ranking [AT73, BR79, Bis78, SR75]. Ranks [HM77, Mer79]. Rao [Ban73, Sen75]. Raphson [JS76]. Rasch [Ens77]. Rate [Bai74b, Dan70, FR72, HH77, LR76, LE73, McL74b, Sha78]. Rates [Ano80c, Bry71, Con74, McD79]. Rathie [Sim76]. Ratio [BL71, Dav79a, DAH73, Har73a, Jae70, ST72]. Rationality [Lut73]. Ratios [BK70b]. Ratti [Spr72]. Ray [Sau73]. Raymond [Jon73, Kno71, Man73]. Rayner [Urq73]. Re [Sin73]. Re-Parameterizations [Sin73]. Reaction [SLM74]. Reactor [HS71]. Readings [Fre73a, K.73c, Qua71]. Real [Boa79a]. Reasoning [Hac78]. Received [Ano75k, Ano76k, Ano76l, Ano76m, Ano76n, Ano77m, Ano77n, Ano77o, Ano77p, Ano78k, Ano78l, Ano78m, Ano78n, Ano79l, Ano79m, Ano79n, Ano79o]. Reciprocals [BK70b]. Recognition [Dol70]. Recorders [Nea70]. Recursive [LF79]. Reduction [Gen77, Par78b, Rel70, Ens76]. Redundancy [Gre70a]. Redundant [TC75a]. Reflected [Coh73]. Regime [Rif71, WB74]. Regions [Dun78b, Luc76, Obe77, LR78]. Regression [Ait74, ABA72, And74, Ano74f, Ano84, BC71, Bea70, BL74b, Ber78, Bro77, Coo77, Cro78, DM77, Dun78b, Far75, Far78, For72, Fra77, Gaf78, Gor76, Gre71, GWM76, Hah72, Hal70, Haw75, Hem75, Hil78, HJF78, Hoc72, HSL76, HK70a, HK70b, HK75, HR75, Hoo79, IC79, J.75d, KME71, LH70, LM72, LS71a, LSM72, Mar70, MK70, MS73, McK79, MHL74, MHL77a, MHL77b, NW77, NW79, NH72, NH73, Obe77, Par70, Par78a, PW72, Rah76, Sac76, SH79, SS71a, Sne77, SD75, SD77, TMB73, WGM74, WG79, Wio70, YF76, Zel72, DV79, Ell78, McC78a, For71, Fre75, Hah79]. Regressions [BC79b, DH74, Fur71, FW74, MT72, Rub76, SS71b]. Regular [Ber72b, Gue72, TS79, GH75]. Reinhardt [Hac78]. Rejecting [Ste72a]. Related [Ano72e, Ano73h, BD71, For73, J.79u, JS76, Wol77a]. Relating [MS73]. Relations [Haw75]. Relationship [All74, McL74b, NH72, NH73, Zel72]. Relationships [CH70, JH73a]. Relative [LR76, NW77]. Relevant [YC71]. Reliability [Bha73, BGH73, Bod70, Bri74, BWL78, CB75, CH70, DY70, EB78c, FB75, Gro71a, Gru71b, Gru71c, HBA71, Hei76, Her70, HM76b, J.78j, MG74, Pie73, Rea71, Sar71, Smi77, TBA70, TC75a, Var70, WK71, WG78, Whi70, Ziz78, J.77n, Kab74, Eas72, J.75f, J.77j]. Remark [Rea71]. Remarks [AvdL70, Nel75, SB72a]. REML [CS76]. Renewal [BGH73, HD70]. Renyi [Rog70c]. Reordering [BS72]. repair [Bha73]. Repairing [HM72]. Replacement [CH74, IS76, RR76, Sch71b, TC75b]. Replacements [IR77]. Replication [Gre71]. Report [Ale72, Ano74g, Com70, Com71]. Representation [Lac71a]. Required [SdV73]. Requiring [Dic74]. Research [Cha70a, Gle78, J.79n, Ger75, Gle77, J.77a, Lac72b]. Residual [LM72, MT72]. Residuals [BD72b, Dan78, DL70, GW75, JM79, Nel73, Sne71, Woo73]. Resolution [AT79, LS71b, SC71, SLM74]. Response [CP75, CO75, HM72, HS73, Luc76, MB78b, MC73, ML75, Roq76, Sne72b, Sne75a, Nay72, Man73]. Responses [Sta79]. Restricted [AF76, CS76, Joh70a, SN77, Wat74]. Restrictions [Kal75]. Results [Bai73, BB79, Bis79, DM78, Dun78a, EB73, EB74, Hol75, LL78, McL72, Nel72a, Weg74, YC71]. Review [Add72, Add79, And70, And75, Ano71f, Ano74e, Bak74, Ban73, Bed74, Bel70, Ben76, Ber72b, Ber73, Bha72, Bis72a, Bis72b, Boa79b, Bon76a, Bon76b, Bon77, Bow77, Bow78, Boy74, Bri73a, Buc73, Cam77, Car77a, Car77b, Car74, Cas76, Cha75a, Cha75b, Che71, Cle72a, Con74, Coo72, Cor71, Cor73, Cra74, Cro78, Cru78a, Cru78b, Dal77, Dav70, Dav77, Dil71, Doo72, Dra79, Dud79, Dun72, Dun77a, Dun77b, Dye74, Dyk76, Eas72, Ehr77, Ens76, Ens77, Ens78, Eri72, Fed74a, Fed77, Fie73, Fin75, Fle76, For71, For73, FW72, Fre73a, Fre73b, Fre73c, Fre74b, Fre74a, Fre75, Gen77, Geo72, Ger75, Gib74, Gie74, Gil74, Gle78, Gle77, Gol74, Goo75, Goo76, Gor76, Gou73, Gow72, Gra72, Gri73, Gro77, Gru71a, Gua71, Hac78, Hac79, Hah79, Hal77b]. Review [Ham75, Han73, Har77, Har70, Har75, Har72, Hay73, Hea78, Hei76, Hei78, Hei79, Hel71a, Hel71b, Hil74, Hil78, Hog74, Hol73, How77, Hub78, Hul73, Hun70, Hun74, Ing75, Ing78a, Ing78b, J.75d, J.75c, J.75g, J.75b, J.75h, J75e, J.75a, J.75f, J.76c, J.76b, J.76g, J.76d, J.76e, J.76f, J.76a, J.76i, J.76h, J.77b, J.77d, J.77f, J.77e, J.77l, J.77c, J.77n, J.77j, J.77g, J.77m, J.77i, J.77k, J.77h, J.77a, J.77o, J.77q, J.77p, J.78n, J.78h, J.78l, J.78k, J.78e, J.78f, J.78g, J.78d, J.78p, J.78q, J.78i, J.78m, J.78j, J.78c, J.78b, J.78a, J.78o, J.79q, J.79i, J.79p, J.79g, J.79e, J.79j, J.79c, J.79l, J.79r, J.79k, J.79o, J.79t, J.79b, J.79v, J.79d, J.79n, J.79a, J.79s, J.79u, J.79f, J.79h, J.79m, Jac80, Jer70, Joh75, Joh71b, JH73b, Joh73c, Jon73, Jon76b, K.71a, K.71d, K.71b, K.71c, K.71e, K.73d, K.73b, K.73c, K.73e, K.73a, K.74c, K.74e]. Review [K.74a, K.74b, K.74d, Kab74, Kal75, Kee75, Ket75, Kha74, Kin78, Kin79a, Kle77, Kno71, Kra76, Kup71, Lac71b, Lac71a, Lac72b, Lac72a, Lac74a, Lan73, Lan75, Law75b, Leh71, Lev79, Lin75, Luk73, Lut73, Lyn78, Man71c, Man73, Mar75, McC72, McC78b, Mer77, Mer79, Mer72, Mih75, Mih77, Mil72a, Moe72, Mye74, Nay72, Nel74, O’F73, Odo77, Odo78, Pre77, Qua71, Qua74b, Qua74a, Qua75, Ram73, Ran70a, Ran70c, Ran70b, Rog70d, Rog70b, Rog70a, Rog70c, Ron72, Ros75a, Rus73, Sac75, Sal75a, Sam73, Sax72, Sch74b, Sch75, Sen70, Sen74, Sen75, Ser73, Sha71, Sid75, Sie75, Sim76, Sim79, Sim73, Smi70a, Smi77, Sne72a, Sne74, Sne76, Sne70, Spr72, Spr78, Spr79, SD75, Sta70, Ste75, Sto76, Str75, Sym71, Sys75, Sys76, Sys78, Tan70, Ton72]. Review [Tsi78, Urq73, Wat73, Web73, Weg76, Wei71, Wil79a, Wil70, Win71, Woo75, Woo74b, Woo74a, Zac74, Zie78, Ziz78, Zys70, vdV75, DV79, Ron72]. Reviews [Boa79a, Kin79b]. Revised [Zah75b]. Revuz [Sys76]. Richard [Boa79a, Car77a, Ens78, Gle77, Hub78, Joh75, Kno71, Odo77, Qua71, Smi77, Woo75]. Ridge [BC71, Bro77, DV79, Far75, Far78, GHW79, Haw75, Hem75, HB78, HK70a, HK70b, HK75, KM79, Mar70, McD75, Obe75, Obe77, WC78]. Ries [JK71]. Ries-Smith [JK71]. Rising [Lac74a]. Risk [HP71, Hun70, J.78j]. Risks [Gru73a]. Rivlin [Jer70]. Roach [Man71c]. Robbins [Sim73]. Robert [Cru78a, Hei78, Hol73, J.77i, Kin79a, Moe72, Nel74, Qua75]. Roberts [And75]. Robin [Con70]. Robust [And74, CG76, Dav79b, DM77, For72, HM77, Law75b]. Robustness [Ano73a, HBA71, MHL77b]. Rodger [Sch75]. Roessler [Woo74b]. Roger [Kin79a]. Rogers [Law75b]. Rohatgi [Hac79]. Romier [J.78m]. Romig [KO77]. Ronald [Dye74, Har77]. Root [GWM76, WGM74, WG79]. Roquemore [Ano77f]. Rosenblatt [Hog74]. Ross [J.77f, J.77e, J.77q, J.77p, Kin79b]. Rotatable [HP78]. Roubens [J.77a]. Rough [Kow70]. ’Round [Con70]. Rounding [SB72b]. Rourke [Qua75]. Roy [Gri73, K.71e]. Rozanov [Sys78]. RST [Ros77]. Rudolf [Lev79]. Rule [AB71, Sor73, Wil79a]. Rules [GS71, Koy78]. Run [Dic74, JC76, Rey75, RH78, VM79]. Runs [Bla77, Wat70a]. Runyon [Qua71]. Rutemiller [Bis78]. Ruymgaart [Kra76]. Ryan [J.78g, J.78g].

S [Add79, And75, Ben76, Bri73a, Cam77, Che71, For71, Gen77, Gra72, Gri73, Hah79, Har75, Hul73, Ing78b, J.75h, J.76f, J.78e, J.78p, J.79q, J.79d, J.79h, J.79m, Joh75, K.71a, K.74d, Lac71a, Lac74a, Lan75, Lin75, Mih75, Mih77, Sax72, Sim73, Sne72a, Sne74, Spr72, Ste75, Ton72, Wil70, KB75]. Sachs [Bed74, JH73b]. Safe [Law73]. Sakarovitch [Coo72]. Sakradi [J75e]. Same [Leo70, WK71]. Sample [AB71, Ang74, Ano73g, Ber72b, BF74, BR75, BM78, BH73, CHH^{+}71, CC71, Cho79, DH74, Dye73, EB79, FD70, FR72, GH70, GB72, Gue71a, Gue74, Gut73b, Hah77, Has72, IM70, Jae76, JL79, Kam73, KB75, Lin74, NL75, Nel72a, Ode71, SdV73, Ste75, TJ79, Ury76, WG79, Wio70, Wol77b, J.76g]. Sampled [BF72]. Sampler [J.79a]. Samples [AM79, Ano73k, Bai73, BAB72, Coh65, Coh72, DM78, Ede74, GS71, Hah75, HS72, Kam74, Lac74b, Law71, Lem75, LF79, MF77, McL72, MD71, McM71, MS79, MM73, MHL74, MHL77a, NS79, OL77, SW72, Sin71, TW72b, Zei73, BD73]. Sampling [AB79, Ano72g, Ano77h, Aus73, Bai72, BL74a, Bin74b, BLB73, Car77b, Chi74, Coh70, Coh75, Coh76, CN77, Ehr72, Eld74, EB73, Gro72, Gue71b, HM76a, Hal77a, Hal79, J.78c, J.79t, Joh70a, KO77, Koy78, Lad77, Lev70, Mac76, PS70, Rel70, Sac72, Sch72, SP72, SP74, Sha79, Tak72, Ten72, VM79, Whe70, WB71, Zac70, J.79a, J.78n, Sen70, Tsi78]. Samuel [Fre74a, Smi70a, Spr78]. Sanford [And75]. Satterthwaite [DW72]. Saul [Ran70a]. Savage [J.75h]. Saxena [J.79v]. Scale [DAH73, Dye73, EB77b, Har73a, Kam72, Kam73, Kam74, LR78]. Scaling [Bro77]. Scanning [San72, San73]. Scatterplots [CK75]. Schafer [Bis78]. Schatzoff [Ano71g]. Scheffe [Par78b]. Scheffé’s [Cor77]. Scheme [BJ70b, DS79, Gro72, Luc73, Rea71, Ron77, Ten72]. Schemes [Koy78]. Scheuer [Rea71, WG78]. Schleifer [Ham75]. Schmidt [J.77c]. Schucany [Sal75a]. Science [J.78e, J.78p, Jon76b, Bon76b]. Sciences [Cru78b, Hei76, Leh71, vdV75]. Scientific [Bon77]. Scientist [J.79p]. Scientists [Boa79a, Cas76, Fed74a, Rog70b]. Sclove [Cro78, Lan75]. Scope [ML75]. Scott [Fre73b]. Screening [Kao72, Kle75, LO79, MB74, OB76, OlHS77, Par78b, SM76]. Seal [Hun70]. Sealed [PBT76]. Searle [Hul73]. Seber [J.76c]. Second [BD74, Hok74, OM72, SWD70, SWD74]. Second-Order [BD74, Hok74]. Secondary [Tho73]. Seeding [DV78]. Segment [Yue78]. Segmented [Par78a]. Selected [BMH76, CGO07, Ing78b, J.75c, K.71a, Kam72, Fre73a, Har72, J.75g, J.76h, J.77k, J.78o, K.73c]. Selecting [Ala71, All71, MBC79, Par78b, RRH73, Dud79]. Selection [All74, Ano80c, Bea70, GG70a, Gor76, HH77, Hel74, Hoc72, Kap72, LH70, Man70b, Man71d, McC75a, McD79, MRR78, NW79, Reg76, Riz71, Tho73]. Selection-Multiple [Gor76]. Selection-of-Variables [Hel74]. Self [Boa79a, LS71b, LSM72]. Self-Teaching [Boa79a]. semigraphic [Tuk70]. Sen [Ano70f, Dun72]. Sensitivity [Ant77, MHC77]. Separate [McL74b]. Separating [YB76]. Separation [Sch70]. Sequence [Rif71, LH77]. Sequences [HS71]. Sequencing [Bag70]. Sequential [AS73, Aro76, Bar72, BS79, Cho73, CL70, CL71, DeG70, DS79, Ehr72, Igl73, IR77, KKM79, MB74, O’B73, PBT76, PS70, ST72, SI70, Ehr77, Fle76]. Serfling [J.75f]. Serial [BJ75, JB74]. Series [AB74, BGH73, Cle71, Cle72b, Dav79b, Dun77a, Gru71b, J.79e, Kle77, LR76, Mal74, Mar74, MH72, Nel74, Par72, Pre77, Sar71, Shu71, Spe71, Var70, Weg76, BT74, Cle72a, Hea78, McC78b]. Service [Bag70]. Services [Hay73]. Session [Ano70a]. Set [CR79, Hed71, Leo70]. Sets [BCK61, BCK78]. Sevast’yanov [Spr79]. Several [Ano79g, BF74, Gut73b, Luc76, TM72, BN72]. Shafer [Ing78a]. Shape [AJ77, EB77b, FP77, O’B73, Rou73]. Shaul [J.77i]. shift [LH77]. Shifts [Dav79b]. Shimi [J.77n]. Shipman [And75]. Shooting [And75]. Short [Bla77, Cro78, Nel72a, Str75]. Shortest [AJ77, SS71a]. Should [Hoc72]. Shrunken [McD75, Obe75]. Side [See73]. Sided [FM77, LO79, SS75, SdV73, Wes71, BN72, GPU76, HP73]. Sidney [Gil74]. Siegmund [Dav77, Sim73]. Sigma [Tak72]. Sign [Sta79]. Signal [Shu71]. Significance [Ano73g, Gar71, GB72]. Simple [ABA72, Ano74f, Eng75, Gue72, Hoo79, KKM79, MH72, NH72, Rah76, Ron77, SH79, TMB73, Zel72, GH75]. Simplex [GK77b, ON75, SN77]. Simplified [Box71, EB77a, GH79, MF75b, RS71, Reg76, SP72]. Simulation [Cha70a, GLV74, HK77, J.79m, K.73a, Mer72, Nay72, Sch75, Vin76, Mih75, Mih77]. Simultaneous [Ait74, Bis79, Hah72, Hah75, Has72, HM76b, McK76, Tar76]. Sines [Cha70b]. Single [Ano72g, Ano74f, Gue71b, HM76a, Lad77, MW78, TMB73, Zac70]. Single-Stage [Lad77]. Singly [MN77, NS79]. Singpurwalla [J.77j]. Singularities [MPBD79]. Situations [Hin77b, Ste72b]. Sixth [Fre73b]. Size [AB71, BS72, BSH71, CHH^{+}71, FD70, Gue74, HG78, J.76g, MF77, NL75, Sch72, Wio70]. Size-Biased [Sch72]. Sizes [Ang74, Ano73g, BM78, Fuc78, GB72, Gut73b, SdV73, Ury76]. Skewness [JL79, Rou73]. Slope [HP78, OM72]. Slope-Rotatable [HP78]. Small [Ano73f, BF74, FR72, GH70, GS71, NS79, Web71]. Small-Sample [GH70]. Smallest [Aus73, Nád70a]. Smirnov [BD73, DM78, LR78, PS77]. Smith [JK71, Smi76]. Smoothing [Gro71a, LW71]. Snedecor [K.73d]. Snell [Ham75]. Sobel [Dud79]. Social [Leh71]. Soforthelfer [JH73b]. Solution [And73, AF76, Gri72, Hem75, Kal71, Ros75c, Sam78]. Solutions [Haw72, J.79o]. Some [AvdL70, Ano72e, Ano73h, Ano79g, BH70, BD72a, Bis79, Box70b, BD71, BHME73, BF74, CB75, CG76, Cor75, DW73, Dic74, Dol70, Dun78a, Dun71, Dun78c, EB73, EB74, FR72, Gue74, GM79, HB71, HM76b, Hin77a, Hol75, Kow70, Kow71, Krz79, LL78, Low76, Low74, Mal73, MH72, MCL76, Nel75, OB76, Riz71, Rub76, San73, SR75, SS75, TM72, Tuk70, Weg74, YC71]. Sonquist [Gle78]. Sources [PBT76, Sau73]. Spaces [Ano76e, Sne75b]. spare [Bha73]. Spares [DY70]. Späth [Ens76]. Specific [Car77b, ML75]. Specifications [Car77b]. Spectra [WRLI75]. Spectral [Bri73b, SCSB76, Ran70c]. Spectroscopic [Mal74, Mar74, BT74]. Spectrum [Nea70, Yue78]. Speizer [Ens78]. SPES [Ano70a]. Spline [Lin79, Wol74]. Sports [J.77i, J.78e]. Sprott [Ano79f]. SPRT [Kow71]. Spuriosity [Gut73a, GDF78]. Square [All71, DeG70, Far75, Low74, McL74a, McL74b, ML75, SR75]. Squared [Mie74, Vin76, HK77]. Squares [AF76, BD72a, BCK61, GZ79, GWM76, GPS73, Har73b, Hed71, HM77, J.77b, JK71, MT72, MHL77a, RJ78, See73, Wat74, BCK78]. Srinivasan [Bri73a]. Srivastava [Gri73, J.75a]. Stage [ABA72, Ede74, GG70b, Lad77, Sam78, TA75, Wol77b]. Standard [BMH76, Joi77]. Stanley [Cro78, Gie74, vdV75]. Starr [J.78h]. State [Ano80c, McD79]. Statements [Ano74g]. Stationarily [Shu71]. Stationary [Weg74]. Statistic [Ano73i, Beh72, DW72, HH77, Ken71, Krz79, PS77, PW72, Som70, TW72b, YM74, LR78, Sal75a]. Statistical [Ano70a, Ano73a, BKRS78, Bel70, BAB72, Bon76a, Cha70a, CG76, Dol70, Dud79, EB77a, Eva71, Eva72, Fed74b, Fre73c, FH79, Gen77, Gib74, Gol74, Gru73b, Hac78, HN70, Har77, Hei78, Hil74, Hub78, J.75a, J.75f, J.76d, J.76e, J.78l, J.78f, Joh71b, Joh73c, K.71a, K.71c, K.74c, Lan75, LSYF79, McC78b, Mer79, Ram73, Ran70b, Sau73, Sax72, Spr76, Sta70, TJ79, Wat70b, Wil79a, Woo74a, Boa79a, GW75, Add72, Boy74, Con74, Dav77, Fie73, Fre74b, J.78p, J.79q, K.71e, K.73d, Kal75, Kin78, Lac72b, Rus73, Smi77, Ste75, Win71]. Statisticians [Sne74, Ano71f, K.74e]. Statistics [Ano76f, Ano79g, BH70, BMH76, Boa79a, BF74, Che71, CK75, Cor71, Dav79a, DM78, Fed74a, Fre73b, Fre74a, Gro77, Gue77, Har72, Ing78b, J.75g, J.75h, J.76f, J.76h, J.77d, J.77f, J.77e, J.77l, J.77k, J.77o, J.77q, J.77p, J.78k, J.78o, J.79g, J.79v, J.79f, Joh75, JK74, Jon73, K.74b, Kam72, Kam74, KB75, Kup71, Lan73, Lyn78, MS74, McC72, Mer77, Mil72a, Pap72, Qua74b, Qua74a, Qua75, SR75, Sim76, TM72, Urq73, ZNL79, Zie78, Boa79a, Kin79b, Bha72, Bow77, Bow78, Car77a, Cru78a, Dye74, Ehr77, Eri72, Geo72, Gra72, Gru71a, Hac79, Hal77b, Hei79, Hog74, How77, Ing75, J.75b, J75e, J.76c, J.76a, J.78m, J.78c, J.78a, J.79l, J.79k, J.79o, K.74d, Kin79b, Lac72a, Mye74, Qua71, Qua75, Ros75a, Sch74b, Sen75, Web73, Wei71, Woo74b, Zys70, Add79, Boa79a, Cru78b]. Statistics [Dal77, Dyk76, Gle77, J.79r, J.79s, K.73e, Kin79a, Lac71b, Lac74a, Wat73]. Statistik [Ens77]. Statistische [Bed74, JH73b]. STD [Koy78, Koy79, BS79]. Stein [Wil79a, DV79, Vin76]. Stein-Rule [Wil79a]. Stein’s [HK77]. Stellar [Sau73]. Stepdown [Man70b]. Stephen [Dra79, Odo77]. Stepwise [Goo71, PW72, YC78]. Steutel [Luk73]. Stochastic [Ano73j, BM74, Gil74, Mar73, TW72a, Fre73c, Hun70]. Stone [Gil74, Sax72, Ton72]. Stopping [Sim73]. Storage [Ven70, K.74a]. Straight [For72, WB74]. Straight-Line [WB74]. Strategies [Bon77, J.77i]. Stratified [Gro72, Zac70]. Stratum [CHH^{+}71]. Strength [CH70, Hal77a]. Stress [CH70, Law77b, Man72, MW78, MH77, MH78]. Stress-Strength [CH70]. Stroup [Boa79a]. Structural [Gri79, Kal71]. Structure [FW72, Rog70d, Wil70]. Structures [K.71d, TC75b]. Stuart [Fie73, Gra72]. Student [Boa79a, J.78g, CC71, Rel70, YM74]. studentized [Smi76, Fed75, GKS75, SU79]. Studies [Ano73a, Dil71, J.78p, J.75h, J.77l].

Study [BJN71, Dun78b, Ens78, FP77, Kee75, Lev70, Zah75a]. Sturdy [Qua75]. Stuttering [Pat76b]. Stuttering-Poisson [Pat76b]. Style [Dil71]. Subject [Cha72, Cha70b, HS73]. Subrahmaniam [J.75c, J.75c]. Subsampling [BF72]. Subset [Ber78, DH74, Gor76, Hoc72, McC75b, Pas79]. Subsets [Ait74, HH74, KN74, McK79]. Substantive [Sto76]. Subsystem [Bri74]. Successes [Gre70b, Sap73]. Such [Koy78]. Sufficient [Ing78b]. Suitable [WB75]. Sujit [Ban73]. Sum [BJN71, MT72, NW77, NW79, Rey75, YC71]. Sum-of-Product [YC71]. Summaries [Ano70k, Ano70l, Ano70m, Ano70n]. Summary [Weg72]. Superimposed [BGH73]. Superiority [Hah77]. Surface [HM72, HS73, Luc76, Man73, MC73, ML75]. Surfaces [CO75, Roq76, Sne75a]. Survey [J.75a, J.77m, J.78c, Law75b, Ste75, Gle78, J.79t]. Survival [Law77b, MH77, MH78, Hei76]. Switching [Koy78, Koy79]. Symmetric [Beh72, Luc76, SN77, K.74d]. Symmetric-Simplex [SN77]. Symmetrical [Cor77]. Symposium [Fre73b]. Syntheses [BA70]. System [Bod70, BWL78, DY70, EP71, GC79, Gru71b, Jae74, MG74, Sar71, ST73, Var70, Wan71, WK71, Bha73]. Systems [BH70, BGH73, BM74, Bri74, Bry71, Buc73, CK79, J.78j, JB76, K.73a, MC73, PW77, TC75a, Wig76, Kev75, Bak74, Mer72, Dav70, Mar75].

T [Bow78, Gow72, J.78g, J.79m, K.73d, Lan75, Rog70a, Ron72]. Table [GW75, JK71, TC72]. Tables [Ano71f, BD78, BK70b, BG78, Gar71, Goo71, Har72, J.75g, J.76h, J.77k, J.78p, J.78o, JD78, Lun75, MF73, MDM77, Sne74, SU79, J.78l, Odo78, J.78q, J.79u]. Tabular [Joh73b]. Tabulated [Sim79]. Tailed [Bry74]. Tails [RSO72]. Tanis [Hei78]. Tanur [J.79k]. Tapiero [J.79d]. Tatsuoka [Ger75]. Taylor [Bow77]. Teaching [Boa79a]. Technique [CK75, CM71, MW75a, Som70, Whi70]. Techniques [Eva71, Eva72, Fed74b, Ger75, Gua71, J.76i, J.79i, K.71c, Lin74, MC73, Rel70, SCSB76, Sne73, For73, J.78n, Mar75]. Technology [Dal77, J.79s]. Technometrics [BCK78, GH75, HK77, JTS75, Smi76, Kit81, Ton75, FW72]. Telescoping [HS71]. Tell [Bon77]. Terms [DS77a, DS77b, GG70a, McL74b]. Ternary [CK79]. Test [AS73, Ano75e, Bar72, BJN71, BD78, BF74, BS79, CC71, CF76, DW72, DAH73, Fil75, Fuc78, GR72, Har73a, J.77b, KP71, Lun75, Man72, MG74, MF75a, MS74, MW78, MCM75, McD75, MDM77, MH78, Mie74, Nel72a, Ng76, Obe75, Ode71, PK76, PK79, Pre75, Rah76, ST72, Sch71a, SW72, Sta79, Ste78, SI70, TK72, WDFH73, WB75, Wes71, Wil79a, Wol77a, Wol77b, BD73, LH74]. Testimation [Rif71]. Testing [Ano74f, Ant77, Bai74b, BKRS78, BC79b, Bli72, BK70b, CH73a, Dun74, FP77, GR72, Gre71, Gru73a, HJF78, Igl73, JD78, KA71, Law71, Leo70, MHC77, MS79, NL75, PBT76, Sam78, SLK79, Sha78, Smi77, TMB73, Ziz78]. Testing-Probability [Smi77]. Tests [Ano73g, BJ75, BS71, Bry74, BH73, Cho73, Cho79, CL70, CL71, DG72, Dan70, DW73, Dub73, EB75, EB77b, FR72, Gar71, GC74, GB72, IR77, Jae76, JB74, Kam74, KKM79, KARR75, Kow70, Kra76, LaB77, Law77b, MD71, McM71, MH77, Moe74, MM73, Nád73, Nai74, NK76, NM78, Obe77, SS75, SS74, Sin71, Sin73, YC78, Bel70]. Text [Boa79a]. th [For72, Lik74]. Thanks [Ano75m]. Theil [Fre74b, Han73]. Their [BD72b, Bri73a, Cle72b, CG76, HR75, NH72, Par72, Spr78, Zel72]. Theodore [Jer70]. Theoretical [Gua71, J.79o, Ing75]. Theory [Fre73c, Hac79, Hun70, Ing78a, K.73c, Kra76, Lev79, Lut73, Man71c, Nel72b, NH73, NK76, NM78, Odo77, Pre77, Rog70d, Sam73, Sid75, Sim76, Sim73, Smi77, Sta70, Ste75, Zac74, Ben76, Dun77b, Hun74, Kee75, Moe72, Sax72, Sen70, Spr78, Sys75, Tan70, Ton72, Weg76, J.77n]. Third [K.74e]. Thomas [Mer72]. Thomasion [Rog70d]. Thompson [Ham75]. Three [Ano73f, BLB73, Coh75, Coh76, CN77, DS77a, Ede74, EB75, GG70b, Hed71, Lem75, Luc78, MF75a, MBC79, McL74b, MB78a, MB78b, SP77, Sor72b, Web71, Jac80]. Three-Level [Ano73f, Luc78, MB78a, MB78b, Web71]. Three-Parameter [CN77, EB75]. Three-Stage [Ede74]. Threshold [MF75a]. Thumb [AB71]. Tiao [Hil74]. Time [AvdL70, AB74, CB75, CF76, Cle71, Cle72a, Cle72b, CM71, Dav79b, Dun77a, Gru71b, HD70, J.76f, J.79e, LR76, McC78b, MH72, Nel74, Par72, Pre77, Rif71, Sar71, Wan71, Kle77, Weg76]. Time-to-Fail [Gru71b]. Times [AJ77, Bag70, DY70, HT74, Kam77, Man72, Nád73, PK76, SS74, Bha73, J.78q]. Tinker [Ens78]. Title [K.71a]. Titles [J.77f, J.77q]. Toeplitz [Ray70]. Tolerance [Dun78c, EB78c, Fau70, FD70, Fra77, FKRO71, Law75a, MF73, MF77, RSO72, TBA70, TJ79, GPU76]. Tolerances [JR72]. Tolerancing [Eva71, Eva72]. Tong [JTS75]. Tool [BG78, Fed75]. Tools [Fed74b]. Total [CHH^{+}71]. Toutenburg [Boa79b]. Tracking [Pas79]. Tracy [K.74d]. Trade [Kee75, Kno71]. Traffic [Ano73j, Ano80c, Mar73, McD79]. Training [Wat70b]. Trajectory [Bus71]. Transform [Ano84, HL76, IC79]. Transformation [BH74, Cha79, GH79, HH73, Tho75, WY78]. Transformations [BWL78, Kow70, Sen74, TK72]. Transformed [Fuc78]. Transient [BGH73]. Transients [Bri73b]. Transition [WB74]. Treatment [Fed75, SOE75]. Treatments [BH77, SdV73, BN72]. Tree [J.77j]. Trending [LR76]. Trials [Gre70b, Sap73, Fle76]. Triangles [CK79]. Trimmed [MHL77a, MHL77b]. Triple [Pat76b]. Trirogoff [Bis72b]. trivariate [TC72]. Truncated [AS73, Coh72, DM78, Gro71b, Gue71a, Ham72a, KKM79, KB75, SI70]. Truncation [Man75]. Truth [Bon76b, Bon77]. Tsokos [J.77n, Zac74]. Tucker [Har70]. Tukey [J.77d, J.77f, J.77e, J.77o, J.77q, J.77p, Law75b, Qua74b, Qua74a, Wes71]. Tuma [K.71b]. Turner [vdV75]. Turoff [J.76i]. Two [Ano73f, AJ77, BE73, BL71, BC79b, Beh70, BH77, BG78, Cha79, CH73a, Cho79, Dan78, DM77, DL70, DAH73, Dun78c, EB74, EB75, EB78c, EB79, GW75, Gri70, Gru71c, Hah77, HW75, Har73a, HJ76, Jae70, Jae76, Joh70c, Joh70b, JD78, Kam74, KB75, KP71, Law77a, Leo70, LO79, Lik74, Low76, Man72, MF75a, MDM77, McK76, MD71, McM71, MPB70, MM73, MA78, Nea70, Pie73, Ray72, Sal75b, Sam78, SG71, SS76, SdV73, TW72b, Tho73, TA75, Var70, WB74, Web71, WDFH73, Wol77b, AB71, GPU76, And75]. Two- [Ano73f, Web71]. Two-Parameter [BE73, EB74, EB75, EB78c, Lik74, Man72, TW72b, GPU76]. Two-Phase [BC79b]. Two-Point [And75]. Two-Regime [WB74]. Two-Sample [Cho79, EB79, Wol77b]. Two-Sided [LO79, SdV73]. Two-Stage [Sam78, TA75, Wol77b]. Two-Way [BG78, Dan78, GW75, HW75, HJ76, JD78, Low76, MA78]. Tyn [Cha75a]. Type [Ano79g, Bry71, Gue74, MC72, Sha79, Sha78, Sin73, TM72, DW72, Har71, TW72b]. Type-I [DW72]. Types [Luc76, McL74b].

U [Cha75a]. Udny [Fie73]. Unbalanced [Hin77b, See70]. Unbiased [BR73, EB77b, EB78a, MF77, NH73, SG71, Sch76, SS78, WK77]. Unconditional [MDM77]. Underlying [Law73]. Unequal [BD78, Bis79, Leo75, vN79]. Uniform [McD76, TC73, TS79, TC72]. Uniformity [Ng76]. Uniformly [EB77b]. Unit [MW78]. Units [Har74]. Univariate [Che71, Gut73a, Sor73, GDF78]. Unknown [DAH73, Har73a, McM71, MM73, OB76, Tak72]. Unknown-Sigma [Tak72]. Unmodeled [Bus71]. Unweighted [Cor77]. Update [Cor79]. upon [Ano72g, Gue71b]. Upper [KARR75]. Upton [Sim79]. Uranium [CBI79]. Urn [Spr78]. Use [Ano71f, Ano84, Bai74a, BH78, BMH76, CO75, Hub78, IC79, Kin79b, Kit79, LM72, NGM70, Nea70, PW72, Sne71, Sne75a, Woo73, Kit81]. Used [Hoc72, JH73a, Leo70]. Useful [BK70b, Krz79, Whi70]. Uses [RTDM79]. Using [ABA72, AA74, Aus73, Bod70, BCK61, BR75, Cor77, Gre71, Har74, HH73, JB76, Kam72, KN74, Mar75, NW79, Pat76a, Rah76, Sin73, SLM74, TS79, WB74, Wig76, Zei73, BCK78, McC78a].

V [Bel70, Bis72b, Hac79, Hei78, Ing75, J.76b, J.79o, J.79u, Luc73, Sac75, Sen75, Spr79, Wat73]. V. [HH73]. Validation [GHW79, Sne77]. Validatory [Wol78]. Validity [Whe75]. Value [And75, Bai72, Bai73, Eas78, EB73, EB74, Eng75, EB77a, Has72, Law75a, Law78, MF73, MF77, NM78, TW72b]. Values [BK70b, BDH70, Dun71, Har74, HG73, JH79, Joh73b, MBC79, Pre71, Rub76, Sin72, You72a, Zah75b]. Variable [Ait74, All74, Bea70, BMP77, Hun74, Kus79, Man70b, Man71d, McC75a, McK79]. Variables [All71, BL71, BK70b, Cas76, Cha72, Fuc78, HH77, Har79, Hel74, LH70, MBC79, NW79, OlHS77, Sch70, SB72b, Tho73, Whe70, WY78, LH77]. Variance [BD72b, Bin74b, BD78, Bis79, BH74, BR73, CS76, CK79, Dav79b, GG70b, GH79, GS71, Hal79, Hel74, HL76, JS76, Krz79, Low76, Man71a, MD71, McM71, MM73, MA78, Nai74, Rel70, Sal75b, Sch76, See70, SW72, Sne76, SS78, TA71, WC70, WB75, Whe70, Wio70, WK77]. Variances [BD78, Bis79, Cha70b, Cor77, GG70a, HW75, Leo75, Low70, MN77, TJ79]. Variate [McD76]. Variates [AR73]. Variation [AR73, Ber72b, IM70, Zei73].

Variational [FM71]. Varieties [Joh78, Joh73a]. Vasant [Ing78b]. Vasudevan [Bri73a]. Vaughn [Woo75]. Vector [Nád70a]. Vectors [JH73a]. Verification [Lev79]. Version [Wes71]. versus [CM71, Cor77]. Vertices [SM74]. Via [Mar76, Bin74a]. Vick [Bon76a]. Viewed [WB71]. Viewpoint [WG78]. Vincze [J75e]. Vinod [HK77]. Virginia [Hei76]. Visual [San72, San73]. Vital [Car77a]. Vocabulary [Dil71]. Vol [J.75g, J.76h, J.78o]. Volume [Ano70o, Ano71l, Ano72l, Ano73p, Ano74l, Ano75l, Ano76o, Ano77q, Ano78o, Ano78p, Ano79p, Bor72, Cra74, Fre73b, Har72, J.77k, J.79q, J.79g, J.79f, K.71d, K.73b, K.73a, Lac71b, Sne74]. VOQL [Hal79]. Vote [Ano75m]. vs [LW71, MF75a, Scl72].

W [Add72, Add79, Ano74e, Bha72, Boy74, Dav70, Fre73a, Fre73c, Gle77, GH75, Hay73, Hea78, Hel71a, Hub78, J.76d, J.76e, J.76a, J.77d, J.77f, J.77e, J.77o, J.77q, J.77p, J.78n, J.78l, K.73d, Lac71b, Lac74a, Lan75, Law75b, Leh71, Luk73, Mil72a, O’F73, Odo77, Qua74b, Qua74a, Sne76, Spr72]. Waiting [CM71, HD70]. Waiting-Time [HD70]. Wald [Sha79]. Walpole [Fed74a, Dye74, Har77]. Wani [Gol74]. Wasan [Rog70a]. Water [K.74a, Ven70]. Watts [Ano75f, Ran70c]. Way [BG78, Dan78, GW75, HW75, HJ76, JD78, Kus79, Low76, MA78, Nai74]. Wayne [McC78b]. Weak [Ber72b]. Weather [Dav79a, Mie74, SOE75]. Weibull [Ano73k, Ant77, Eas78, AJ77, Bai72, Bai73, BL74b, BAB72, Coh65, Coh73, Coh75, D’A71a, D’A71b, Dan70, DA73, EB73, EB74, Eng75, EB75, EB77a, EB78b, EB79, Fal70, Fin76, Gro71b, HBA71, KARR75, Law73, Law75a, Law78, LL78, Lem75, Man70a, Man71b, Man72, MF73, MF75a, MF75b, MF77, MHC77, MN77, NM78, Rou73, SS76, SA79, SW75, TBA70, TW72b]. Weighing [Cam77]. Weight [Bin74b]. Weighted [Cor77, Har73b, Hin77a, JK71, NW79, Ste70, WC70]. Weighting [BH74]. Weights [GZ79]. Weischedel [Gor76]. Welch [Ros75c]. Wetherill [Boa79a, Ehr77, Tsi78]. Wheeler [J.76b]. Where [Sar71]. Whether [Leo70]. Which [Bri73b, BF74, Cha70b, DY70, Gru71b, Hoc72, Luc76, RSO72, Sta79, Ste72b]. White [Ano80b]. Whitehouse [Mar75]. Whittemore [J.79h]. Wide [Tak72]. Wilf [J.79q]. Wilkinson [J.79m]. Wilks [Gra72]. Willems [Bak74]. William [Gle78, J.79k, Mil72a, Sal75a, Sam73, Woo74a]. Williams [Dil71, J.77l, J.79a]. Windowing [Yue78]. Winkler [Hol73, Lac71b]. Winn [Kin79b]. Winsorized [MHL77a, YF76]. Wiorkowski [Cro71]. Within [BR79, J.76e, Sap73]. Within-Group [BR79]. Within-Groups [J.76e]. without [Gre71]. Wolfe [Ano78e, Gib74]. Wolfowitz [Sha79]. Wolock [Ano74e]. Wonnacott [Bow78, Bow78]. Wood [Sne72a]. Work [Buc73]. Working [Ens78]. Works [CGO07]. World [Boa79a].

X [Sau73, Ury72]. X-Ray [Sau73].

Yates [Ber72a]. Year [Ens78, Ven70]. Yonathan [Cha75b]. Youden [DeG70]. Yu [Sen75]. Yule [Fie73, Fie73]. Yvonne [Odo77].

Z [J.78l]. Zeleny [J.78h]. Zellner [Goo76, Goo75, J.75h]. Zemroch [J.79u]. Zero [Nád73, PK76, SS74]. Ziegler [J.76a]. Zionts [Lin75].

References

Alldredge:1974:MLE

[AA74] J. Richard Alldredge and David W. Armstrong. Maximum likelihood estimation for the multinomial distribution using geometric progamming. Technometrics, 16(4):585–587, November 1974. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267611.

Al-Bayyati:1971:RTD

[AB71] H. A. Al-Bayyati. A rule of thumb for determining a sample size in comparing two proportions. Technometrics, 13(3):675–677, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267178.

Anderssen:1974:TSA

[AB74] R. S. Anderssen and Peter Bloomfield. A time series approach to numerical differentiation. Technometrics, 16(1):69–75, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267494.

Abraham:1979:SIF

[AB79] Bovas Abraham and George E. P. Box. Sampling interval and feedback control. Technometrics, 21(1):1–8, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268575.

Al-Bayyati:1972:DSE

[ABA72] H. A. Al-Bayyati and J. C. Arnold. On double-stage estimation in simple linear regression using prior knowledge. Technometrics, 14(2):405–414, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267431.

Addelman:1972:BRB

[Add72] Sidney Addelman. Book review: Statistical Design and Analysis of Experiments by Peter W. M. John. Technometrics, 14(1):234–235, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266937.

Addelman:1979:BRB

[Add79] Sidney Addelman. Book review: Statistics for Experimenters by G. E. P. Box; W. G. Hunter; J. S. Hunter. Technometrics, 21(3):387–388, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267766.

Armstrong:1976:BBS

[AF76] R. D. Armstrong and E. L. Frome. A branch-and-bound solution of a restricted least squares problem. Technometrics, 18(4):447–450, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268660.

Aitkin:1974:SIC

[Ait74] Murray A. Aitkin. Simultaneous inference and the choice of variable subsets in multiple regression. Technometrics, 16(2):221–227, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267942.

Ashour:1977:SCI

[AJ77] S. K. Ashour and P. W. Jones. Shortest confidence intervals for the shape parameter of a Weibull distribution based on two adjacent failure times. Technometrics, 19(2):201–203, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268630.

Antle:1971:NCM

[AK71] Charles E. Antle and Lawrence A. Klimko. A note on a conjecture of Mantel concerning quadratic programming. Technometrics, 13(2):425–427, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266803.

Alam:1971:SMP

[Ala71] Khursheed Alam. On selecting the most probable category. Technometrics, 13(4):843–850, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266959.

AGC:1972:AR

[Ale72] Alexander Grant & Co. Auditors’ report. Technometrics, 14(2):??, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267446.

Allen:1971:MSE

[All71] David M. Allen. Mean square error of prediction as a criterion for selecting variables. Technometrics, 13(3):469–475, August 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267161.

Allen:1974:RBV

[All74] David M. Allen. The relationship between variable selection and data augmentation and a method for prediction. Technometrics, 16(1):125–127, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267500.

Angers:1979:NPI

[AM79] Claude Angers and Georges McLaughlin. A note on prediction intervals for large future samples. Technometrics, 21(3):383–385, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267765.

Anderson:1970:BRB

[And70] Virgil L. Anderson. Book review: Evolutionary Operation: A Method for Increasing Industrial Productivity by George E. P. Box; Norman R. Draper. Technometrics, 12(1):172–173, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267365.

Andrew:1973:SEI

[And73] A. L. Andrew. Solution of equations involving centrosymmetric matrices. Technometrics, 15(2):405–407, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266998.

Andrews:1974:RMM

[And74] D. F. Andrews. A robust method for multiple linear regression. Technometrics, 16(4):523–531, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267603.

Andrew:1975:BRB

[And75] A. L. Andrew. Book review: Two-Point Boundary Value Problems: Shooting Methods by Sanford M. Roberts; Jerome S. Shipman. Technometrics, 17(2):277–278, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268370.

Anderson:1976:LE

[And76] O. D. Anderson. Letter to the Editor. Technometrics, 18(2):245, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267541. See errata

[Ano77f].

Anderson:1977:LE

[And77] O. D. Anderson. Letter to the Editor. Technometrics, 19(2):223, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268645.

Angers:1974:GME

[Ang74] Claude Angers. A graphical method to evaluate sample sizes for the multinomial distribution. Technometrics, 16(3):469–471, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267680.

Anonymous:1970:APP

[Ano70a] Anonymous. Abstracts of papers presented at SPES Session of the 129th Annual Meeting of the American Statistical Association. Technometrics, 12(1):191–206, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267376.

Anonymous:1970:BMa

[Ano70b] Anonymous. Back matter. Technometrics, 12(1):??, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267377.

Anonymous:1970:BMb

[Ano70c] Anonymous. Back matter. Technometrics, 12(2):438, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267255.

Anonymous:1970:BMc

[Ano70d] Anonymous. Back matter. Technometrics, 12(3):725–726, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267227.

Anonymous:1970:BMd

[Ano70e] Anonymous. Back matter. Technometrics, 12(4):939–940, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267345.

Anonymous:1970:EMH

[Ano70f] Anonymous. Errata to Margolin, Harter, Good, Hoerl and Sen. Technometrics, 12(3):721–724, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267226. See [HK70a].

Anonymous:1970:FMa

[Ano70g] Anonymous. Front matter. Technometrics, 12(1):181–190, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267346.

Anonymous:1970:FMb

[Ano70h] Anonymous. Front matter. Technometrics, 12(2):421–422, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267228.

Anonymous:1970:FMc

[Ano70i] Anonymous. Front matter. Technometrics, 12(3):??, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267194.

Anonymous:1970:FMd

[Ano70j] Anonymous. Front matter. Technometrics, 12(4):i–ii, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267317.

Anonymous:1970:SNBa

[Ano70k] Anonymous. Summaries of new books. Technometrics, 12(1): 183–186, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267374.

Anonymous:1970:SNBb

[Ano70l] Anonymous. Summaries of new books. Technometrics, 12(2):427– 431, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267252.

Anonymous:1970:SNBc

[Ano70m] Anonymous. Summaries of new books. Technometrics, 12(3):715– 718, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267224.

Anonymous:1970:SNBd

[Ano70n] Anonymous. Summaries of new books. Technometrics, 12(4): 933–936, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267343.

Anonymous:1970:VI

[Ano70o] Anonymous. Volume information. Technometrics, 12(4):941–944, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267316.

Anonymous:1971:AAE

[Ano71a] Anonymous. Addendum: The augmentation of experimental data to maximize |X′X|. Technometrics, 13(4):927, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266979.

Anonymous:1971:BMa

[Ano71b] Anonymous. Back matter. Technometrics, 13(1):219–220, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267098.

Anonymous:1971:BMb

[Ano71c] Anonymous. Back matter. Technometrics, 13(2):??, May 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266817.

Anonymous:1971:BMc

[Ano71d] Anonymous. Back matter. Technometrics, 13(3):??, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267193.

Anonymous:1971:BMd

[Ano71e] Anonymous. Back matter. Technometrics, 13(4):929–938, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266980.

Anonymous:1971:BRB

[Ano71f] Anonymous. Book review: Abridged Tables for Use of Examination Candidates by The Institute of Statisticians. Technometrics, 13(1):213–214, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267095.

Anonymous:1971:EKS

[Ano71g] Anonymous. Errata to Kumar and Schatzoff. Technometrics, 13(3): 712, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267192.

See [KP71].

Anonymous:1971:FMa

[Ano71h] Anonymous. Front matter. Technometrics, 13(1):215, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267071.

Anonymous:1971:FMb

[Ano71i] Anonymous. Front matter. Technometrics, 13(2):459–460, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266784.

Anonymous:1971:FMc

[Ano71j] Anonymous. Front matter. Technometrics, 13(3):??, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267159.

Anonymous:1971:FMd

[Ano71k] Anonymous. Front matter. Technometrics, 13(4):??, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266947.

Anonymous:1971:VI

[Ano71l] Anonymous. Volume information. Technometrics, 13(4):933–937, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266946.

Anonymous:1972:BMa

[Ano72a] Anonymous. Back matter. Technometrics, 14(1):246, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266945.

Anonymous:1972:BMb

[Ano72b] Anonymous. Back matter. Technometrics, 14(2):??, May 1972. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267447.

Anonymous:1972:BMc

[Ano72c] Anonymous. Back matter. Technometrics, 14(3):811, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267315.

Anonymous:1972:BMd

[Ano72d] Anonymous. Back matter. Technometrics, 14(4):988–990, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267158.

Anonymous:1972:CFD

[Ano72e] Anonymous. Corrigendum: “Factorial Designs, the |X′X| Criterion, and Some Related Matters”. Technometrics, 14(2):511, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267445. See [BD71].

Anonymous:1972:CLF

[Ano72f] Anonymous. Corrigendum: “On Lack of Fit”. Technometrics, 14(1):245, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266944. See [DH71].

Anonymous:1972:CDS

[Ano72g] Anonymous. Corrigendum: “On the Determination of Single Sampling Attribute Plans Based upon a Linear Cost Model and a Prior Distribution”. Technometrics, 14(4):987, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267157. See [Gue71b].

Anonymous:1972:FMa

[Ano72h] Anonymous. Front matter. Technometrics, 14(1):239–240, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266912.

Anonymous:1972:FMb

[Ano72i] Anonymous. Front matter. Technometrics, 14(2):509–510, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267417.

Anonymous:1972:FMc

[Ano72j] Anonymous. Front matter. Technometrics, 14(3):??, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267280.

Anonymous:1972:FMd

[Ano72k] Anonymous. Front matter. Technometrics, 14(4):??, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267130.

Anonymous:1972:VI

[Ano72l] Anonymous. Volume information. Technometrics, 14(4):991–994, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267129.

Anonymous:1973:ABR

[Ano73a] Anonymous. Annotated bibliography on robustness studies of statistical procedures. Technometrics, 15(2):424, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267008.

Anonymous:1973:BMa

[Ano73b] Anonymous. Back matter. Technometrics, 15(1):??, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266848.

Anonymous:1973:BMb

[Ano73c] Anonymous. Back matter. Technometrics, 15(2):433–436, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267015.

Anonymous:1973:BMc

[Ano73d] Anonymous. Back matter. Technometrics, 15(3):658, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266882.

Anonymous:1973:BMd

[Ano73e] Anonymous. Back matter. Technometrics, 15(4):953–960, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267416.

Anonymous:1973:CSI

[Ano73f] Anonymous. Corrigenda: “Small Incomplete Factorial Experiment Designs for Two- and Three-Level Factors”. Technometrics, 15(4):951, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267415. See [Web71].

Anonymous:1973:CES

[Ano73g] Anonymous. Corrigendum: “Extension of Sample Sizes and Percentage Points for Significance Tests of Outlying Observations”. Technometrics, 15(2):429, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267012. See [GB72].

Anonymous:1973:CFD

[Ano73h] Anonymous. Corrigendum: “Factorial Designs, the |X′X| Criterion and Some Related Matters”. Technometrics, 15(2):430, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267013. See [BD71].

Anonymous:1973:ENS

[Ano73i] Anonymous. Erratum: A note on the C_{p} statistic. Technometrics, 15(3):657, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266881. See [Ken71].

Anonymous:1973:ESM

[Ano73j] Anonymous. Erratum: “A Stochastic Model of Microscale Air Pollution from Highway Traffic”. Technometrics, 15(4):949, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267414. See

[Mar73].

Anonymous:1973:EML

[Ano73k] Anonymous. Erratum: “Maximum-Likelihood Estimation of the Parameters of Gamma and Weibull Populations from Complete and from Censored Samples”. Technometrics, 15(2):431, May 1973. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267014. See [Coh65].

Anonymous:1973:FMa

[Ano73l] Anonymous. Front matter. Technometrics, 15(1):205–206, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266818.

Anonymous:1973:FMb

[Ano73m] Anonymous. Front matter. Technometrics, 15(2):??, May 1973. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266981.

Anonymous:1973:FMc

[Ano73n] Anonymous. Front matter. Technometrics, 15(3):??, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266849.

Anonymous:1973:FMd

[Ano73o] Anonymous. Front matter. Technometrics, 15(4):659, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267379.

Anonymous:1973:VI

[Ano73p] Anonymous. Volume information. Technometrics, 15(4):955–959, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267378.

Anonymous:1974:BMa

[Ano74a] Anonymous. Back matter. Technometrics, 16(1):143–144, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267514.

Anonymous:1974:BMb

[Ano74b] Anonymous. Back matter. Technometrics, 16(2):??, May 1974. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267967.

Anonymous:1974:BMc

[Ano74c] Anonymous. Back matter. Technometrics, 16(3):??, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267690.

Anonymous:1974:BMd

[Ano74d] Anonymous. Back matter. Technometrics, 16(4):643–648, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267630.

Anonymous:1974:BRB

[Ano74e] Anonymous. Book review: Cyclic Designs by J. A. John; F. W. Wolock; H. A. David. Technometrics, 16(1):138, February 1974. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267506.

Anonymous:1974:CTS

[Ano74f] Anonymous. Corrigendum: “Testing for a Single Outlier in Simple Linear Regression”. Technometrics, 16(2):331, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267965. See [TMB73].

Anonymous:1974:FSA

[Ano74g] Anonymous. Financial statements and auditors’ report. Technometrics, 16(2):??, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267966.

Anonymous:1974:FMa

[Ano74h] Anonymous. Front matter. Technometrics, 16(1):??, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267484.

Anonymous:1974:FMb

[Ano74i] Anonymous. Front matter. Technometrics, 16(2):333, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267934.

Anonymous:1974:FMc

[Ano74j] Anonymous. Front matter. Technometrics, 16(3):??, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267662.

Anonymous:1974:FMd

[Ano74k] Anonymous. Front matter. Technometrics, 16(4):??, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267598.

Anonymous:1974:VI

[Ano74l] Anonymous. Volume information. Technometrics, 16(4):639–647, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267597.

Anonymous:1975:BMa

[Ano75a] Anonymous. Back matter. Technometrics, 17(1):??, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268025.

Anonymous:1975:BMb

[Ano75b] Anonymous. Back matter. Technometrics, 17(2):283–285, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268374.

Anonymous:1975:BMc

[Ano75c] Anonymous. Back matter. Technometrics, 17(3):396, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268095.

Anonymous:1975:BMd

[Ano75d] Anonymous. Back matter. Technometrics, 17(4):521, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268451.

Anonymous:1975:CPP

[Ano75e] Anonymous. Corrigendum: “The Probability Plot Correlation Coefficient Test for Normality”. Technometrics, 17(4):520, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268450. See [Fil75].

Anonymous:1975:EHW

[Ano75f] Anonymous. Errata to Hamdan and Watts. Technometrics, 17 (1):147, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268024. See [Ham74, WB74].

Anonymous:1975:FMa

[Ano75g] Anonymous. Front matter. Technometrics, 17(1):??, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267992.

Anonymous:1975:FMb

[Ano75h] Anonymous. Front matter. Technometrics, 17(2):??, May 1975. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268344.

Anonymous:1975:FMc

[Ano75i] Anonymous. Front matter. Technometrics, 17(3):??, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268061.

Anonymous:1975:FMd

[Ano75j] Anonymous. Front matter. Technometrics, 17(4):397–398, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268425.

Anonymous:1975:PR

[Ano75k] Anonymous. Publications received. Technometrics, 17(4):516, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268448.

Anonymous:1975:VI

[Ano75l] Anonymous. Volume information. Technometrics, 17(4):517–523, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268424.

Anonymous:1975:VT

[Ano75m] Anonymous. A vote of thanks. Technometrics, 17(1):??, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267993.

Anonymous:1976:BMa

[Ano76a] Anonymous. Back matter. Technometrics, 18(1):126, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267933.

Anonymous:1976:BMb

[Ano76b] Anonymous. Back matter. Technometrics, 18(2):247–248, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267543.

Anonymous:1976:BMc

[Ano76c] Anonymous. Back matter. Technometrics, 18(3):366, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268755.

Anonymous:1976:BMd

[Ano76d] Anonymous. Back matter. Technometrics, 18(4):506, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268674.

Anonymous:1976:EED

[Ano76e] Anonymous. Erratum: “Experimental Designs for Quadratic Models in Constrained Mixture Spaces”. Technometrics, 18(2):246, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267542. See

[Sne75b].

Anonymous:1976:EPS

[Ano76f] Anonymous. Erratum: “Progress in Statistics”. Technometrics, 18(3):364, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268753. See [J75e].

Anonymous:1976:FMa

[Ano76g] Anonymous. Front matter. Technometrics, 18(1):??, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267909.

Anonymous:1976:FMb

[Ano76h] Anonymous. Front matter. Technometrics, 18(2):??, May 1976. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267515.

Anonymous:1976:FMc

[Ano76i] Anonymous. Front matter. Technometrics, 18(3):??, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268733.

Anonymous:1976:FMd

[Ano76j] Anonymous. Front matter. Technometrics, 18(4):369–370, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268649.

Anonymous:1976:PRa

[Ano76k] Anonymous. Publications received. Technometrics, 18(1):122–123, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267930.

Anonymous:1976:PRb

[Ano76l] Anonymous. Publications received. Technometrics, 18(2):244, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267540.

Anonymous:1976:PRc

[Ano76m] Anonymous. Publications received. Technometrics, 18(3):364, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268752.

Anonymous:1976:PRd

[Ano76n] Anonymous. Publications received. Technometrics, 18(4):503, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268672.

Anonymous:1976:VI

[Ano76o] Anonymous. Volume information. Technometrics, 18(4):507–508, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268648.

Anonymous:1977:ACO

[Ano77a] Anonymous. Addendum and corrigendum to Odoroff and Ehrenfeld. Technometrics, 19(4):520, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267905. See [Odo77, Ehr77].

Anonymous:1977:BMa

[Ano77b] Anonymous. Back matter. Technometrics, 19(1):107–108, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268274.

Anonymous:1977:BMb

[Ano77c] Anonymous. Back matter. Technometrics, 19(2):225–226, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268647.

Anonymous:1977:BMc

[Ano77d] Anonymous. Back matter. Technometrics, 19(3):352, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267718.

Anonymous:1977:BMd

[Ano77e] Anonymous. Back matter. Technometrics, 19(4):523–525, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267908.

Anonymous:1977:ERA

[Ano77f] Anonymous. Errata to Roquemore and Anderson. Technometrics, 19(1):106, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268273. See [Roq76, And76].

Anonymous:1977:ELE

[Ano77g] Anonymous. Erratum: Letter to the Editor. Technometrics, 19(4):522, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267907.

Anonymous:1977:EOC

[Ano77h] Anonymous. Erratum: “Optimal Choice of the Sampling Interval for Discrete Process Control”. Technometrics, 19(2):224, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268646. See [Mac76].

Anonymous:1977:FMa

[Ano77i] Anonymous. Front matter. Technometrics, 19(1):??, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268247.

Anonymous:1977:FMb

[Ano77j] Anonymous. Front matter. Technometrics, 19(2):??, May 1977. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268619.

Anonymous:1977:FMc

[Ano77k] Anonymous. Front matter. Technometrics, 19(3):??, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267691.

Anonymous:1977:FMd

[Ano77l] Anonymous. Front matter. Technometrics, 19(4):??, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267875.

Anonymous:1977:PRa

[Ano77m] Anonymous. Publications received. Technometrics, 19(1):104, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268271.

Anonymous:1977:PRb

[Ano77n] Anonymous. Publications received. Technometrics, 19(2):221, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268643.

Anonymous:1977:PRc

[Ano77o] Anonymous. Publications received. Technometrics, 19(3):347, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267716.

Anonymous:1977:PRd

[Ano77p] Anonymous. Publications received. Technometrics, 19(4):520–521, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267906.

Anonymous:1977:VI

[Ano77q] Anonymous. Volume information. Technometrics, 19(4):526–527, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267874.

Anonymous:1978:BMa

[Ano78a] Anonymous. Back matter. Technometrics, 20(1):108, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268177.

Anonymous:1978:BMb

[Ano78b] Anonymous. Back matter. Technometrics, 20(2):220, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268732.

Anonymous:1978:BMc

[Ano78c] Anonymous. Back matter. Technometrics, 20(3):??, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268151.

Anonymous:1978:BMd

[Ano78d] Anonymous. Back matter. Technometrics, 20(4):505, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267661.

Anonymous:1978:EWB

[Ano78e] Anonymous. Errata to Wolfe, Bose and Bradu. Technometrics, 20(2): 219, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268731.

See [Wol77a, BCK61, BG78].

Anonymous:1978:FMa

[Ano78f] Anonymous. Front matter. Technometrics, 20(1):??, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268152.

Anonymous:1978:FMb

[Ano78g] Anonymous. Front matter. Technometrics, 20(2):i–iii, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268700.

Anonymous:1978:FMc

[Ano78h] Anonymous. Front matter. Technometrics, 20(3):??, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268127.

Anonymous:1978:FMd

[Ano78i] Anonymous. Front matter. Technometrics, 20(4):353–354, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267632.

Anonymous:1978:FMe

[Ano78j] Anonymous. Front matter. Technometrics, 20(4):??, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267570.

Anonymous:1978:PRa

[Ano78k] Anonymous. Publications received. Technometrics, 20(1):107, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268176.

Anonymous:1978:PRb

[Ano78l] Anonymous. Publications received. Technometrics, 20(2):217, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268729.

Anonymous:1978:PRc

[Ano78m] Anonymous. Publications received. Technometrics, 20(3):323, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268150.

Anonymous:1978:PRd

[Ano78n] Anonymous. Publications received. Technometrics, 20(4):504, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267660.

Anonymous:1978:VIa

[Ano78o] Anonymous. Volume information. Technometrics, 20(4):506–508, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267631.

Anonymous:1978:VIb

[Ano78p] Anonymous. Volume information. Technometrics, 20(4):509–629, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267569.

Anonymous:1979:BMa

[Ano79a] Anonymous. Back matter. Technometrics, 21(1):140, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268601.

Anonymous:1979:BMb

[Ano79b] Anonymous. Back matter. Technometrics, 21(2):276, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268538.

Anonymous:1979:BMc

[Ano79c] Anonymous. Back matter. Technometrics, 21(3):??, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267782.

Anonymous:1979:BMd

[Ano79d] Anonymous. Back matter. Technometrics, 21(4):588–594, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268311.

Anonymous:1979:CMB

[Ano79e] Anonymous. Corrigenda to Muse and Bishop. Technometrics, 21(1):139, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268600. See [MA78, Bis78].

Anonymous:1979:CSO

[Ano79f] Anonymous. Corrigenda to Sprott and Odoroff. Technometrics, 21 (2):275, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268537.

See [Spr78, Odo78].

Anonymous:1979:CSG

[Ano79g] Anonymous. Corrigendum: “Some Grubbs-Type Statistics for the Detection of Several Outliers”. Technometrics, 21(3):396, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267781. See

[TM72].

Anonymous:1979:FMa

[Ano79h] Anonymous. Front matter. Technometrics, 21(1):??, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268574.

Anonymous:1979:FMb

[Ano79i] Anonymous. Front matter. Technometrics, 21(2):??, May 1979. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268510.

Anonymous:1979:FMc

[Ano79j] Anonymous. Front matter. Technometrics, 21(3):??, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267747.

Anonymous:1979:FMd

[Ano79k] Anonymous. Front matter. Technometrics, 21(4):i–vi, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268276.

Anonymous:1979:PRa

[Ano79l] Anonymous. Publications received. Technometrics, 21(1):136–137, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268598.

Anonymous:1979:PRb

[Ano79m] Anonymous. Publications received. Technometrics, 21(2):274, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268535.

Anonymous:1979:PRc

[Ano79n] Anonymous. Publications received. Technometrics, 21(3):394, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267779.

Anonymous:1979:PRd

[Ano79o] Anonymous. Publications received. Technometrics, 21(4):586, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268309.

Anonymous:1979:VI

[Ano79p] Anonymous. Volume information. Technometrics, 21(4):590–593, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268275.

Anonymous:1980:CGJ

[Ano80a] Anonymous. Corrigenda to Gabriel, Jackson and Davis. Technometrics, 22(1):136, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268407. See [GZ79, Jac80, Dav79b].

Anonymous:1980:CLW

[Ano80b] Anonymous. Corrigenda to Leone, White, Narula and Narula. Technometrics, 22(3):452, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268342. See [LNN61, WG79].

Anonymous:1980:CNS

[Ano80c] Anonymous. Corrigendum: “Nonparametric Selection Procedures Applied to State Traffic Fatality Rates”. Technometrics, 22(2):285, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268482. See

[McD79].

Anonymous:1984:CUR

[Ano84] Anonymous. Corrigendum: “The Use of the Rank Transform in Regression”. Technometrics, 26(4):419, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269514. See [IC79].

Antle:1977:DED

[Ant77] Charles E. Antle. Discussion of: “Experimental Design for Sensitivity Testing: The Weibull Model”. Technometrics, 19(4):413, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267880. See

[MHC77].

Azen:1973:MLE

[AR73] S. P. Azen and A. H. Reed. Maximum likelihood estimation of correlation between variates having equal coefficients of variation. Technometrics, 15(3):457–462, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266851.

Anderson:1979:LDB

[AR79] J. A. Anderson and S. C. Richardson. Logistic discrimination and bias correction in maximum likelihood estimation. Technometrics, 21(1):71–78, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268582.

Aroian:1976:ADM

[Aro76] Leo A. Aroian. Applications of the direct method in sequential analysis. Technometrics, 18(3):301–306, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268739.

Alexander:1973:TST

[AS73] Robert Alexander and Ronald Suich. A truncated sequential t-test for general α and β. Technometrics, 15(1):79–86, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266826.

Alam:1973:PRE

[AT73] Khursheed Alam and James R. Thompson. A problem of ranking and estimation with Poisson processes. Technometrics, 15(4):801–808, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267390.

Anderson:1979:NMR

[AT79] Donald A. Anderson and Ann M. Thomas. Near minimal resolution IV designs for the s^{n} factorial experiment. Technometrics, 21(3):331– 336, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267755.

Austin:1973:CCC

[Aus73] John A. Austin, Jr. Control chart constants for largest and smallest in sampling from a normal distribution using the generalized Burr distribution. Technometrics, 15(4):931–933, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267403.

Adriaanse:1970:SRG

[AvdL70] R. P. Adriaanse and P. van der Laan. Some remarks on a general class of Markov processes with discrete time parameter and dependent increments. Technometrics, 12(4):851–866, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267330.

Bobis:1970:AED

[BA70] A. H. Bobis and L. Bryce Andersen. An approach for economic discrimination between alternative chemical syntheses. Technometrics, 12(3):439–455, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267195.

Billmann:1972:SIC

[BAB72] Barry R. Billmann, Charles E. Antle, and Lee J. Bain. Statistical inference from censored Weibull samples. Technometrics, 14 (4):831–840, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267132.

Badenius:1970:QPW

[Bad70] Duncan Badenius. Query 3: In probability, why does A ⊂ B mean, “A Implies B”? Technometrics, 12(3):707–708, August 1970. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267218.

Bagga:1970:SRS

[Bag70] P. C. Bagga. Sequencing with random service times. Technometrics, 12(2):327–334, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267237.

Bain:1972:IBC

[Bai72] Lee J. Bain. Inferences based on censored sampling from the Weibull or extreme-value distribution. Technometrics, 14(3):693–702, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267297.

Bain:1973:ROM

[Bai73] Lee J. Bain. Results for one or more independent censored samples from the Weibull or extreme-value distribution. Technometrics, 15(2):279–288, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266988.

Bailey:1974:NUD

[Bai74a] William A. Bailey. A note on the use of discrete convolutions. Technometrics, 16(3):473–475, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267681.

Bain:1974:ALF

[Bai74b] Lee J. Bain. Analysis for the linear failure-rate life-testing distribution. Technometrics, 16(4):551–559, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267607.

Baker:1974:BRB

[Bak74] Charles R. Baker. Book review: The Analysis of Feedback Systems by Jan C. Willems. Technometrics, 16(1):141–142, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267511.

Banerjee:1973:BRB

[Ban73] K. S. Banerjee. Book review: Generalized Inverse of Matrices and Its Applications by C. Radhakrishna Rao; Sujit Kumar Mitra. Technometrics, 15(1):197, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266840.

Barnett:1972:BSL

[Bar72] V. D. Barnett. A Bayesian sequential life test. Technometrics, 14(2):453–468, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267435.

Banerjee:1979:BRI

[BB79] Asit K. Banerjee and G. K. Bhattacharyya. Bayesian results for the inverse Gaussian distribution with an application. Technometrics, 21(2):247–251, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268523.

Banerjee:1971:CRR

[BC71] K. S. Banerjee and R. N. Carr. A comment on ridge regression. Biased estimation for non-orthogonal problems. Technometrics, 13 (4):895–898, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266967.

Beckman:1979:LE

[BC79a] R. J. Beckman and R. D. Cook. Letter to the Editor. Technometrics, 21(4):587, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268310.

Beckman:1979:TTP

[BC79b] R. J. Beckman and R. D. Cook. Testing for two-phase regressions. Technometrics, 21(1):65–69, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268581.

Bose:1961:MCS

[BCK61] R. C. Bose, I. M. Chakravarti, and D. E. Knuth. On methods of constructing sets of mutually orthogonal Latin squares using a computer. II. Technometrics, 3(1):111–117, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www. jstor.org/stable/1266482. See errata [Ano78e, BCK78].

Bose:1978:EMC

[BCK78] R. C. Bose, I. M. Chakravarti, and D. E. Knuth. Errata: “On methods of constructing sets of mutually orthogonal Latin squares using a computer. II” (Technometrics 3 (1961), 111–117). Technometrics, 20 (2):219, 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). See [BCK61].

Box:1971:FDC

[BD71] M. J. Box and N. R. Draper. Factorial designs, the |X′X| criterion, and some related matters. Technometrics, 13(4):731–742, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266950. See corri-

genda [Ano72e, Ano73h].

Barham:1972:ALS

[BD72a] R. H. Barham and Wanzer Drane. An algorithm for least squares estimation of nonlinear parameters when some of the parameters are linear. Technometrics, 14(3):757–766, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267303.

Behnken:1972:RTV

[BD72b] Donald W. Behnken and Norman R. Draper. Residuals and their variance patterns. Technometrics, 14(1):101–111, February 1972. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266922.

Barr:1973:KST

[BD73] Donald R. Barr and Teddy Davidson. A Kolmogorov-Smirnov test for censored samples. Technometrics, 15(4):739–757, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267385.

Box:1974:MPS

[BD74] Michael J. Box and Norman R. Draper. On minimum-point secondorder designs. Technometrics, 16(4):613–616, November 1974. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267614.

Bishop:1978:EAV

[BD78] Thomas A. Bishop and Edward J. Dudewicz. Exact analysis of variance with unequal variances: Test procedures and tables. Technometrics, 20(4):419–430, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267642.

Box:1970:MVM

[BDH70] Michael J. Box, Norman R. Draper, and William G. Hunter. Missing values in multiresponse nonlinear model fitting. Technometrics, 12(3):613–620, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267206.

Bain:1973:IET

[BE73] Lee J. Bain and Max Engelhardt. Interval estimation for the two-parameter double exponential distribution. Technometrics, 15 (4):875–887, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267397.

Beale:1970:NPV

[Bea70] E. M. L. Beale. Note on procedures for variable selection in multiple regression. Technometrics, 12(4):909–914, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267335.

Beck:1974:PEC

[Bec74] J. V. Beck. Parameter estimation with cumulative errors. Technometrics, 16(1):85–92, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267496.

Bedall:1974:BRB

[Bed74] Fritz K. Bedall. Book review: Statistische Auswertungsmethoden by Lothar Sachs. Technometrics, 16(1):141, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267510.

Behboodian:1970:MMT

[Beh70] Javad Behboodian. On the modes of a mixture of two normal distributions. Technometrics, 12(1):131–139, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267357.

Behboodian:1972:DSS

[Beh72] Javad Behboodian. On the distribution of a symmetric statistic from a mixed population. Technometrics, 14(4):919–923, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267140.

Bell:1970:BRB

[Bel70] C. B. Bell. Book review: Distribution-Free Statistical Tests by James V. Bradley. Technometrics, 12(4):929, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267340.

Bender:1976:BRB

[Ben76] Filmore E. Bender. Book review: Cost of Living Index Numbers: Practice, Precision and Theory by Kali S. Banerjee. Technometrics, 18(2):243–244, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267538.

Berger:1972:YOF

[Ber72a] Paul D. Berger. On Yates’ order in fractional factorial designs. Technometrics, 14(4):971–972, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267146.

Berman:1972:BRB

[Ber72b] Simeon M. Berman. Book review: On Regular Variation and Its Application to the Weak Convergence of Sample Extremes by L. de Haan. Technometrics, 14(1):236–237, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266940.

Berry:1973:BRB

[Ber73] Brian J. L. Berry. Book review: Explorations in Mathematical Anthropology by Paul Kay. Technometrics, 15(1):201–202, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266844.

Berk:1978:CSR

[Ber78] Kenneth N. Berk. Comparing subset regression procedures. Technometrics, 20(1):1–6, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268153.

Brown:1972:SMS

[BF72] G. H. Brown and N. I. Fisher. Subsampling a mixture of sampled material. Technometrics, 14(3):663–668, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267294.

Brown:1974:SSB

[BF74] Morton B. Brown and Alan B. Forsythe. The small sample behavior of some statistics which test the equality of several means. Technometrics, 16(1):129–132, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267501.

Best:1974:IGA

[BG74] D. J. Best and P. G. Gipps. An improved gamma approximation to the negative binomial. Technometrics, 16(4):621–624, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267616.

Bradu:1978:BDT

[BG78] Dan Bradu and K. Ruben Gabriel. The biplot as a diagnostic tool for models of two-way tables. Technometrics, 20(1):47–68, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268161. See errata

[Ano78e].

Blumenthal:1973:TRB

[BGH73] S. Blumenthal, J. A. Greenwood, and L. Herbach. The transient reliability behavior of series systems or superimposed renewal processes. Technometrics, 15(2):255–269, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266986.

Ball:1970:SII

[BH70] Geoffrey H. Ball and David J. Hall. Some implications of interactive graphic computer systems for data analysis and statistics. Technometrics, 12(1):17–31, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267348.

Bulgren:1973:DST

[BH73] William G. Bulgren and John E. Hewett. Double sample tests for hypotheses about the mean of an exponential distribution. Technometrics, 15(1):187–190, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266836.

Box:1974:CIV

[BH74] George E. P. Box and William J. Hill. Correcting inhomogeneity of variance with power transformation weighting. Technometrics, 16(3):385–389, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267668.

Bradley:1977:ODT

[BH77] Ralph A. Bradley and A. Hedayat. Optimal designs for two noninteractive treatments. Technometrics, 19(1):53–57, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268254.

Bauer:1978:UMQ

[BH78] Peter Bauer and Peter Hackl. The use of MOSUMS for quality control. Technometrics, 20(4):431–436, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267643.

Bhapkar:1972:BRB

[Bha72] Vasant Bhapkar. Book review: Practical Nonparametric Statistics by W. J. Conover. Technometrics, 14(4):977–979, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267149.

Bhat:1973:RIC

[Bha73] U. Narayan Bhat. Reliability of an independent component, s-spare system with exponential life times and general repair times. Technometrics, 15(3):529–539, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266858.

Box:1973:SPA

[BHME73] G. E. P. Box, W. G. Hunter, J. F. MacGregor, and J. Erjavec. Some problems associated with the analysis of multiresponse data. Technometrics, 15(1):33–51, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266823.

Binns:1974:ANB

[Bin74a] M. R. Binns. Approximating the negative bionomial via the positive binomial. Technometrics, 16(2):323–324, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267956.

Binns:1974:DSP

[Bin74b] M. R. Binns. A double sampling procedure to control the variance in net weight acceptance sampling. Technometrics, 16(4):539–543, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267605.

Bishir:1972:BRB

[Bis72a] John Bishir. Book review: Physical Concepts of Probability by J. A. McFadden. Technometrics, 14(4):977, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267148.

Bishop:1972:BRB

[Bis72b] Albert Bishop. Book review: Mathematical Methods of Optimal Control by V. G. Boltyanskii; K. N. Trirogoff. Technometrics, 14(4):981–983, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267153.

Bishop:1978:NSR

[Bis78] Thomas A. Bishop. A note on Schafer and Rutemiller’s complete ranking procedure. Technometrics, 20(4):495–496, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267651. See corri-

genda [Ano79e] and [SR75].

Bishop:1979:SRS

[Bis79] Thomas A. Bishop. Some results on simultaneous inference for analysis of variance with unequal variances. Technometrics, 21(3):337–340, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267756.

Basson:1970:LE

[BJ70a] Rodney P. Basson and Jay W. Johnson. Letters to the Editor. Technometrics, 12(1):187–188, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267375.

Bellmore:1970:IES

[BJ70b] M. Bellmore and P. A. Jensen. An implicit enumeration scheme for proper cut generation. Technometrics, 12(4):775–788, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267324.

Bagshaw:1975:ESC

[BJ75] Michael Bagshaw and Richard A. Johnson. The effect of serial correlation on the performance of CUSUM tests II. Technometrics, 17(1):73– 80, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268003.

Bhattacharyya:1971:CPS

[BJN71] G. K. Bhattacharyya, Richard A. Johnson, and Henry R. Neave. A comparative power study of the bivariate rank sum test and T^{2}. Technometrics, 13(1):191–198, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267086.

Boardman:1970:ECE

[BK70a] Thomas J. Boardman and Peter J. Kendell. Estimation in compound exponential failure models. Technometrics, 12(4):891–900, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267333.

Boardman:1970:TEA

[BK70b] Thomas J. Boardman and Peter J. Kendell. Tables of exact and approximate expected values of certain ratios and reciprocals of multinomial random variables useful in life testing. Technometrics, 12(4):901–908, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267334.

Bowman:1974:PPP

[BK74] Kimiko O. Bowman and Marvin A. Kastenbaum. Potential pitfalls of portable power. Technometrics, 16(3):349–352, August 1974. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267664.

Barnett:1978:SPT

[BKRS78] A. Barnett, D. J. Kleitman, D. Rosenbaum, and B. Singer. A statistical procedure for testing nuclear powered cardiac pacemakers. Technometrics, 20(3):221–226, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268128.

Basu:1971:DRT

[BL71] A. P. Basu and Robert H. Lochner. On the distribution of the ratio of two random variables having generalized life distributions. Technometrics, 13(2):281–287, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266790.

Britt:1973:EPN

[BL73] H. I. Britt and R. H. Luecke. The estimation of parameters in nonlinear, implicit models. Technometrics, 15(2):233–247, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266984.

Beja:1974:ESA

[BL74a] Avraham Beja and Shaul P. Ladany. Efficient sampling by artificial attributes. Technometrics, 16(4):601–611, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267613.

Berger:1974:EWP

[BL74b] Roger W. Berger and Kenneth Lawrence. Estimating Weibull parameters by linear and nonlinear regression. Technometrics, 16 (4):617–619, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267615.

Blackwell:1977:ESP

[Bla77] M. T. R. Blackwell. The effect of short production runs on CSP-1. Technometrics, 19(3):259–263, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267695.

Bray:1973:TCA

[BLB73] D. F. Bray, D. A. Lyon, and I. W. Burr. Three class attributes plans in acceptance sampling. Technometrics, 15(3):575–585, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266864.

Blight:1972:MEC

[Bli72] B. J. N. Blight. On the most economical choice of a life testing procedure for exponentially distributed data. Technometrics, 14(3):613– 618, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267289.

Box:1972:LE

[BM72] George E. P. Box and John F. MacGregor. Letter to the Editor. Technometrics, 14(4):985–986, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267156.

Box:1974:ACL

[BM74] George E. P. Box and John F. MacGregor. The analysis of closed-loop dynamic-stochastic systems. Technometrics, 16(3):391–398, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267669.

Box:1976:PEC

[BM76] George E. P. Box and John F. MacGregor. Parameter estimation with closed-loop operating data. Technometrics, 18(4):371–380, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268651. See discus-

sion [Ell76].

Bruvold:1978:SSC

[BM78] Norman T. Bruvold and Richard A. Murphy. Sample sizes for comparison of proportions. Technometrics, 20(4):437–440, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267644.

Beyer:1976:ESD

[BMH76] Jon N. Beyer, Albert H. Moore, and H. Leon Harter. Estimation of the standard deviation of the logistic distribution by the use of selected order statistics. Technometrics, 18(3):313–332, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268741.

Breiman:1977:VKE

[BMP77] Leo Breiman, William Meisel, and Edward Purcell. Variable kernel estimates of multivariate densities. Technometrics, 19(2):135–144, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268623.

Bechhofer:1972:OAO

[BN72] Robert E. Bechhofer and Dominique Jean-Marie Nocturne. Optimal allocation of observations when comparing several treatments with a control. II. 2-sided comparisons. Technometrics, 14(2):423–436, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267433.

Boardman:1973:ECE

[Boa73] Thomas J. Boardman. Estimation in compound exponential failure models—when the data are grouped. Technometrics, 15(2):271–277, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266987.

Boardman:1979:BRBa

[Boa79a] Eileen C. Boardman. Book reviews: Statistics in the Real World by Richard J. Larsen; Donna Fox Stroup; A Programmed Text in Statistics: Books One-Four by J. Hine; G. B. Wetherill; Statistical Methods for Engineers and Scientists: A Students’ Course Book by A. C. Bajpai; I. M. Calus; J. A. Fairley; Statistics: A Self-Teaching Guide by Donald J. Koosis. Technometrics, 21(3):390–391, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267771.

Boardman:1979:BRBb

[Boa79b] Thomas J. Boardman. Book review: Prediction and Improved Estimation in Linear Models by John Bibby; Helge Toutenburg. Technometrics, 21(4):582, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268300.

Bodin:1970:ASR

[Bod70] Lawrence D. Bodin. Approximations to system reliability using a modular decomposition. Technometrics, 12(2):335–344, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267238.

Bonis:1976:BRBa

[Bon76a] Austin J. Bonis. Book review: Comparative Statistical Inference by Vick Barnett. Technometrics, 18(1):119–120, February 1976. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267925.

Bonis:1976:BRBb

[Bon76b] Austin J. Bonis. Book review: Gambling with Truth. An Essay on Induction and the Aims of Science by Isaac Levi. Technometrics, 18 (2):240, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267533.

Bonis:1977:BRB

[Bon77] Austin J. Bonis. Book review: Scientific Strategies in Human Affairs: To Tell the Truth by Irwin D. J. Bross. Technometrics, 19(1):102, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268264.

Borges:1972:QAQ

[Bor72] Rudolf Borges. Query 1: Addition to query 3, In Probability, Why Does A ⊂ B Mean, “A Implies B”? (Volume 13, No. 3, August 1970). Technometrics, 14(1):243, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266943.

Bowerman:1977:BRB

[Bow77] Bruce L. Bowerman. Book review: Probability and Mathematical Statistics by Lester D. Taylor. Technometrics, 19(3):342, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267706.

Bowerman:1978:BRB

[Bow78] Bruce Bowerman. Book review: Introductory Statistics by T. H. Wonnacott; R. J. Wonnacott. Technometrics, 20(4):502–503, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267655.

Box:1970:IPE

[Box70a] M. J. Box. Improved parameter estimation. Technometrics, 12 (2):219–229, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267230.

Box:1970:SEN

[Box70b] M. J. Box. Some experiences with a nonlinear experimental design criterion. Technometrics, 12(3):569–589, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267204.

Box:1971:SED

[Box71] M. J. Box. Simplified experimental design. Technometrics, 13(1):19– 31, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267073.

Boyle:1974:BRB

[Boy74] Carolyn R. Boyle. Book review: Statistical and Mathematical Aspects of Pollution Problems by John W. Pratt. Technometrics, 16(4):636, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267627.

Brown:1973:EML

[BR73] Gerald G. Brown and Herbert C. Rutemiller. The efficiencies of maximum likelihood and minimum variance unbiased estimators of fraction defective in the normal case. Technometrics, 15(4):849–855, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267394.

Brown:1975:EMP

[BR75] G. H. Brown and D. S. Robson. The estimation of mixing proportions by double sample using bulk measurement. Technometrics, 17(1):119–126, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268009.

Bakir:1979:NPP

[BR79] Saad T. Bakir and Marion R. Reynolds, Jr. A nonparametric procedure for process control based on within-group ranking. Technometrics, 21(2):175–183, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268514.

Brillinger:1973:BRB

[Bri73a] David R. Brillinger. Book review: Introduction to Random Differential Equations and Their Applications by S. K. Srinivasan; R. Vasudevan. Technometrics, 15(2):421–422, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267004.

Brillinger:1973:PSE

[Bri73b] David R. Brillinger. A power spectral estimate which is insensitive to transients. Technometrics, 15(3):559–562, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266861.

Britney:1974:RCS

[Bri74] Robert R. Britney. The reliability of complex systems with dependent subsystem failures: An absorbing Markov chain model. Technometrics, 16(2):245–250, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267946.

Brown:1977:CSR

[Bro77] P. J. Brown. Centering and scaling in ridge regression. Technometrics, 19(1):35–36, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268251.

Brown:1979:OCL

[Bro79] G. H. Brown. An optimization criterion for linear inverse estimation. Technometrics, 21(4):575–579, November 1979. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268298.

Bryson:1971:GTP

[Bry71] Maurice C. Bryson. A Gibbs-type phenomenon in the failure rates of parallel systems. Technometrics, 13(3):692–695, August 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267182.

Bryson:1974:HTD

[Bry74] Maurice C. Bryson. Heavy-tailed distributions: Properties and tests. Technometrics, 16(1):61–68, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267493.

Barlow:1971:EAL

[BS71] Richard E. Barlow and Ernest M. Scheuer. Estimation from accelerated life tests. Technometrics, 13(1):145–159, February 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267082. See remark [Rea71].

Bradley:1972:COL

[BS72] E. L. Bradley, Jr. and J. G. Saw. Choosing the optimum lot size for a reordering policy with instantaneous emergency deliveries. Technometrics, 14(4):931–934, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267142.

Bryant:1979:CLM

[BS79] Clifford M. Bryant and Josef Schmee. Confidence limits on MTBF for sequential test plans of MIL-STD 781. Technometrics, 21(1):33–42, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268578.

Bratcher:1971:BPP

[BSH71] T. L. Bratcher, W. R. Schucany, and H. H. Hunt. Bayesian prediction and population size assumptions. Technometrics, 13(3):678–681, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267179.

Beaton:1974:FPS

[BT74] Albert E. Beaton and John W. Tukey. The fitting of power series, meaning polynomials, illustrated on band-spectroscopic data. Technometrics, 16(2):147–185, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1267936. See discussion [Mal74, Mar74].

Buck:1973:BRB

[Buc73] James R. Buck. Book review: Work Design, a Systems Concept by Gerald Nadler. Technometrics, 15(1):200–201, February 1973. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266843.

Burgerhout:1973:CPD

[Bur73] Th. J. Burgerhout. A class of probability distributions with application to aircraft flight path analysis. Technometrics, 15(2):365–378, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266995.

Bush:1971:UEA

[Bus71] Norman Bush. Unmodeled error analysis on trajectory and orbital estimation. Technometrics, 13(2):303–314, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266792.

Butcher:1978:TIM

[BWL78] Allan C. Butcher, Alan Winterbottom, and Harold Lampkin. Transformations improving maximum likelihood confidence intervals for system reliability. Technometrics, 20(4):467–473, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267647.

Cameron:1977:BRB

[Cam77] J. M. Cameron. Book review: Weighing Designs by K. S. Baneriee. Technometrics, 19(2):219–220, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268639.

Carter:1974:BRB

[Car74] Walter H. Carter, Jr. Book review: Fundamental Concepts in the Design of Experiments by Charles Hicks. Technometrics, 16(2):327– 328, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267961.

Carney:1977:BRB

[Car77a] Edward J. Carney. Book review: Vital Statistics by Michael Orkin; Richard Drogin. Technometrics, 19(2):218–219, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268637.

Carroll:1977:BRB

[Car77b] Mavis B. Carroll. Book review: Micro-Organisms in Foods 2: Sampling for Microbiological Analysis; Principles and Specific Applications by The International Commission on Microbiological Specifications for Foods. Technometrics, 19(2):221, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268642.

Castro:1976:BRB

[Cas76] Peter E. Castro. Book review: Complex Variables for Scientists and Engineers by John D. Paliouras. Technometrics, 18(2):244, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267539.

Canfield:1975:SDT

[CB75] Ronald V. Canfield and Leon E. Borgman. Some distributions of time to failure for reliability applications. Technometrics, 17(2):263–268, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268360.

Conover:1979:MDC

[CBI79] W. J. Conover, Thomas R. Bement, and Ronald L. Iman. On a method for detecting clusters of possible uranium deposits. Technometrics, 21(3):277–282, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267748.

Chan:1971:STB

[CC71] Lai K. Chan and Smiley W. H. Cheng. On the Student’s test based on sample percentiles from the normal, logistic and Cauchy distribution. Technometrics, 13(1):127–137, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267080.

Chacko:1979:OME

[CDR79] Annie Chacko, A. Dey, and G. V. S. Ramakrishna. Orthogonal maineffect plans for asymmetrical factorials. Technometrics, 21(2):269– 270, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268528.

Chhikara:1976:OTP

[CF76] Raj S. Chhikara and J. Leroy Folks. Optimum test procedures for the mean of first passage time distribution in Brownian motion with positive drift (inverse Gaussian distribution). Technometrics, 18(2):189– 193, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267521.

Chhikara:1977:IGD

[CF77] R. S. Chhikara and J. L. Folks. The inverse Gaussian distribution as a lifetime model. Technometrics, 19(4):461–468, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267886.

Cleveland:1976:SRS

[CG76] William S. Cleveland and Roberta Guarino. Some robust statistical procedures and their application to air pollution data. Technometrics, 18(4):401–409, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268655.

Chan:2007:MMC

[CGO07] Raymond H. Chan, Chen Greif, and Dianne P. O’Leary, editors. Milestones in Matrix Computation: the Selected Works of Gene H. Golub with Commentaries. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920681-3. xi + 565 + 3

pp. LCCN QA188 .G67 2007. URL http://www.loc.gov/catdir/
enhancements/fy0737/2007276086-d.html.

Church:1970:ERS

[CH70] J. D. Church and Bernard Harris. The estimation of reliability from stress-strength relationships. Technometrics, 12(1):49–54, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267350.

Chase:1973:TET

[CH73a] Gerald R. Chase and John E. Hewett. On testing for equality of two availabilities. Technometrics, 15(4):889–896, November 1973. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267398.

Cohen:1973:EED

[CH73b] A. Clifford Cohen and Frederick Russell Helm. Estimation in the exponential distribution. Technometrics, 15(2):415–418, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267000.

Cleroux:1974:ARA

[CH74] R. Cleroux and M. Hanscom. Age replacement with adjustment and depreciation costs and interest charges. Technometrics, 16(2):235– 239, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267944.

Chambers:1970:CSR

[Cha70a] J. M. Chambers. Computers in statistical research: Simulation and computer-aided mathematics. Technometrics, 12(1):1–15, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267347.

Chapman:1970:MVC

[Cha70b] J. W. Chapman. Moments, variances, and covariances of sines and cosines of arguments which are subject to random error. Technometrics, 12(3):693–694, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267213.

Chandler:1972:IPE

[Cha72] J. P. Chandler. On an iterative procedure for estimating functions when both variables are subject to error. Technometrics, 14(1):71–76, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266919.

Chafee:1975:BRB

[Cha75a] Nathaniel Chafee. Book review: Partial Differential Equations of Mathematical Physics by Tyn Myint-U. Technometrics, 17(3):389, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268089.

Chambers:1975:BRB

[Cha75b] John M. Chambers. Book review: Nonlinear Parameter Estimation by Yonathan Bard. Technometrics, 17(2):271–272, May 1975. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268363.

Chang:1979:CIE

[Cha79] Wei-Chien Chang. Confidence interval estimation and transformation of data in a mixture of two multivariate normal distributions with any given large dimension. Technometrics, 21(3):351–355, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267758.

Chew:1971:BRB

[Che71] Milton C. Chew, Jr. Book review: Distributions in Statistics: Continuous Univariate Distributions — 1 and 2 by N. L. Johnson; S. Kotz. Technometrics, 13(4):919–920, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266973.

Chaddha:1971:ATS

[CHH^{+}71] R. L. Chaddha, W. W. Hardgrave, D. J. Hudson, M. Segal, and J. W. Suurballe. Allocation of total sample size when only the stratum means are of interest. Technometrics, 13(4):817–831, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266957.

Chiu:1974:NPD

[Chi74] W. K. Chiu. A new prior distribution for attributes sampling. Technometrics, 16(1):93–102, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267497.

Chiu:1975:EDA

[Chi75] W. K. Chiu. Economic design of attribute control charts. Technometrics, 17(1):81–87, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268004.

Choi:1973:NST

[Cho73] S. C. Choi. On nonparametric sequential tests for independence. Technometrics, 15(3):625–629, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266868.

Choi:1979:TST

[Cho79] S. C. Choi. Two-sample tests for compound distributions for homogeneity of mixing proportions. Technometrics, 21(3):361–365, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267760.

Consul:1973:GPD

[CJ73] P. C. Consul and G. C. Jain. A generalization of the Poisson distribution. Technometrics, 15(4):791–799, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267389.

Cleveland:1975:GTE

[CK75] William S. Cleveland and Beat Kleiner. A graphical technique for enhancing scatterplots with moving statistics. Technometrics, 17(4):447–454, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268431.

Cornell:1979:OCP

[CK79] John A. Cornell and Andre I. Khuri. Obtaining constant prediction variance on concentric triangles for ternary mixture systems. Technometrics, 21(2):147–157, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268512.

Corneliussen:1970:STB

[CL70] Arvid Corneliussen and Daniel W. Ladd. On sequential tests of the binomial distribution. Technometrics, 12(3):635–646, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267209. See adden-

dum [CL71].

Corneliussen:1971:AST

[CL71] Arvid H. Corneliussen and Daniel W. Ladd. Addendum to “On Sequential Tests of the Binomial Distribution”. Technometrics, 13(1):205–210, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267090. See [CL70].

Cleveland:1971:FTS

[Cle71] William S. Cleveland. Fitting time series models for prediction. Technometrics, 13(4):713–723, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266948.

Cleveland:1972:BRB

[Cle72a] William S. Cleveland. Book review: Multiple Time Series by E. J. Hannan. Technometrics, 14(1):229–231, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266935.

Cleveland:1972:IAT

[Cle72b] William S. Cleveland. The inverse autocorrelations of a time series and their applications. Technometrics, 14(2):277–293, May 1972. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267420. See discussion [Par72].

Connell:1971:CWT

[CM71] Terrence L. Connell and Piotr W. Mikulski. On comparison of waiting time versus counting technique in construction of confidence intervals for a parameter of a Poisson process. Technometrics, 13(3):699–703, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267184.

Cohen:1977:PCS

[CN77] A. Clifford Cohen and Nicholas J. Norgaard. Progressively censored sampling in the three-parameter gamma distribution. Technometrics, 19(3):333–340, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267704.

Cornell:1975:UGA

[CO75] John A. Cornell and Lyman Ott. The use of gradients to aid in the interpretation of mixture response surfaces. Technometrics, 17(4):409–424, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268427. See discussion [Sne75a].

Cohen:1965:MLE

[Coh65] A. Clifford Cohen. Maximum likelihood estimation in the Weibull distribution based on complete and on censored samples. Technometrics, 7(4):579–588, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1266397. See errata [Ano73k].

Cohen:1970:CAS

[Coh70] A. Clifford Cohen. Curtailed attribute sampling. Technometrics, 12(2):295–298, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267234.

Cohen:1972:EPP

[Coh72] A. Clifford Cohen. Estimation in a Poisson process based on combinations of complete and truncated samples. Technometrics, 14 (4):841–846, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267133.

Cohen:1973:RWD

[Coh73] A. Clifford Cohen. The reflected Weibull distribution. Technometrics, 15(4):867–873, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267396.

Cohen:1975:MCS

[Coh75] A. Clifford Cohen. Multi-censored sampling in the three parameter Weibull distribution. Technometrics, 17(3):347–351, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268072.

Cohen:1976:PCS

[Coh76] A. Clifford Cohen. Progressively censored sampling in the three parameter log-normal distribution. Technometrics, 18(1):99–103, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267922.

Company:1970:AR

[Com70] Alexander Grant & Company. Auditor’s report. Technometrics, 12(2):435–437, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267254.

Company:1971:AR

[Com71] Alexander Grant & Company. Auditors’ report. Technometrics, 13 (2):461, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266816.

Conniffe:1970:NRR

[Con70] D. Conniffe. A note on the ’round robin’ design. Technometrics, 12(3):695–696, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267214.

Conover:1974:BRB

[Con74] W. J. Conover. Book review: Statistical Methods for Rates and Proportions by Joseph L. Fleiss. Technometrics, 16(2):326–327, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267960.

Cooper:1972:BRB

[Coo72] Leon Cooper. Book review: Notes on Linear Programming by M. Sakarovitch. Technometrics, 14(1):237–238, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266941.

Cook:1977:DIO

[Coo77] R. Dennis Cook. Detection of influential observation in linear regression. Technometrics, 19(1):15–18, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268249.

Cornell:1971:BRB

[Cor71] John A. Cornell. Book review: Introductory Mathematical Statistics: Principles and Methods by Erwin Kreyszig. Technometrics, 13(4):922–925, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266977.

Cornell:1973:EMR

[Cor73] John A. Cornell. Experiments with mixtures: A review. Technometrics, 15(3):437–455, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266850.

Cornell:1975:SCD

[Cor75] John A. Cornell. Some comments on designs for Cox’s mixture polynomial. Technometrics, 17(1):25–35, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267996.

Cornell:1977:WVU

[Cor77] John A. Cornell. Weighted versus unweighted estimates using Scheffé’s mixture model for symmetrical error variances patterns. Technometrics, 19(3):237–247, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267693.

Cornell:1979:EMU

[Cor79] John A. Cornell. Experiments with mixtures: An update and bibliography. Technometrics, 21(1):95–106, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268585.

Cleveland:1975:ECF

[CP75] William S. Cleveland and Emanuel Parzen. The estimation of coherence, frequency response, and envelope delay. Technometrics, 17(2):167–172, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268347.

Chhikara:1979:NCM

[CR79] R. S. Chhikara and D. T. Register. A numerical classification method for partitioning of a large multidimensional mixed data set. Technometrics, 21(4):531–537, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268293.

Crain:1974:BRB

[Cra74] Bradford R. Crain. Book review: Approximating Countable Markov Chains, Volume III by David Freedman. Technometrics, 16(4):635, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267624.

Crocker:1971:CNJ

[Cro71] Douglas C. Crocker. Comment on a note by John J. Wiorkowski. Technometrics, 13(2):455, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266814. See [Wio70].

Crocker:1978:BRB

[Cro78] Douglas C. Crocker. Book review: A Short Course on Applied Regression Analysis by Stanley L. Sclove. Technometrics, 20(2):215, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268727.

Crump:1978:BRBa

[Cru78a] Phelps P. Crump. Book review: Introduction to Statistics by Robert Fried. Technometrics, 20(1):103–104, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268166.

Crump:1978:BRBb

[Cru78b] Phelps P. Crump. Book review: Statistics and Experimental Design in Engineering and the Physical Sciences by Norman L. Johnson; Fred C. Leone. Technometrics, 20(3):319–320, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268143.

Corbeil:1976:RML

[CS76] R. R. Corbeil and S. R. Searle. Restricted maximum likelihood (REML) estimation of variance components in the mixed model. Technometrics, 18(1):31–38, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267913.

DAgostino:1971:ALE

[D’A71a] Ralph B. D’Agostino. An addendum to “linear estimation of the Weibull parameters”. Technometrics, 13(2):428–429, May 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266804.

DAgostino:1971:LEW

[D’A71b] Ralph B. D’Agostino. Linear estimation of the Weibull parameters. Technometrics, 13(1):171–182, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267084.

Dumonceaux:1973:DBL

[DA73] Robert Dumonceaux and Charles E. Antle. Discrimination between the log-normal and the Weibull distributions. Technometrics, 15(4):923–926, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267401.

Dumonceaux:1973:LRT

[DAH73] Robert Dumonceaux, Charles E. Antle, and Gerald Haas. Likelihood ratio test for discrimination between two models with unknown location and scale parameters. Technometrics, 15(1):19–27, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266821. See discus-

sion [Har73a].

Dalal:1977:BRB

[Dal77] Jayesh G. Dalal. Book review: Statistics for Technology by Christopher Chatfield. Technometrics, 19(3):342, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267707.

Danziger:1970:PCL

[Dan70] L. Danziger. Planning censored life tests for estimation of the hazard rate of a Weibull distribution with prescribed precision. Technometrics, 12(2):408–412, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267246.

Daniel:1978:PRT

[Dan78] Cuthbert Daniel. Patterns in residuals in the two-way layout. Technometrics, 20(4):385–395, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267638.

David:1970:BRB

[Dav70] F. N. David. Book review: Systems of Frequency Curves by W. P. Elderton; N. L. Johnson. Technometrics, 12(1):173–174, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267366.

Davis:1977:BRB

[Dav77] Henry T. Davis. Book review: Statistical and Computational Methods in Data Analysis by Siegmund Brandt. Technometrics, 19 (4):516–517, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267898.

Davis:1979:CRS

[Dav79a] A. W. Davis. On certain ratio statistics in weather modification experiments. Technometrics, 21(3):283–289, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267749.

Davis:1979:RMD

[Dav79b] William W. Davis. Robust methods for detection of shifts of the innovation variance of a time series. Technometrics, 21(3):313–320, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267753.

See [Ano80a].

DeGray:1970:SYS

[DeG70] R. J. DeGray. The sequential Youden square. Technometrics, 12(1):147–152, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267359.

Draper:1971:BEB

[DG71] N. Draper and I. Guttman. Bayesian estimation of the binomial parameter. Technometrics, 13(3):667–673, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267177.

Dahiya:1972:GFT

[DG72] Ram C. Dahiya and John Gurland. Goodness of fit tests for the gamma and exponential distributions. Technometrics, 14(3):791–801, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267306.

Draper:1971:LF

[DH71] Norman R. Draper and Agnes M. Herzberg. On lack of fit. Technometrics, 13(2):231–241, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1266786. See corrigendum [Ano72f].

Diehr:1974:ADS

[DH74] George Diehr and Donald R. Hoflin. Approximating the distribution of the sample R^{2} in best subset regressions. Technometrics, 16(2):317–320, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267954.

Dickinson:1974:SRO

[Dic74] A. W. Dickinson. Some run orders requiring a minimum number of factor level changes for the 2^{4} and 2^{5} main effect plans. Technometrics, 16(1):31–37, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267489.

Dillon:1971:BRB

[Dil71] Martin Dillon. Book review: Style and Vocabulary: Numerical Studies by C. B. Williams. Technometrics, 13(3):708–709, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267188.

Draper:1970:NRT

[DL70] Norman R. Draper and Willard E. Lawrence. A note on residuals in two dimensions. Technometrics, 12(2):394–398, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267244.

Denby:1977:TDD

[DM77] L. Denby and C. L. Mallows. Two diagnostic displays for robust regression analysis. Technometrics, 19(1):1–13, February 1977. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268248.

Dufour:1978:DRM

[DM78] R. Dufour and U. R. Maag. Distribution results for modified Kolmogorov–Smirnov statistics for truncated or censored samples. Technometrics, 20(1):29–32, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268157.

Dolby:1970:SSA

[Dol70] James L. Dolby. Some statistical aspects of character recognition. Technometrics, 12(2):231–245, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267231.

Doob:1972:BRB

[Doo72] J. L. Doob. Book review: Markov Chains by David Freedman. Technometrics, 14(2):503, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267441.

Downton:1973:ENC

[Dow73] F. Downton. The estimation of Pr(Y < X) in the normal case. Technometrics, 15(3):551–558, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266860.

Dey:1977:NOM

[DR77] A. Dey and G. V. S. Ramakrishna. A note on orthogonal maineffect plans. Technometrics, 19(4):511–512, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267894.

Drane:1979:BRB

[Dra79] J. Wanzer Drane. Book review: The Analysis of Cross-Classified Categorical Data by Stephen E. Fienberg. Technometrics, 21(3):388, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267767.

Draper:1977:DTF

[DS77a] Norman R. Draper and Ralph C. St. John. Designs in three and four components for mixtures models with inverse terms. Technometrics, 19(2):117–130, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268621.

Draper:1977:MMI

[DS77b] Norman R. Draper and Ralph C. St. John. A mixtures model with inverse terms. Technometrics, 19(1):37–46, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268252.

Doviak:1979:SSE

[DS79] M. J. Doviak and R. L. Scheaffer. A sequential scheme for estimating average quality in a mixed product production process. Technometrics, 21(4):549–554, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268295.

Dubey:1973:PPT

[Dub73] Satya D. Dubey. On the performance of power of tests for the comparison of exponential distributions. Technometrics, 15(1):183–186, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266835.

Dudewicz:1979:BRB

[Dud79] Edward J. Dudewicz. Book review: Selecting and Ordering Populations: A New Statistical Methodology by J. D. Gibbons; I. Olkin; M. Sobel. Technometrics, 21(4):582–583, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268301.

Dunn:1971:SEV

[Dun71] Olive Jean Dunn. Some expected values for probabilities of correct classification in discriminant analysis. Technometrics, 13(2):345–353, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266796.

Dunn:1972:BRB

[Dun72] Olive Jean Dunn. Book review: Nonparametric Methods in Multivariate Analysis by M. L. Puri; P. K. Sen. Technometrics, 14(2):502–503, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267440.

Dunsmore:1974:BPD

[Dun74] I. R. Dunsmore. The Bayesian predictive distribution in life testing models. Technometrics, 16(3):455–460, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267677.

Dunn:1977:BRBa

[Dun77a] Douglas M. Dunn. Book review: Forecasting and Time Series Analysis by Douglas C. Montgomery; Lynwood A. Johnson. Technometrics, 19(3):344, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267711.

Dunn:1977:BRBb

[Dun77b] Douglas M. Dunn. Book review: Management Applications of Modern Control Theory by A. Bensoussan; E. Gerald Hurst,; B. Naslund. Technometrics, 19(3):344–345, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267712.

Duncan:1978:EDC

[Dun78a] Acheson J. Duncan. The economic design of p-charts to maintain current control of a process: Some numerical results. Technometrics, 20(3):235–243, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268131.

Duncan:1978:ESJ

[Dun78b] George T. Duncan. An empirical study of jackknife-constructed confidence regions in nonlinear regression. Technometrics, 20(2):123–129, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268703.

Dunsmore:1978:SAT

[Dun78c] Ian R. Dunsmore. Some approximations for tolerance factors for the two parameter exponential distribution. Technometrics, 20(3):317– 318, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268141.

Davis:1978:ACI

[DV78] A. W. Davis and L. G. Veitch. Approximate confidence intervals for nonlinear functions, with applications to a cloud-seeding experiment. Technometrics, 20(3):227–230, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268129.

Draper:1979:RRJ

[DV79] Norman R. Draper and R. Craig Van Nostrand. Ridge regression and James-Stein estimation: review and comments. Technometrics, 21(4):451–466, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268284.

Davenport:1972:TEP

[DW72] James M. Davenport and J. T. Webster. Type-I error and power of a test involving a Satterthwaite’s approximate F -statistic. Technometrics, 14(3):555–569, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267284.

Davenport:1973:CSA

[DW73] James M. Davenport and J. T. Webster. A comparison of some approximate F -tests. Technometrics, 15(4):779–789, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267388.

Dyer:1970:NRS

[DY70] M. R. Dyer and D. H. Young. A note on the reliability of a system with spares which operates at discrete times. Technometrics, 12(3):702–705, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267217.

Dyer:1973:ESP

[Dye73] D. D. Dyer. Estimation of the scale parameter of the chi distribution based on sample quantiles. Technometrics, 15(3):489–496, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266854.

Dyer:1974:BRB

[Dye74] Alan R. Dyer. Book review: Introduction to Statistics by Ronald E. Walpole. Technometrics, 16(4):633–634, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267620.

Dykstra:1971:AED

[Dyk71] Otto Dykstra, Jr. The augmentation of experimental data to maximize |X′X|. Technometrics, 13(3):682–688, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267180.

Dykstra:1976:BRB

[Dyk76] Otto Dykstra, Jr. Book review: Statistics: An Introduction by Harold J. Larson. Technometrics, 18(4):502, November 1976. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268670.

Easterling:1972:BRB

[Eas72] Robert G. Easterling. Book review: Reliability Mathematics: Fundamentals; Practices; Procedures by Bertram L. Amstadter. Technometrics, 14(1):232–234, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266936.

Easterling:1976:GFP

[Eas76] Robert G. Easterling. Goodness of fit and parameter estimation. Technometrics, 18(1):1–9, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267910.

Easterling:1978:DCI

[Eas78] Robert G. Easterling. Discussion of: “Confidence Interval Estimation for the Weibull and Extreme Value Distributions”. Technometrics, 20(4):365–368, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267634. See [Law78].

Engelhardt:1973:SCC

[EB73] Max Engelhardt and Lee J. Bain. Some complete and censored sampling results for the Weibull or extreme-value distribution. Technometrics, 15(3):541–549, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266859.

Engelhardt:1974:SRP

[EB74] Max Engelhardt and Lee J. Bain. Some results on point estimation for the two-parameter Weibull or extreme-value distribution. Technometrics, 16(1):49–56, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267491.

Engelhardt:1975:TTP

[EB75] Max Engelhardt and Lee J. Bain. Tests of two-parameter exponentiality against three-parameter Weibull alternatives. Technometrics, 17(3):353–356, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268073.

Engelhardt:1977:SSP

[EB77a] Max Engelhardt and Lee J. Bain. Simplified statistical procedures for the Weibull or extreme-value distribution. Technometrics, 19(3):323– 331, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267703.

Engelhardt:1977:UMP

[EB77b] Max Engelhardt and Lee J. Bain. Uniformly most powerful unbiased tests on the scale parameter of a gamma distribution with a nuisance shape parameter. Technometrics, 19(1):77–81, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268258.

Engelhardt:1978:COU

[EB78a] Max Engelhardt and Lee J. Bain. Construction of optimal unbiased inference procedures for the parameters of the gamma distribution. Technometrics, 20(4):485–489, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267649.

Engelhardt:1978:PIW

[EB78b] Max Engelhardt and Lee J. Bain. Prediction intervals for the Weibull process. Technometrics, 20(2):167–169, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268709.

Engelhardt:1978:TLC

[EB78c] Max Engelhardt and Lee J. Bain. Tolerance limits and confidence limits on reliability for the two-parameter exponential distribution. Technometrics, 20(1):37–39, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268159.

Engelhardt:1979:PLT

[EB79] Max Engelhardt and Lee J. Bain. Prediction limits and two-sample problems with complete or censored Weibull data. Technometrics, 21(2):233–237, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268521.

Edelman:1974:TSN

[Ede74] David A. Edelman. Three-stage nested designs with composited samples. Technometrics, 16(3):409–417, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267671.

Ehrenfeld:1972:GSS

[Ehr72] S. Ehrenfeld. On group sequential sampling. Technometrics, 14 (1):167–174, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266928.

Ehrenfeld:1977:BRB

[Ehr77] Sylrain Ehrenfeld. Book review: Sequential Methods in Statistics by G. B. Wetherill. Technometrics, 19(2):220, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268640. See addendum and corri-

gendum [Ano77a].

Elder:1974:DSL

[Eld74] Robert S. Elder. Double sampling for lot average. Technometrics, 16(3):435–439, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267674.

Ellingsen:1976:DPE

[Ell76] Walter R. Ellingsen. Discussion of: “Parameter Estimation with Closed-Loop Operating Data”. Technometrics, 18(4):381–384, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268652.

See [BM76].

Ellerton:1978:REA

[Ell78] Roger R. W. Ellerton. Is the regression equation adequate?—a generalization. Technometrics, 20(3):313–315, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268140.

Engelhardt:1975:SEP

[Eng75] Max Engelhardt. On simple estimation of the parameters of the Weibull or extreme-value distribution. Technometrics, 17(3):369–374, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268076.

Enslein:1976:BRB

[Ens76] Kurt Enslein. Book review: Cluster Analysis Algorithms for Classification of Objects and Data Reduction by Helmuth Späth. Technometrics, 18(2):240–241, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267534.

Enslein:1977:BRB

[Ens77] Kurt Enslein. Book review: Einfuhrung in die Mathematische Statistik II, Anwendungen by D. Rasch. Technometrics, 19(4):517, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267899.

Enslein:1978:BRB

[Ens78] Kurt Enslein. Book review: The Natural History of Chronic Bronchitis and Emphysema: An Eight Year Study of Early Chronic Obstructive Lung Disease in Working Men in London by Charles Fletcher; Richard Peto; Cecily Tinker; Frank E. Speizer. Technometrics, 20(2):212–213, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268722.

Easterling:1971:CCS

[EP71] R. G. Easterling and R. R. Prairie. Combining component and system information. Technometrics, 13(2):271–280, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266789.

Ericson:1972:BRB

[Eri72] W. A. Ericson. Book review: Introduction to Mathematical Statistics by Paul G. Hoel. Technometrics, 14(3):807–809, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267313.

Evans:1971:ANI

[Eva71] David H. Evans. An application of numerical integration techniques to statistical tolerancing, II: A note on the error. Technometrics, 13(2):315–324, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266793.

Evans:1972:ANI

[Eva72] David H. Evans. An application of numerical integration techniques to statistical tolerancing, III: General distributions. Technometrics, 14(1):23–35, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266915.

Evans:1979:CAE

[Eva79] James W. Evans. Computer augmentation of experimental designs to maximize |X′X|. Technometrics, 21(3):321–330, August 1979. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267754.

Falls:1970:EPC

[Fal70] Lee W. Falls. Estimation of parameters in compound Weibull distributions. Technometrics, 12(2):399–407, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267245.

Farebrother:1975:MMS

[Far75] R. W. Farebrother. The minimum mean square error linear estimator and ridge regression. Technometrics, 17(1):127–128, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268010.

Farebrother:1978:PRR

[Far78] R. W. Farebrother. Partitioned ridge regression. Technometrics, 20(2):121–122, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268702.

Faulkenberry:1970:NTL

[Fau70] G. David Faulkenberry. A note on tolerance limits for a binomial distribution. Technometrics, 12(4):920–922, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267337.

Folks:1975:IOC

[FB75] J. Leroy Folks and Richard H. Browne. On the interpretation of the observed confidence in certain reliability assessments. Technometrics, 17(3):287–290, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268062.

Faulkenberry:1970:SST

[FD70] G. David Faulkenberry and James C. Daly. Sample size for tolerance limits on a normal distribution. Technometrics, 12(4):813–821, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267327.

Feder:1974:BRB

[Fed74a] Paul I. Feder. Book review: Probability and Statistics for Engineers and Scientists by R. E. Walpole; R. H. Myers. Technometrics, 16(2): 325, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267957.

Feder:1974:GTS

[Fed74b] Paul I. Feder. Graphical techniques in statistical data analysis: Tools for extracting information from data. Technometrics, 16(2):287–299, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267951.

Feder:1975:SRG

[Fed75] Paul I. Feder. Studentized range graph paper: A graphical tool for the comparison of treatment means. Technometrics, 17(2):181–188, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268350.

Feder:1977:BRB

[Fed77] Paul I. Feder. Book review: The Analysis of Categorical Data by R. L. Plackett. Technometrics, 19(4):515–516, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267896.

Fung:1979:DSC

[FH79] Conrad A. Fung and William G. Hunter. Discussion of: “Statistical Comparison of Multiple Analytic Procedures: Application to Clinical Chemistry”. Technometrics, 21(4):411–416, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268279. See [LSYF79].

Fienberg:1973:BRB

[Fie73] Stephen E. Fienberg. Book review: Statistical Papers of George Udny Yule (1871–1951) by Alan Stuart; Maurice G. Kendall; George Udny Yule. Technometrics, 15(3):647–648, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266873.

Filliben:1975:PPC

[Fil75] James J. Filliben. The probability plot correlation coefficient test for normality. Technometrics, 17(1):111–117, February 1975. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268008. See corrigendum

[Ano75e].

Fink:1975:BRB

[Fin75] James P. Fink. Book review: Elementary Differential Equations by Earl D. Rainville; Phillip E. Bedient. Technometrics, 17(2):278–279, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268372.

Finkelstein:1976:CBP

[Fin76] J. M. Finkelstein. Confidence bounds on the parameters of the Weibull process. Technometrics, 18(1):115–117, February 1976. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267924.

Frawley:1971:TLB

[FKRO71] W. H. Frawley, C. H. Kapadia, J. N. K. Rao, and D. B. Owen. Tolerance limits based on range and mean range. Technometrics, 13(3):651–656, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267175.

Fleiss:1971:LE

[Fle71] Joseph L. Fleiss. Letter to the Editor. Technometrics, 13(1):217, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267097.

Flehinger:1976:BRB

[Fle76] B. J. Flehinger. Book review: Sequential Medical Trials by P. Armitage. Technometrics, 18(3):362, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268748.

Feder:1971:AVM

[FM71] P. I. Feder and Reiji Mezaki. An application of variational methods to experimental design. Technometrics, 13(4):771–793, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266954.

Fertig:1977:OSP

[FM77] Kenneth W. Fertig and Nancy R. Mann. One-sided prediction intervals for at least p out of m future observations from a normal population. Technometrics, 19(2):167–177, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268626.

Forthofer:1971:BRB

[For71] Ron Forthofer. Book review: Regression and Econometric Methods by David S. Huang. Technometrics, 13(2):449, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266807.

Forsythe:1972:RES

[For72] Alan B. Forsythe. Robust estimation of straight line regression coefficients by minimizing p th power deviations. Technometrics, 14(1):159–166, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266927.

Forthofer:1973:BRB

[For73] R. Forthofer. Book review: Linear Programming and Related Techniques by C. van de Panne. Technometrics, 15(3):653, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266878.

Farewell:1977:SDS

[FP77] Vern T. Farewell and Ross L. Prentice. A study of distributional shape in life testing. Technometrics, 19(1):69–75, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268257.

Fercho:1972:SSP

[FR72] W. W. Fercho and L. J. Ringer. Small sample power of some tests of the constant failure rate. Technometrics, 14(3):713–724, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267299.

Frane:1977:NCT

[Fra77] James W. Frane. A note on checking tolerance in matrix inversion and regression. Technometrics, 19(4):513–514, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267895.

Freeman:1973:BRBa

[Fre73a] Jean L. Freeman. Book review: Selected Readings in Econometrics from Econometrica by John W. Hooper; Marc Nerlove. Technometrics, 15(1):203, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266846.

Freeman:1973:BRBb

[Fre73b] Jean L. Freeman. Book review: Sixth Berkeley Symposium on Mathematical Statistics and Probability, Volume I by Lucien Le Cam; Jerzy Neyman; Elizabeth L. Scott. Technometrics, 15(2):425–426, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267010.

Freeman:1973:BRBc

[Fre73c] Jean L. Freeman. Book review: Stochastic Point Processes: Statistical Analysis, Theory, and Applications by Peter A. W. Lewis. Technometrics, 15(4):943–944, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267411.

Freeman:1974:BRBb

[Fre74a] Daniel H. Freeman, Jr. Book review: Distributions in Statistics: Continuous Multivariate Distributions by Norman L. Johnson; Samuel Kotz. Technometrics, 16(4):636, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267626.

Freeman:1974:BRBa

[Fre74b] Daniel H. Freeman, Jr. Book review: Statistical Decomposition Analysis by Henri Theil. Technometrics, 16(2):328, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267962.

Freeman:1975:BRB

[Fre75] Daniel H. Freeman, Jr. Book review: Regression Estimation from Grouped Observations by Y. Haitovsky. Technometrics, 17(1):142, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268021.

Fuchs:1978:TSL

[Fuc78] Camil Fuchs. On test sizes in linear models for transformed variables. Technometrics, 20(3):291–299, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268138.

Furnival:1971:APR

[Fur71] George M. Furnival. All possible regressions with less computation. Technometrics, 13(2):403–408, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266801.

Fraser:1972:BRS

[FW72] D. A. S. Fraser and J. S. Williams. On the Technometrics review of “The Structure of Inference”. Technometrics, 14(3):809–810, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267314. See [Wil70].

Furnival:1974:RLB

[FW74] George M. Furnival and Robert W. Wilson, Jr. Regressions by leaps and bounds. Technometrics, 16(4):499–511, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267601.

Gafarian:1978:CBM

[Gaf78] A. V. Gafarian. Confidence bands in multivariate polynomial regression. Technometrics, 20(2):141–149, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268705.

Gart:1971:OCT

[Gar71] John J. Gart. On the ordering of contingency tables for significance tests. Technometrics, 13(4):910–911, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266971.

Gardiner:1974:NE

[Gar74] Donald A. Gardiner. New Editor. Technometrics, 16(2):145, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267935.

Grubbs:1972:ESS

[GB72] Frank E. Grubbs and Glenn Beck. Extension of sample sizes and percentage points for significance tests of outlying observations. Technometrics, 14(4):847–854, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1267134. See corrigendum [Ano73g].

Greenstreet:1974:PTE

[GC74] Richard L. Greenstreet and Robert J. Connor. Power of tests for equality of covariance matrices. Technometrics, 16(1):27–30, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267488.

Gaver:1979:JEC

[GC79] Donald P. Gaver and Boyer B. Chu. Jackknife estimates of component and system availability. Technometrics, 21(4):443–450, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268283.

Guttman:1978:CHU

[GDF78] Irwin Guttman, Rudolf Dutter, and Peter R. Freeman. Care and handling of univariate outliers in the general linear model to detect spuriosity—a Bayesian approach. Technometrics, 20(2):187–193, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268712.

Gentleman:1977:BRB

[Gen77] Jane F. Gentleman. Book review: Data Reduction. Analysing and Interpreting Statistical Data by A. S. C. Ehrenberg. Technometrics, 19(4):516, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267897.

George:1972:BRB

[Geo72] Stephen L. George. Book review: A First Course in Statistics by F. N. David. Technometrics, 14(4):983–984, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267155.

Gerig:1975:BRB

[Ger75] Thomas M. Gerig. Book review: Multivariate Analysis: Techniques for Educational and Psychological Research by Maurice M. Tatsuoka. Technometrics, 17(1):141–142, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268020.

Gnanadesikan:1970:SPM

[GG70a] Mrudulla Gnanadesikan and S. S. Gupta. A selection procedure for multivariate normal distributions in terms of the generalized variances. Technometrics, 12(1):103–117, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267355.

Goldsmith:1970:ACD

[GG70b] Charles H. Goldsmith and David W. Gaylor. Three stage nested designs for estimating variance components. Technometrics, 12(3):487– 498, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267198.

ASQC Chemical Division Technical Conference 1968 Prize Winning Paper.

Gaver:1970:CCS

[GH70] Donald P. Gaver, Jr. and David G. Hoel. Comparison of certain smallsample Poisson probability estimates. Technometrics, 12(4):835–850, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267329.

Guenther:1975:LE

[GH75] William C. Guenther and M. A. Hamdan. Letters to the Editor: “An improved gamma approximation to the negative binomial” (Technometrics 16 (1974), 621–624) by D. J. Best and P. G. Gipps and “A simple approximation to the negative binomial (and regular binomial)” (ibid. 14 (1972), no. 2, 385–389) by W. C. Guenther. Technometrics, 17(4):518–519, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268449.

Goodnight:1979:SAT

[GH79] J. H. Goodnight and W. J. Hemmerle. A simplified algorithm for the W transformation in variance component estimation. Technometrics, 21(2):265–267, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268527.

Golub:1979:GCV

[GHW79] Gene H. Golub, Michael T. Heath, and Grace Wahba. Generalized cross-validation as a method for choosing a good ridge parameter. Technometrics, 21(2):215–223, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1268518. Cited in Åke Björck’s bibliography on least

squares, which is available at ftp://math.liu.se/pub/references.

Reprinted in [CGO07].

Gibbons:1974:BRB

[Gib74] Jean D. Gibbons. Book review: Nonparametric Statistical Methods by M. Hollander; D. A. Wolfe. Technometrics, 16(3):477–478, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267682.

Giesbrecht:1974:BRB

[Gie74] Francis Giesbrecht. Book review: GPSS Primer by Stanley Greenberg. Technometrics, 16(1):142, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267512.

Gillings:1974:BRB

[Gil74] Dennis Gillings. Book review: Introduction to Stochastic Processes by Paul G. Hoel; Sidney C. Port; Charles J. Stone. Technometrics, 16(1):137–138, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267505.

Galil:1977:CBD

[GK77a] Z. Galil and J. Kiefer. Comparison of Box–Draper and Doptimum designs for experiments with mixtures. Technometrics, 19(4):441–444, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267883.

Galil:1977:CSD

[GK77b] Z. Galil and J. Kiefer. Comparison of simplex designs for quadratic mixture models. Technometrics, 19(4):445–453, November 1977. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267884.

Gentle:1975:DSB

[GKS75] James E. Gentle, Ralph L. Kodell, and Patricia L. Smith. On the distribution of the Studentized bivariate range. Technometrics, 17(4):501–505, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268439.

Glenn:1977:BRB

[Gle77] William A. Glenn. Book review: Statistics in Research by Bernard Ostle; Richard W. Mensing. Technometrics, 19(1):103, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268266.

Gleason:1978:BRB

[Gle78] Terry C. Gleason. Book review: Survey and Opinion Research: Procedures for Processing and Analysis by John A. Sonquist; William C. Dunkelberg. Technometrics, 20(4):502, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267654.

Goldman:1974:RPP

[GLV74] Aaron S. Goldman, Homer D. Lewis, and William M. Visscher. Random packing of particles: Simulation with a one dimensional parking model. Technometrics, 16(2):301–309, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267952.

Gunst:1979:SCE

[GM79] Richard F. Gunst and Robert L. Mason. Some considerations in the evaluation of alternate prediction equations. Technometrics, 21(1):55–63, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268580.

Griffiths:1975:LE

[GMD^{+}75] David Griffiths, Alan Miller, Norman R. Draper, Michael J. Box, and Edward R. Jones. Letters to the Editor. Technometrics, 17(2):281–282, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268373.

Goldsmith:1974:BRB

[Gol74] C. H. Goldsmith. Book review: Probability and Statistical Inference by J. K. Wani. Technometrics, 16(1):140–141, February 1974. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267509.

Good:1970:ICM

[Goo70] I. J. Good. The inverse of a centrosymmetric matrix. Technometrics, 12(4):925–928, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267339.

Goodman:1971:AMC

[Goo71] Leo A. Goodman. The analysis of multidimensional contingency tables: Stepwise procedures and direct estimation methods for building models for multiple classifications. Technometrics, 13(1):33–61, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267074.

Good:1975:BRB

[Goo75] I. J. Good. Book review: An Introduction to Bayesian Inference in Econometrics by Arnold Zellner. Technometrics, 17(1):137–138, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268014.

See amendment [Goo76].

Good:1976:ARA

[Goo76] I. J. Good. Amendment to review of Arnold Zellner’s An Introduction to Bayesian Inference in Econometrics. Technometrics, 18(1):123, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267931.

See [Goo75].

Gorman:1976:BRB

[Gor76] John W. Gorman. Book review: Optimal Subset Selection-Multiple Regression, Interdependence and Optimal Network Algorithms by D. E. Boyce; A. Farhi; R. Weischedel. Technometrics, 18(2):239–240, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267530.

Gould:1973:BRB

[Gou73] A. Lawrence Gould. Book review: Probit Analysis by D. J. Finney. Technometrics, 15(1):197–199, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266841.

Gower:1972:BRB

[Gow72] J. C. Gower. Book review: Generalized Inverse Matrices by T. L. Boullion; P. L. Odell. Technometrics, 14(3):806–807, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267311.

Guttman:1973:LSE

[GPS73] Irwin Guttman, Victor Pereyra, and Hugo D. Scolnik. Least squares estimation for a class of non-linear models. Technometrics, 15(2):209– 218, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266982.

Guenther:1976:OSC

[GPU76] William C. Guenther, S. A. Patil, and V. R. R. Uppuluri. Onesided β-content tolerance factors for the two-parameter exponential distribution. Technometrics, 18(3):333–340, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268742.

Graff:1972:GTP

[GR72] Lois E. Graff and Robert Roeloffs. Group testing in the presence of test error; an extension of the Dorfman procedure. Technometrics, 14(1):113–122, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266923.

Grandage:1972:BRB

[Gra72] A. Grandage. Book review: Introductory Engineering Statistics by Irwin Guttman; S. S. Wilks; J. Stuart Hunter. Technometrics, 14(4):979, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267150.

Greenberg:1970:ALP

[Gre70a] Harvey J. Greenberg. An application of a Lagrangian penalty function to obtain optimal redundancy. Technometrics, 12(3):545–552, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267201.

Greenberg:1970:FOS

[Gre70b] Irwin Greenberg. The first occurrence of n successes in N trials. Technometrics, 12(3):627–634, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267208.

Green:1971:TDR

[Gre71] J. R. Green. Testing departure from a regression, without using replication. Technometrics, 13(3):609–615, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267172.

Gridgeman:1970:CTM

[Gri70] N. T. Gridgeman. A comparison of two methods of analysis of mixtures of normal distributions. Technometrics, 12(4):823–833, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267328.

Griffeath:1972:CSD

[Gri72] David S. Griffeath. Computer solution of the discrete maximum entropy problem. Technometrics, 14(4):891–897, November 1972. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267137.

Grizzle:1973:BRB

[Gri73] James E. Grizzle. Book review: Analysis and Design of Certain Quantitative Multiresponse Experiments by S. N. Roy; R. Gnanadesikan; J. N. Srivastava. Technometrics, 15(1):196–197, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266839.

Griffiths:1979:SIC

[Gri79] David Griffiths. Structural identifiability for compartmental models. Technometrics, 21(2):257–259, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268525.

Gross:1971:AES

[Gro71a] Alan J. Gross. The application of exponential smoothing to reliability assessment. Technometrics, 13(4):877–883, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266964.

Gross:1971:MPM

[Gro71b] Alan J. Gross. Monotonicity properties of the moments of truncated gamma and Weibull density functions. Technometrics, 13(4):851–857, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266960.

Grosh:1972:BSA

[Gro72] Doris Lloyd Grosh. A Bayes sampling allocation scheme for stratified finite populations with hyperbinomial prior distributions. Technometrics, 14(3):599–612, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267288.

Groeneveld:1977:BRB

[Gro77] Richard A. Groeneveld. Book review: Introduction to Statistics and Computer Programming by Carl F. Kossack; Claudia I. Henschke. Technometrics, 19(3):341–342, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267705.

Grubbs:1971:BRB

[Gru71a] F. E. Grubbs. Book review: Order Statistics by Herbert A. David. Technometrics, 13(3):706–708, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267186.

Grubbs:1971:AFBa

[Gru71b] Frank E. Grubbs. Approximate fiducial bounds for the reliability of a series system for which each component has an exponential timeto-fail distribution. Technometrics, 13(4):865–871, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266962.

Grubbs:1971:AFBb

[Gru71c] Frank E. Grubbs. Approximate fiducial bounds on reliability for the two parameter negative exponential distribution. Technometrics, 13(4):873–876, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266963.

Grubbs:1973:DNF

[Gru73a] Frank E. Grubbs. Determination of number of failures to control risks of erroneous judgments in exponential life testing. Technometrics, 15(1):191–193, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266837.

Grubbs:1973:EMP

[Gru73b] Frank E. Grubbs. Errors of measurement, precision, accuracy and the statistical comparison of measuring instruments. Technometrics, 15(1):53–66, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266824.

Guttman:1971:IRD

[GS71] Irwin Guttman and Dennis E. Smith. Investigation of rules for dealing with outliers in small samples from the normal distribution II: Estimation of the variance. Technometrics, 13(1):101–111, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267078.

Gualtierotti:1971:BRB

[Gua71] Antonio F. Gualtierotti. Book review: Gestion de la Qualité: Aspects Économiques et Techniques (Management of Quality: Economical and Theoretical Features) by L. Larrieu. Technometrics, 13(2): 452, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266811.

Guenther:1971:ASN

[Gue71a] William C. Guenther. The average sample number for truncated double sample attribute plans. Technometrics, 13(4):811–816, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266956.

Guenther:1971:DSS

[Gue71b] William C. Guenther. On the determination of single sampling attribute plans based upon a linear cost model and a prior distribution. Technometrics, 13(3):483–498, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267163. See corrigendum [Ano72g].

Guenther:1972:SAN

[Gue72] William C. Guenther. A simple approximation to the negative binomial (and regular binomial). Technometrics, 14(2):385–389, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267429.

Guenther:1974:SSF

[Gue74] William C. Guenther. Sample size formulas for some binomial type problems. Technometrics, 16(3):465–467, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267679.

Guenther:1977:EMO

[Gue77] William C. Guenther. An easy method for obtaining percentage points of order statistics. Technometrics, 19(3):319–321, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267702.

Guttman:1973:CHU

[Gut73a] Irwin Guttman. Care and handling of univariate or multivariate outliners in detecting spuriosity: A Bayesian approach. Technometrics, 15(4):723–738, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267384.

Guttman:1973:PPS

[Gut73b] Irwin Guttman. Premium and protection of several procedures for dealing with outliers when sample sizes are moderate to large. Technometrics, 15(2):385–404, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266997.

Goel:1971:DRC

[GW71] Amrit L. Goel and S. M. Wu. Determination of A. R. L. and a contour nomogram for Cusum charts to control normal mean. Technometrics, 13(2):221–230, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266785.

Gentleman:1975:DOT

[GW75] J. F. Gentleman and M. B. Wilk. Detecting outliers in a two-way table: I. Statistical behavior of residuals. Technometrics, 17(1):1–14, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267994.

Gunst:1976:CLS

[GWM76] R. F. Gunst, J. T. Webster, and R. L. Mason. A comparison of least squares and latent root regression estimators. Technometrics, 18(1):75–83, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267919.

Gabriel:1979:LRA

[GZ79] K. Ruben Gabriel and S. Zamir. Lower rank approximation of matrices by least squares with any choice of weights. Technometrics, 21(4):489–498, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268288. See corrigenda [Ano80a].

Hackney:1978:BRB

[Hac78] Olga P. Hackney. Book review: Elementary Probability and Statistical Reasoning by Howard E. Reinhardt; Don O. Loftsgaarden. Technometrics, 20(4):503, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267657.

Hackney:1979:BRB

[Hac79] Olga P. Hackney. Book review: An Introduction to Probability Theory and Mathematical Statistics by V. K. Rohatgi. Technometrics, 21(3): 391, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267772.

Hahn:1972:SPI

[Hah72] Gerald J. Hahn. Simultaneous prediction intervals for a regression model. Technometrics, 14(1):203–214, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266931.

Hahn:1975:SPL

[Hah75] Gerald J. Hahn. A simultaneous prediction limit on the means of future samples from an exponential distribution. Technometrics, 17(3):341–345, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268071.

Hahn:1977:PID

[Hah77] Gerald J. Hahn. A prediction interval on the difference between two future sample means and its application to a claim of product superiority. Technometrics, 19(2):131–134, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268622.

Hahn:1979:BRB

[Hah79] Gerald J. Hahn. Book review: Regression Analysis by Example by S. Chatterjee; B. Price. Technometrics, 21(2):271–272, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268529.

Halperin:1970:IEL

[Hal70] Max Halperin. On inverse estimation in linear regression. Technometrics, 12(4):727–736, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267319.

Hald:1977:NDA

[Hal77a] Anders Hald. A note on the determination of attribute sampling plans of given strength. Technometrics, 19(2):211–212, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268632.

Haller:1977:BRB

[Hal77b] H. Smith Haller. Book review: An Introduction to Mathematical Statistics by H. D. Brunk. Technometrics, 19(2):217–218, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268634.

Hall:1979:MVV

[Hal79] James E. Hall. Minimum variance and VOQL sampling plans. Technometrics, 21(4):555–565, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268296.

Hamdan:1970:LE

[Ham70] M. A. Hamdan. Letter to the Editor. Technometrics, 12(3):719, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267225.

Hamdan:1972:ETB

[Ham72a] M. A. Hamdan. Estimation in the truncated bivariate Poisson distribution. Technometrics, 14(1):37–45, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266916.

Hamdan:1972:CCE

[Ham72b] M. A. Hamdan. On a characterization by conditional expectations. Technometrics, 14(2):497–499, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267438.

Hamdan:1974:CC

[Ham74] M. A. Hamdan. On the continuity correction. Technometrics, 16(4):631–632, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1267619. See errata [Ano75f].

Hammood:1975:BRB

[Ham75] Abdul W. Hammood. Book review: Finite Mathematics with Business Applications by J. G. Kemeny; A. Schleifer; J. L. Snell; G. L. Thompson. Technometrics, 17(2):276–277, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268369.

Hannan:1973:BRB

[Han73] E. J. Hannan. Book review: Principles of Econometrics by Henri Theil. Technometrics, 15(1):195–196, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266838.

Hanumara:1975:EIM

[Han75] R. Choudary Hanumara. Estimating imprecisions of measuring instruments. Technometrics, 17(3):299–302, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268064.

Harris:1970:BRB

[Har70] Bernard Harris. Book review: A Graduate Course in Probability by Howard G. Tucker. Technometrics, 12(1):180, February 1970. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267373.

Harter:1971:MPP

[Har71] H. Leon Harter. More percentage points of the Pearson Type III distribution. Technometrics, 13(1):203–204, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267089.

Hartley:1972:BRB

[Har72] H. O. Hartley. Book review: Selected Tables in Mathematical Statistics, Volume I by H. L. Harter; D. B. Owen. Technometrics, 14(3): 803, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267307.

Harter:1973:LRT

[Har73a] H. Leon Harter. [Likelihood Ratio Test for Discrimination between Two Models with Unknown Location and Scale Parameters]: Discussion no. 2. Technometrics, 15(1):29–31, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266822. See [DAH73].

Harville:1973:FPL

[Har73b] David A. Harville. Fitting partially linear models by weighted least squares. Technometrics, 15(3):509–515, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266856.

Harville:1974:NOA

[Har74] David A. Harville. Nearly optimal allocation of experimental units using observed covariate values. Technometrics, 16(4):589–599, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267612.

Harris:1975:BRB

[Har75] Bernard Harris. Book review: Random Data: Analysis and Measurement Procedures by Julius S. Bendat; Allan G. Piersol. Technometrics, 17(2):271, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268362.

Hare:1977:BRB

[Har77] Lynne B. Hare. Book review: Elementary Statistical Concepts by Ronald E. Walpole. Technometrics, 19(2):218, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268636.

Hare:1979:DME

[Har79] Lynne B. Hare. Designs for mixture experiments involving process variables. Technometrics, 21(2):159–173, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268513.

Hassanein:1972:SEP

[Has72] Khatab M. Hassanein. Simultaneous estimation of the parameters of the extreme value distribution by sample quantiles. Technometrics, 14(1):63–70, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266918.

Hawkins:1972:NMS

[Haw72] Richard H. Hawkins. A note on multiple solutions to the mixed distribution problem. Technometrics, 14(4):973–976, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267147.

Hawkins:1975:RBR

[Haw75] Douglas M. Hawkins. Relations between ridge regression and eigenanalysis of the augmented correlation matrix. Technometrics, 17 (4):477–480, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268435.

Hawkins:1978:LE

[Haw78] Douglas Hawkins. Letter to the Editor. Technometrics, 20(2):218, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268730.

Hayes:1973:BRB

[Hay73] Robert M. Hayes. Book review: The Evaluation of Information Services and Products by Donald W. King; Edward C. Bryant. Technometrics, 15(4):940–941, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267408.

Hewett:1971:ISM

[HB71] John E. Hewett and William G. Bulgren. Inequalities for some multivariate f-distributions with applications. Technometrics, 13(2):397– 402, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266800.

Hemmerle:1978:ECG

[HB78] W. J. Hemmerle and T. F. Brantle. Explicit and constrained generalized ridge estimation. Technometrics, 20(2):109–120, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268701.

Hager:1971:REG

[HBA71] Harold W. Hager, Lee J. Bain, and Charles E. Antle. Reliability estimation for the generalized gamma distribution and robustness of the Weibull model. Technometrics, 13(3):547–557, August 1971.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267167.

Haynes:1970:WTL

[HD70] Tyler Haynes and Edward A. Davis. Waiting-time for a large gap in an alternating renewal process. Technometrics, 12(3):697–699, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267215.

Hearne:1978:BRB

[Hea78] F. Terry Hearne. Book review: BMDP-77 Biomedical Computer Programs, P Series by W. J. Dixon; M. B. Brown. Technometrics, 20(3): 320, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268144.

Hedayat:1971:STM

[Hed71] A. Hedayat. A set of three mutually orthogonal Latin squares of order 15. Technometrics, 13(3):696–698, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267183.

Heil:1976:BRB

[Hei76] Peter F. Heil. Book review: Survival Distributions: Reliability Applications in the Biomedical Sciences by Alan J. Gross; Virginia A. Clark. Technometrics, 18(4):501, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268667.

Heil:1978:BRB

[Hei78] Peter F. Heil. Book review: Probability and Statistical Inference by Robert V. Hogg; Elliot A. Tanis. Technometrics, 20(4):503, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267656.

Heil:1979:BRB

[Hei79] Peter F. Heil. Book review: Introduction to Probability and Statistics by Narayan C. Giri. Technometrics, 21(1):135, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268593.

Helms:1971:BRBa

[Hel71a] Ronald W. Helms. Book review: Approximations to the Poisson, Binomial, and Hypergeometric Distribution Functions by W. Molenaar.

Technometrics, 13(1):214, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267096.

Helms:1971:BRBb

[Hel71b] Ronald W. Helms. Book review: Artificial Intelligence by Bernard Meltzer. Technometrics, 13(2):451–452, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266810.

Helms:1974:AEV

[Hel74] Ronald W. Helms. The average estimated variance criterion for the selection-of-variables problem in general linear models. Technometrics, 16(2):261–273, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267949.

Hemmerle:1975:ESG

[Hem75] William J. Hemmerle. An explicit solution for generalized ridge regression. Technometrics, 17(3):309–314, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268066.

Herbach:1970:EAE

[Her70] L. H. Herbach. An exact asymptotically efficient confidence bound for reliability in the case of the Gumbel distribution. Technometrics, 12(3):700–701, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267216.

Haseman:1973:ANI

[HG73] J. K. Haseman and D. W. Gaylor. An algorithm for non-iterative estimation of multiple missing values for crossed classifications. Technometrics, 15(3):631–636, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266869.

Hunter:1978:BAE

[HG78] A. J. Hunter and H. J. Griffiths. Bayesian approach to estimation of insect population size. Technometrics, 20(3):231–234, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268130.

Hemmerle:1973:CML

[HH73] W. J. Hemmerle and H. O. Hartley. Computing maximum likelihood estimates for the mixed A. O. V. model using the W transformation. Technometrics, 15(4):819–831, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267392.

Hill:1974:DES

[HH74] William J. Hill and William G. Hunter. Design of experiments for subsets of parameters. Technometrics, 16(3):425–434, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267673.

Habbema:1977:SVD

[HH77] J. D. F. Habbema and J. Hermans. Selection of variables in discriminant analysis by F -statistic and error rate. Technometrics, 19(4):487–493, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267890.

Hill:1974:BRB

[Hil74] Bruce M. Hill. Book review: Bayesian Inference in Statistical Analysis by George E. P. Box; George C. Tiao. Technometrics, 16(3):478– 479, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267683.

Hill:1978:RED

[Hil78] Peter D. H. Hill. A review of experimental design procedures for regression model discrimination. Technometrics, 20(1):15–21, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268155.

Hines:1977:CSW

[Hin77a] W. G. S. Hines. A comparison of some weighted averages. Technometrics, 19(3):313–318, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267701.

Hinkley:1977:JUS

[Hin77b] David V. Hinkley. Jackknifing in unbalanced situations. Technometrics, 19(3):285–292, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267698.

Holms:1973:LE

[HJ73] Arthur G. Holms and Eric E. Johnson. Letters to the Editor. Technometrics, 15(4):947, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267413.

Hegemann:1976:ATW

[HJ76] Victor Hegemann and Dallas E. Johnson. On analyzing two-way AoV data with interaction. Technometrics, 18(3):273–281, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268736.

Hill:1978:TAR

[HJF78] R. Carter Hill, George G. Judge, and Thomas B. Fomby. On testing the adequacy of a regression model. Technometrics, 20(4):491–494, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267650.

Hoerl:1970:RRA

[HK70a] Arthur E. Hoerl and Robert W. Kennard. Ridge regression: Applications to nonorthogonal problems. Technometrics, 12(1):69–82, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267352.

See [Ano70f].

Hoerl:1970:RRB

[HK70b] Arthur E. Hoerl and Robert W. Kennard. Ridge regression: Biased estimation for nonorthogonal problems. Technometrics, 12(1):55–67, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267351.

Hoerl:1975:NPG

[HK75] Arthur E. Hoerl and Robert W. Kennard. A note on a power generalization of ridge regression. Technometrics, 17(2):269, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268361.

Hoerl:1977:LEC

[HK77] Arthur E. Hoerl and Robert W. Kennard. Letter to the editor: [Comments on: “Simulation and extension of a minimum mean squared error estimator in comparison with Stein’s” (Technometrics 18 (1976), no. 4, 491–496) by H. D. Vinod]. Technometrics, 19(1):105, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268272.

Hemmerle:1976:IAT

[HL76] W. J. Hemmerle and J. A. Lorens. Improved algorithm for the W -transform in variance component estimation. Technometrics, 18(2):207–211, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267524.

Helms:1971:DIP

[HM71] Ronald W. Helms and Scott E. Michaels. Discussion of invited papers. Technometrics, 13(3):477–481, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267162.

Hebble:1972:RRS

[HM72] T. L. Hebble and T. J. Mitchell. “repairing” response surface designs. Technometrics, 14(3):767–779, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267304.

Hinich:1973:FNM

[HM73] M. J. Hinich and H. Montes. Fitting a nonstationary model of ionospheric motions. Technometrics, 15(4):759–767, November 1973. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267386.

Hald:1976:AFC

[HM76a] Anders Hald and Uffe Møller. The ASN function for curtailed single sampling by attributes. Technometrics, 18(3):307–312, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268740.

Hewett:1976:SAS

[HM76b] John E. Hewett and M. L. Moeschberger. Some approximate simultaneous prediction intervals for reliability analysis. Technometrics, 18(2):227–229, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267527.

Hettmansperger:1977:RAB

[HM77] Thomas P. Hettmansperger and Joseph W. McKean. A robust alternative based on ranks to least squares in analyzing linear models. Technometrics, 19(3):275–284, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267697.

Hahn:1970:PSC

[HN70] Gerald J. Hahn and Wayne Nelson. A problem in the statistical comparison of measuring devices. Technometrics, 12(1):95–102, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267354.

Hocking:1972:CSS

[Hoc72] R. R. Hocking. Criteria for selection of a subset regression: Which one should be used? Technometrics, 14(4):967–970, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267145.

Halfon:1977:LE

[HOC^{+}77] Efraim Halfon, R. L. Obenchain, R. Dennis Cook, D. G. Kabe, Richard L. Scheaffer, G. K. Shukla, and William C. Guenther. Letters to the Editor. Technometrics, 19(3):348–351, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267717.

Hogg:1974:BRB

[Hog74] Robert V. Hogg. Book review: Probability and Statistics by Julius R. Blum; Judah I. Rosenblatt. Technometrics, 16(2):326, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267958.

Hoke:1974:ESO

[Hok74] Albert T. Hoke. Economical second-order designs based on irregular fractions of the 3^{n} factorial. Technometrics, 16(3):375–384, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267667.

Holland:1973:BRB

[Hol73] Paul W. Holland. Book review: Introduction to Bayesian Inference and Decision by Robert L. Winkler. Technometrics, 15(4):938–939, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267406.

Holland:1975:SRD

[Hol75] Burt S. Holland. Some results on the discretization of continuous probability distributions. Technometrics, 17(3):333–339, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268070.

Hooke:1979:DSM

[Hoo79] Robert Hooke. Discussion of: “A Simple Method for Regression Analysis with Censored Data”. Technometrics, 21(4):433–434, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268281. See [SH79].

Howe:1977:BRB

[How77] William G. Howe. Book review: Probability and Statistics by M. H. DeGroot. Technometrics, 19(2):218, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268635.

Herman:1971:MLE

[HP71] R. J. Herman and Rusi K. N. Patell. Maximum likelihood estimation for multi-risk model. Technometrics, 13(2):385–396, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266799.

Hall:1973:OSP

[HP73] I. J. Hall and R. R. Prairie. One-sided prediction intervals to contain at least m out of k future observations. Technometrics, 15(4):897–914, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267399.

Hader:1978:SRC

[HP78] R. J. Hader and Sung H. Park. Slope-rotatable central composite designs. Technometrics, 20(4):413–417, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267641.

Hogg:1975:ADF

[HR75] Robert V. Hogg and Ronald H. Randles. Adaptive distributionfree regression methods and their applications. Technometrics, 17 (4):399–407, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268426.

Holms:1971:DED

[HS71] Arthur G. Holms and Steven M. Sidik. Design of experiments as “doubly telescoping” sequences of blocks with application to a nuclear reactor experiment. Technometrics, 13(3):559–574, August 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267168.

Hocking:1972:OIM

[HS72] R. R. Hocking and W. B. Smith. Optimum incomplete multinormal samples. Technometrics, 14(2):299–307, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267422.

Heller:1973:RSO

[HS73] Nelson B. Heller and Glenn E. Staats. Response surface optimization when experimental factors are subject to costs and constraints. Technometrics, 15(1):113–123, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266829.

Hocking:1976:CBE

[HSL76] R. R. Hocking, F. M. Speed, and M. J. Lynn. A class of biased estimators in linear regression. Technometrics, 18(4):425–437, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268658.

Hannan:1974:EET

[HT74] E. J. Hannan and P. J. Thomson. Estimating echo times. Technometrics, 16(1):77–84, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267495.

Huber:1978:BRB

[Hub78] John Huber. Book review: Introduction to the Use of Computer Packages for Statistical Analyses by Richard W. Moore. Technometrics, 20(4):504, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267659.

Hultquist:1973:BRB

[Hul73] Robert Hultquist. Book review: Linear Models by S. R. Searle. Technometrics, 15(2):419, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267001.

Hunt:1970:BRB

[Hun70] Jerry G. Hunt. Book review: Stochastic Theory of a Risk Business by Hilary L. Seal. Technometrics, 12(1):176–178, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267371.

Hunter:1974:BRB

[Hun74] Jeffrey J. Hunter. Book review: Problems in Complex Variable Theory by Jan G. Krzyz. Technometrics, 16(3):481–482, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267687.

Haile:1975:CFD

[HW75] Alemayehu Haile and J. T. Webster. A comparison of four designs for estimating the variances of random effects in a two-way classification. Technometrics, 17(4):495–500, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268438.

Iman:1979:URT

[IC79] Ronald L. Iman and W. J. Conover. The use of the rank transform in regression. Technometrics, 21(4):499–509, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268289. See corrigendum

[Ano84].

Iglewicz:1973:AME

[Igl73] Boris Iglewicz. An alternative measure of the efficiency of sequential testing procedures. Technometrics, 15(3):571–574, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266863.

Iglewicz:1970:CAP

[IM70] Boris Iglewicz and Raymond H. Myers. Comparisons of approximations to the percentage points of the sample coefficient of variation. Technometrics, 12(1):166–169, February 1970. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267363.

Inglis:1975:BRB

[Ing75] James Inglis. Book review: Theoretical Statistics by D. R. Cox; D. V. Hinkley. Technometrics, 17(4):513, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268442.

Inglis:1978:BRBa

[Ing78a] James Inglis. Book review: A Mathematical Theory of Evidence by Glenn Shafer. Technometrics, 20(1):106, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268172.

Inglis:1978:BRBb

[Ing78b] James Inglis. Book review: Sufficient Statistics: Selected Contributions by Vasant S. Hazurbazar; Anant M. Kshirsagar. Technometrics, 20(1):106, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268173.

Izenman:1977:SDR

[IR77] Alan J. Izenman and Yosef Rinott. Sequential designs for replacements in exponential life tests. Technometrics, 19(4):455–459, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267885.

Ingram:1976:CEA

[IS76] C. R. Ingram and R. L. Scheaffer. On consistent estimation of age replacement intervals. Technometrics, 18(2):213–219, May 1976. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267525.

J:1975:BRBg

[J.75a] J. E. J. Book review: A Survey of Statistical Design and Linear Models by J. N. Srivastava. Technometrics, 17(4):514–515, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268446.

J:1975:BRBd

[J.75b] J. E. J. Book review: Applied Statistics by R. P. Gupta. Technometrics, 17(3):389–390, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268090.

J:1975:BRBb

[J.75c] J. E. J. Book review: Multivariate Analysis, a Selected and Abstracted Bibliography 1957–1972 by Kocherlakota Subrahmaniam; Kathleen Subrahmaniam. Technometrics, 17(3):386, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268084.

J:1975:BRBa

[J.75d] J. E. J. Book review: Point Estimation in the Linear Regression Model by Mario Faliva. Technometrics, 17(3):385–386, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268083.

J:1975:BRBf

[J75e] J. E. J. Book review: Progress in Statistics by J. Gani; K. Sakradi; I. Vincze. Technometrics, 17(3):391–392, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268092. See erratum [Ano76f].

J:1975:BRBh

[J.75f] J. E. J. Book review: Reliability and Biometry, Statistical Analysis of Lifelength by F. Proschan; R. J. Serfling. Technometrics, 17(4):515–516, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268447.

J:1975:BRBc

[J.75g] J. E. J. Book review: Selected Tables in Mathematical Statistics, Vol. II by Institute of Mathematical Statistics; H. L. Harter; D. B. Owen. Technometrics, 17(3):386, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268085.

J:1975:BRBe

[J.75h] J. E. J. Book review: Studies in Bayesian Econometrics and Statistics (In Honor of Leonard J. Savage) by S. E. Fienberg; A. Zellner. Technometrics, 17(3):390–391, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268091.

J:1976:BRBg

[J.76a] J. E. J. Book review: Contributions to Applied Statistics by W. J. Ziegler. Technometrics, 18(3):363, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268750.

J:1976:BRBb

[J.76b] J. E. J. Book review: Developments in Field Experiment Design and Analysis by V. J. Bofinger; J. L. Wheeler. Technometrics, 18(1):122, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267929.

J:1976:BRBa

[J.76c] J. E. J. Book review: Elementary Statistics by G. A. F. Seber. Technometrics, 18(1):120–121, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267926.

J:1976:BRBd

[J.76d] J. E. J. Book review: Multivariate Statistical Methods: Among-Groups Covariation by W. R. Atchley; E. H. Bryant. Technometrics, 18(2):241, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267535.

J:1976:BRBe

[J.76e] J. E. J. Book review: Multivariate Statistical Methods: Within-Groups Covariation by E. H. Bryant; W. R. Atchley. Technometrics, 18(2):241–242, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267536.

J:1976:BRBf

[J.76f] J. E. J. Book review: Probability, Statistics and Time: A Collection of Essays by M. S. Bartlett. Technometrics, 18(3):362, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268749.

J:1976:BRBc

[J.76g] J. E. J. Book review: Sample Size Choice; Charts for Experiments with Linear Models by R. E. Odeh; M. Fox. Technometrics, 18(2):240, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267531.

J:1976:BRBi

[J.76h] J. E. J. Book review: Selected Tables in Mathematical Statistics, Vol. 3 by The Institute of Mathematical Statistics; H. L. Harter; D. B. Owen. Technometrics, 18(4):502–503, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268671.

J:1976:BRBh

[J.76i] J. E. J. Book review: The Delphi Method: Techniques and Applications by H. A. Linstone; M. Turoff. Technometrics, 18(3):363–364, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268751.

J:1977:BRBn

[J.77a] J. E. J. Book review: Advances in Operations Research by Marc Roubens. Technometrics, 19(4):519–520, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267904.

J:1977:BRBa

[J.77b] J. E. J. Book review: An Annotated Bibliography on Methods for Non-Linear Least Squares Computations Including Test Problems by John C. Nash. Technometrics, 19(1):102–103, February 1977. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268265.

J:1977:BRBf

[J.77c] J. E. J. Book review: Econometrics by Peter Schmidt. Technometrics, 19(2):219, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268638.

J:1977:BRBb

[J.77d] J. E. J. Book review: Index to Statistics and Probability: Citation Index by John W. Tukey. Technometrics, 19(1):103, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268267.

J:1977:BRBd

[J.77e] J. E. J. Book review: Index to Statistics and Probability: Locations and Authors by I. C. Ross; J. W. Tukey. Technometrics, 19(1):104, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268269.

J:1977:BRBc

[J.77f] J. E. J. Book review: Index to Statistics and Probability: Permuted Titles by I. C. Ross; J. W. Tukey. Technometrics, 19(1):104, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268268.

J:1977:BRBi

[J.77g] J. E. J. Book review: Multiple Hypergeometric Functions and Applications by Harold Exton. Technometrics, 19(3):347, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267715.

J:1977:BRBm

[J.77h] J. E. J. Book review: Multivariate Analysis IV by P. R. Krishnaiah. Technometrics, 19(4):518–519, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267903.

J:1977:BRBk

[J.77i] J. E. J. Book review: Optimal Strategies in Sports by Shaul P. Ladany; Robert E. Machol. Technometrics, 19(4):517–518, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267901.

J:1977:BRBh

[J.77j] J. E. J. Book review: Reliability and Fault Tree Analysis by R. E. Barlow; J. B. Fussell; N. D. Singpurwalla. Technometrics, 19(3):346– 347, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267714.

J:1977:BRBl

[J.77k] J. E. J. Book review: Selected Tables in Mathematical Statistics, Volume 4 by Institute of Mathematical Statistics; D. B. Owen; R. E. Odeh. Technometrics, 19(4):518, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267902.

J:1977:BRBe

[J.77l] J. E. J. Book review: Studies in Probability and Statistics: Papers in Honour of Edwin J. G. Pitman by E. J. Williams. Technometrics, 19(1):104, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268270.

J:1977:BRBj

[J.77m] J. E. J. Book review: The Analysis of Survey Data by C. A. O’Muircheartaigh; C. Payne. Technometrics, 19(4):517, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267900.

J:1977:BRBg

[J.77n] J. E. J. Book review: Theory and Applications of Reliability by C. P. Tsokos; I. N. Shimi. Technometrics, 19(3):345–346, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267713.

J:1977:RBSa

[J.77o] J. E. J. Review: Index to Statistics and Probability: Citation Index, by John W. Tukey. Technometrics, 19(1):103, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268267.

J:1977:RBSc

[J.77p] J. E. J. Review: Index to Statistics and Probability: Location and Authors, by I. C. Ross and J. W. Tukey. Technometrics, 19(1):104, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268269.

J:1977:RBSb

[J.77q] J. E. J. Review: Index to Statistics and Probability: Permuted Titles, by I. C. Ross and J. W. Tukey. Technometrics, 19(1):104, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268268.

J:1978:BRBp

[J.78a] J. E. J. Book review: Applications of Statistics by P. R. Krishnaiah. Technometrics, 20(3):322–323, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268148.

J:1978:BRBo

[J.78b] J. E. J. Book review: Conflicting Objectives in Decisions by D. E. Bell; R. L. Keeney; H. Raiffa. Technometrics, 20(3):322, August 1978.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268147.

J:1978:BRBn

[J.78c] J. E. J. Book review: Contributions to Survey Sampling and Applied Statistics by H. A. David. Technometrics, 20(3):321–322, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268146.

J:1978:BRBh

[J.78d] J. E. J. Book review: Interactive Data Analysis by D. R. McNeil. Technometrics, 20(2):212, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268720.

J:1978:BRBe

[J.78e] J. E. J. Book review: Management Science in Sports by R. E. Machol; S. P. Ladany; D. G. Morrison. Technometrics, 20(1):106, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268175.

J:1978:BRBf

[J.78f] J. E. J. Book review: Methods for Statistical Data Analysis of Multivariate Observations by R. Gnanadesikan. Technometrics, 20(2):210– 211, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268717.

J:1978:BRBg

[J.78g] J. E. J. Book review: MINITAB: Student Handbook by T. A. Ryan; B. L. Joiner; B. F. Ryan. Technometrics, 20(2):211–212, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268719.

J:1978:BRBb

[J.78h] J. E. J. Book review: Multiple Criteria Decision Making by M. K. Starr; M. Zeleny. Technometrics, 20(1):105, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268169.

J:1978:BRBk

[J.78i] J. E. J. Book review: New Developments in the Applications of Bayesian Methods by A. Aykaç; C. Brumat. Technometrics, 20(2):214–215, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268726.

J:1978:BRBm

[J.78j] J. E. J. Book review: Nuclear Systems Reliability Engineering and Risk Assessment by J. B. Fussell; G. R. Burdick. Technometrics, 20(3):321, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268145.

J:1978:BRBd

[J.78k] J. E. J. Book review: On the History of Statistics and Probability by D. B. Owen. Technometrics, 20(1):106, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268174.

J:1978:BRBc

[J.78l] J. E. J. Book review: Pocket Book of Statistical Tables by R. E. Odeh; D. B. Owen; Z. W. Birnbaum; L. Fisher. Technometrics, 20(1):105, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268170.

J:1978:BRBl

[J.78m] J. E. J. Book review: Recent Developments in Statistics by J. R. Barra; F. Brodeau; G. Romier; B. van Cutsem. Technometrics, 20(2):215–217, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268728.

J:1978:BRBa

[J.78n] J. E. J. Book review: Sampling Techniques by W. G. Cochran. Technometrics, 20(1):104, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268167.

J:1978:BRBq

[J.78o] J. E. J. Book review: Selected Tables in Mathematical Statistics, Vol. 5 by Institute of Mathematical Statistics; D. B. Owen; R. E. Odeh. Technometrics, 20(3):323, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268149.

J:1978:BRBi

[J.78p] J. E. J. Book review: Statistical Tables for Science, Engineering, Management and Business Studies by J. Murdoch; S. A. Baines. Technometrics, 20(2):214, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268724.

J:1978:BRBj

[J.78q] J. E. J. Book review: Tables of Random Times by I. E. Hill; P. A. Hill. Technometrics, 20(2):214, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268725.

J:1979:BRBq

[J.79a] J. E. J. Book review: A Sampler on Sampling by Bill Williams. Technometrics, 21(4):583–584, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268302.

J:1979:BRBm

[J.79b] J. E. J. Book review: Algorithmic Methods in Probability by Marcel F. Neuts. Technometrics, 21(3):392, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267775.

J:1979:BRBg

[J.79c] J. E. J. Book review: Analyzing Qualitative/Categorical Data by Leo A. Goodman. Technometrics, 21(2):272, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268530.

J:1979:BRBo

[J.79d] J. E. J. Book review: Applied Optimal Control by A. Bensoussan; P. R. Kleindorfer; C. S. Tapiero. Technometrics, 21(3):393, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267777.

J:1979:BRBe

[J.79e] J. E. J. Book review: Applied Time Series Analysis by D. F. Findley. Technometrics, 21(1):136, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268596.

J:1979:BRBt

[J.79f] J. E. J. Book review: Developments in Statistics, Volume 2 by P. R. Krishnaiah. Technometrics, 21(4):585, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268306.

J:1979:BRBd

[J.79g] J. E. J. Book review: Developments in Statistics, Volume I by P. R. Krishnaiah. Technometrics, 21(1):136, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268595.

J:1979:BRBu

[J.79h] J. E. J. Book review: Energy and Health by N. E. Breslow; A. S. Whittemore. Technometrics, 21(4):585, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268307.

J:1979:BRBb

[J.79i] J. E. J. Book review: Graphical Techniques for Multivariate Data by Brian Everitt. Technometrics, 21(1):134, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268591.

J:1979:BRBf

[J.79j] J. E. J. Book review: Handbook of Hypergeometric Integrals by Harold Exton. Technometrics, 21(1):136, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268597.

J:1979:BRBj

[J.79k] J. E. J. Book review: International Encyclopedia of Statistics by William H. Kruskal; Judith M. Tanur. Technometrics, 21(3):388–389, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267768.

J:1979:BRBh

[J.79l] J. E. J. Book review: Maps and Statistics by Peter Lewis. Technometrics, 21(2):273, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268532.

J:1979:BRBv

[J.79m] J. E. J. Book review: Modelling and Simulation Practice by M. Cross; R. D. Gibson; M. J. O’Carroll; T. S. Wilkinson. Technometrics, 21(4):586, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268308.

J:1979:BRBp

[J.79n] J. E. J. Book review: Operational Research ’78 by K. B. Haley. Technometrics, 21(3):393–394, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267778.

J:1979:BRBk

[J.79o] J. E. J. Book review: Problems and Solutions in Theoretical Statistics by D. R. Cox; D. V. Hinkley. Technometrics, 21(3):390, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267770.

J:1979:BRBc

[J.79p] J. E. J. Book review: R. A. Fisher, the Life of a Scientist by Joan Fisher Box. Technometrics, 21(1):135, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268594.

J:1979:BRBa

[J.79q] J. E. J. Book review: Statistical Methods for Digital Computers, Volume III by K. Enslein; A. Ralston; H. S. Wilf. Technometrics, 21(1):133–134, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268590.

J:1979:BRBi

[J.79r] J. E. J. Book review: Statistics by R. F. Hirsch. Technometrics, 21 (2):274, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268534.

J:1979:BRBr

[J.79s] J. E. J. Book review: Statistics for Technology by Christopher Chatfield. Technometrics, 21(4):584–585, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268304.

J:1979:BRBl

[J.79t] J. E. J. Book review: Survey Sampling and Measurement by N. K. Namboodiri. Technometrics, 21(3):391–392, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267774.

J:1979:BRBs

[J.79u] J. E. J. Book review: Tables of the F - and Related Distributions with Algorithms by K. V. Mardia; P. J. Zemroch. Technometrics, 21(4):585, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268305.

J:1979:BRBn

[J.79v] J. E. J. Book review: The H-Function with Applications in Statistics and Other Disciplines by A. M. Mathai; R. K. Saxena. Technometrics, 21(3):392–393, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267776.

Jackson:1980:BRBc

[Jac80] J. Edward Jackson. Book review: Three Aspects of Policy and Policymaking: Knowledge, Data and Institutions by Karl Brunner; Allan H. Meltzer. Technometrics, 22(1):134, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268404. See corrigenda [Ano80a].

Jaech:1970:CTM

[Jae70] John L. Jaech. Comparing two methods of obtaining a confidence interval for the ratio of Poisson parameters. Technometrics, 12(2):383– 387, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267242.

Jaech:1974:DSP

[Jae74] John L. Jaech. Determination of system performance by the generation of conditional system moments. Technometrics, 16(4):627–629, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267618.

Jaech:1976:LST

[Jae76] John L. Jaech. Large sample tests for Grubbs’ estimators of instrument precision with more than two instruments. Technometrics, 18(2):127–133, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267516.

Johnson:1974:ESC

[JB74] Richard A. Johnson and Michael Bagshaw. The effect of serial correlation on the performance of CUSUM tests. Technometrics, 16(1):103–112, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267498.

Jennrich:1976:FSL

[JB76] Robert I. Jennrich and Peter B. Bright. Fitting systems of linear differential equations using computer generated exact derivatives. Technometrics, 18(4):385–392, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1268653. See discussion [Wig76].

Joiner:1976:DEW

[JC76] Brian L. Joiner and Cathy Campbell. Designing experiments when run order is important. Technometrics, 18(3):249–259, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268734.

John:1978:TTO

[JD78] J. A. John and N. R. Draper. On testing for two outliers or one outlier in two-way tables. Technometrics, 20(1):69–78, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268162.

Johnson:1974:LE

[JEBE74] Arthur F. Johnson, A. W. F. Edwards, Lee J. Bain, and Max Engelhardt. Letters to the Editor. Technometrics, 16(4):641–642, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267629.

Jerome:1970:BRB

[Jer70] Joseph Jerome. Book review: An Introduction to the Approximation of Functions by Theodore J. Rivlin. Technometrics, 12(2):425–426, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267251.

Jackson:1973:RAC

[JH73a] J. Edward Jackson and F. Terry Hearne. Relationships among coefficients of vectors used in principal components. Technometrics, 15(3):601–610, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266866.

Johnson:1973:BRBa

[JH73b] William D. Johnson and Wolfgang Habermann. Book review: Statistische Methoden Ein Soforthelfer by Luther Sachs. Technometrics, 15(1):202, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266845.

Jackson:1979:HPC

[JH79] J. Edward Jackson and F. Terry Hearne. Hotelling’s T_{M}^{2} for principal components: What about absolute values? Technometrics, 21(2):253–255, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268524.

Johnson:1971:NWL

[JK71] William D. Johnson and Gary G. Koch. A note on the weighted least squares analysis of the ries-Smith contingency table data. Technometrics, 13(2):438–447, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266806.

Johnson:1974:APM

[JK74] Charles A. Johnson and Jerome H. Klotz. The atom probe and Markov chain statistics of clustering. Technometrics, 16(4):483–493, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267599.

Johnson:1979:BSS

[JL79] Mark E. Johnson and Victor W. Lowe, Jr. Bounds on the sample skewness and kurtosis. Technometrics, 21(3):377–378, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267763.

Jackson:1979:CPR

[JM79] J. Edward Jackson and Govind S. Mudholkar. Control procedures for residuals associated with principal component analysis. Technometrics, 21(3):341–349, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267757.

Johnson:1978:EPO

[JMM78] Dallas E. Johnson, Stephen A. McGuire, and George A. Milliken. Estimating σ^{2} in the presence of outliers. Technometrics, 20(4):441–455, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267645.

Johansen:1970:APR

[Joh70a] S. Johansen. Asymptotic properties of the restricted Bayesian double sampling plan. Technometrics, 12(3):647–667, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267210.

John:1970:IPOb

[Joh70b] S. John. On identifying the population of origin of each observation in a mixture of observations from two gamma populations. Technometrics, 12(3):565–568, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267203.

John:1970:IPOa

[Joh70c] S. John. On identifying the population of origin of each observation in a mixture of observations from two normal populations. Technometrics, 12(3):553–563, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267202.

Johnson:1971:LCD

[Joh71a] Arthur F. Johnson. Linear combinations in designing experiments. Technometrics, 13(3):575–587, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267169.

Johnson:1971:BRB

[Joh71b] William D. Johnson. Book review: Introductory Probability and Statistical Applications by Paul L. Meyer. Technometrics, 13(1):212–213, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267093.

John:1973:BDV

[Joh73a] Peter W. M. John. A balanced design for 18 varieties. Technometrics, 15(3):641–642, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266871.

Johnson:1973:EEA

[Joh73b] Eric E. Johnson. Empirical equations for approximating tabular F values. Technometrics, 15(2):379–384, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266996.

Johnson:1973:BRBb

[Joh73c] William D. Johnson. Book review: Nonparametric Statistical Inference by Jean D. Gibbons. Technometrics, 15(2):421, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267003.

Johansson:1975:BRB

[Joh75] J. K. Johansson. Book review: Basic Statistics with Business Applications by Richard C. Clelland; John S. deCani; Francis E. Brown. Technometrics, 17(3):384–385, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268081.

John:1978:NBD

[Joh78] Peter W. M. John. A new balanced design for eighteen varieties. Technometrics, 20(2):155–158, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268707.

John:1979:MPF

[Joh79] Peter W. M. John. Missing points in 2^{n} and 2^{n−k} factorial designs. Technometrics, 21(2):225–228, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268519.

Joiner:1977:ECF

[Joi77] Brian L. Joiner. Evaluation of cryogenic flow meters: An example in non-standard experimental design and analysis. Technometrics, 19(4):353–379, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267876.

Jones:1971:CIL

[Jon71] C. F. Jones. A confidence interval for the lognormal hazard. Technometrics, 13(4):885–888, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266965.

Jones:1973:BRB

[Jon73] Robert C. Jones. Book review: Basic Statistics for Business and Economics by Paul G. Hoel; Raymond J. Jessen. Technometrics, 15(1):199–200, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266842.

Jones:1976:ADO

[Jon76a] Byron Jones. An algorithm for deriving optimal block designs. Technometrics, 18(4):451–458, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268661.

Jones:1976:BRB

[Jon76b] Thomas A. Jones. Book review: Geomathematics: Mathematical Background and Geo-Science Applications by F. P. Agterberg. Technometrics, 18(2):242–243, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267537.

Joshi:1972:EEM

[Jos72] Prakash C. Joshi. Efficient estimation of the mean of an exponential distribution when an outlier is present. Technometrics, 14(1):137–143, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266925.

Jayne:1972:PEF

[JR72] Lance W. Jayne and Edward W. Ross, Jr. Probability of equipment failure when tolerances are incorrect. Technometrics, 14(4):899–909, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267138.

Jennrich:1976:NRR

[JS76] R. I. Jennrich and P. F. Sampson. Newton–Raphson and related algorithms for maximum likelihood variance component estimation. Technometrics, 18(1):11–17, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267911.

Johnson:1975:LEN

[JTS75] Norman L. Johnson, H. Tong, and W. Spendley. Letter to the editor on: “A note on the estimation of Pr{Y < X} in the exponential case” (Technometrics 16 (1974), 625) by H. Tong. Technometrics, 17(3):393–394, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268093. With a reply by Tong.

K:1971:BRBa

[K.71a] G. G. K. Book review: An Author and Permuted Title Index to Selected Statistical Journals by Brian L. Joiner; N. F. Laubscher; Eleanor S. Brown; Bert Levy. Technometrics, 13(2):452–453, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266812.

K:1971:BRBc

[K.71b] G. G. K. Book review: Engineering Mathematics Handbook by Jan J. Tuma. Technometrics, 13(3):709, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267189.

K:1971:BRBd

[K.71c] G. G. K. Book review: Nonparametric Techniques in Statistical Inference by Madan Lai Puri. Technometrics, 13(3):709–710, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267190.

K:1971:BRBb

[K.71d] G. G. K. Book review: Random Counts in Models and Structures (Volume 1) by G. P. Patil. Technometrics, 13(2):453–454, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266813.

K:1971:BRBe

[K.71e] G. G. K. Book review: Statistical Computation by Roy C. Milton; John A. Nelder. Technometrics, 13(3):710–711, August 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267191.

K:1973:BRBe

[K.73a] G. G. K. Book review: Analysis and Simulation of Biochemical Systems, Volume 25 by H. C. Hemker; B. Hess. Technometrics, 15(4):944–945, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267412.

K:1973:BRBb

[K.73b] G. G. K. Book review: Collected Papers of R. A. Fisher, Volume I by J. H. Bennett; R. A. Fisher. Technometrics, 15(2):426–427, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267011.

K:1973:BRBc

[K.73c] G. G. K. Book review: Selected Readings in Economic Theory from Econometrica by Kenneth J. Arrow. Technometrics, 15(3):654–655, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266880.

K:1973:BRBa

[K.73d] G. G. K. Book review: Statistical Papers in Honor of George W. Snedecor by T. A. Bancroft. Technometrics, 15(2):424–425, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267009.

K:1973:BRBd

[K.73e] G. G. K. Book review: Statistics and Probability in Civil Engineering by Peter Lumb. Technometrics, 15(4):942–943, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267410.

K:1974:BRBc

[K.74a] G. G. K. Book review: Inventory Control and Water Storage by A. Prekopa. Technometrics, 16(3):482, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267688.

K:1974:BRBd

[K.74b] G. G. K. Book review: Mathematics and Statistics: Essays in Honour of Harald Bergstrom by Peter Jagers; Lennart Rade. Technometrics, 16(3):482, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267689.

K:1974:BRBa

[K.74c] G. G. K. Book review: Multivariate Statistical Inference by D. G. Kabe; R. P. Gupta. Technometrics, 16(2):328, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267963.

K:1974:BRBe

[K.74d] G. G. K. Book review: Symmetric Functions in Statistics by Derrick S. Tracy. Technometrics, 16(4):637, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267628.

K:1974:BRBb

[K.74e] G. G. K. Book review: Third Conference Advisory Group of Forest Statisticians. Technometrics, 16(2):328–329, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267964.

Kendell:1971:EPL

[KA71] P. J. Kendell and R. L. Anderson. An estimation problem in life-testing. Technometrics, 13(2):289–301, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266791.

Kabe:1974:BRB

[Kab74] D. G. Kabe. Book review: Quality Control and Reliability by Norbert L. Enrick. Technometrics, 16(1):139–140, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267508.

Kalotay:1971:SSL

[Kal71] A. J. Kalotay. Structural solution to the linear calibration problem. Technometrics, 13(4):761–769, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266953.

Kalbfleisch:1975:BRB

[Kal75] J. D. Kalbfleisch. Book review: Statistical Inference under Order Restrictions by R. E. Barlow; D. J. Bartholomew; J. M. Bremner; H. D. Brunk. Technometrics, 17(1):139–140, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268017.

Kaminsky:1972:CIE

[Kam72] Kenneth S. Kaminsky. Confidence intervals for the exponential scale parameter using optimally selected order statistics. Technometrics, 14(2):371–383, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267428.

Kaminsky:1973:CAC

[Kam73] Kenneth S. Kaminsky. Comparison of approximate confidence intervals for the exponential scale parameter from sample quantiles. Technometrics, 15(3):483–487, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266853.

Kaminsky:1974:CIT

[Kam74] Kenneth S. Kaminsky. Confidence intervals and tests for two exponential scale parameters based on order statistics in compressed samples. Technometrics, 16(2):251–254, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267947.

Kaminsky:1977:CPI

[Kam77] Kenneth S. Kaminsky. Comparison of prediction intervals for failure times when life is exponential. Technometrics, 19(1):83–86, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268259.

Kao:1972:ESC

[Kao72] Edward P. C. Kao. Economic screening of a continuously manufactured product. Technometrics, 14(3):653–661, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267293.

Kappenman:1972:NSG

[Kap72] Russell F. Kappenman. A note on selection of the greatest exceedance probability. Technometrics, 14(1):219–222, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266933.

Kappenman:1975:CCI

[Kap75] Russell F. Kappenman. Conditional confidence intervals for double exponential distribution parameters. Technometrics, 17(2):233–235, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268356.

Klimko:1975:UBP

[KARR75] Lawrence A. Klimko, Charles E. Antle, Alfred W. Rademaker, and Howard E. Rockette. Upper bounds for the power of invariant tests for the exponential distribution with Weibull alternatives. Technometrics, 17(3):357–360, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268074.

Koziol:1975:PPA

[KB75] James A. Koziol and David P. Byar. Percentage points of the asymptotic distributions of one and two sample K–S statistics for truncated or censored data. Technometrics, 17(4):507–510, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268440.

Krutchkoff:1972:LE

[KBW72] Richard G. Krutchkoff, A. R. Burgess, and E. B. Wilson. Letters to the Editor. Technometrics, 14(1):241–242, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266942.

Keesing:1975:BRB

[Kee75] Donald B. Keesing. Book review: Economic Growth, Development, and Foreign Trade: A Study in Pure Theory by Pranab K. Bardhan. Technometrics, 17(2):276, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268368.

Kennard:1971:NS

[Ken71] Robert W. Kennard. A note on the C_{p} statistic. Technometrics, 13(4):899–900, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1266968. See erratum [Ano73i].

Kettenring:1975:BRB

[Ket75] J. R. Kettenring. Book review: Cluster Analysis by Brian Everitt. Technometrics, 17(2):275, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268366.

Keviczky:1975:DEI

[Kev75] L. Keviczky. “Design of experiments” for the identification of linear dynamic systems. Technometrics, 17(3):303–308, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268065.

Khatri:1974:BRB

[Kha74] C. G. Khatri. Book review: Multivariate Analysis by Anant M. Kshirsagar. Technometrics, 16(1):138–139, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267507.

Kinory:1978:BRB

[Kin78] Shimshon Kinory. Book review: Statistical Analysis for Decision Making by Morris Hamburg. Technometrics, 20(1):104–105, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268168.

Kinory:1979:BRBa

[Kin79a] Shimshon Kinory. Book review: Statistics by David Freedman; Robert Pisani; Roger Purves. Technometrics, 21(2):273–274, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268533.

Kinory:1979:BRBb

[Kin79b] Shimshon Kinory. Book reviews: How to Use (And Misuse) Statistics by Gregory A. Kimble; Business Statistics by Paul R. Winn; Ross H. Johnson. Technometrics, 21(3):391, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267773.

Kitagawa:1979:UAD

[Kit79] Genshiro Kitagawa. On the use of AIC for the detection of outliers. Technometrics, 21(2):193–199, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1268516. See corrigenda [Kit81].

Kitagawa:1981:CUA

[Kit81] Genshiro Kitagawa. Corrigenda: “On the use of AIC for the detection of outliers” [Technometrics 21 (1979), no. 2, 193–199; MR 80f:62026]. Technometrics, 23(3):320–321, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267810. See [Kit79].

Kao:1979:SPC

[KKM79] Peichuen Kao, Edward P. C. Kao, and Jack M. Mogg. A simple procedure for computing performance characteristics of truncated sequential tests with exponential lifetimes. Technometrics, 21(2):229–232, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268520.

Kelley:1976:EEE

[KKS76] Gary D. Kelley, Judy A. Kelley, and W. R. Schucany. Efficient estimation of P (Y < X) in the exponential case. Technometrics, 18(3):359– 360, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268746.

Kleijnen:1975:SDP

[Kle75] Jack P. C. Kleijnen. Screening designs for poly-factor experimentation. Technometrics, 17(4):487–493, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268437.

Kleiner:1977:BRB

[Kle77] Beat Kleiner. Book review: Time Series Analysis: Forecasting and Control by G. E. P. Box; G. J. Jenkins. Technometrics, 19(3):343–344, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267710.

Khuri:1979:MRA

[KM79] A. I. Khuri and Raymond H. Myers. Modified ridge analysis. Technometrics, 21(4):467–473, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268285.

Kubicek:1971:QR

[KME71] M. Kubicek, M. Marek, and E. Eckert. Quasilinearized regression. Technometrics, 13(3):601–608, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267171.

Kaminsky:1974:PIE

[KN74] Kenneth S. Kaminsky and Paul I. Nelson. Prediction intervals for the exponential distribution using subsets of the data. Technometrics, 16(1):57–59, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267492.

Knox:1971:BRB

[Kno71] Hugh Knox. Book review: Markets, Prices, and Interregional Trade by Raymond G. Bressler,; Richard King. Technometrics, 13(1):211–212, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267092.

Kase:1977:GAD

[KO77] Shigeo Kase and Hiroshi Ohta. GERT analysis of the Dodge–Romig sampling inspection plan. Technometrics, 19(3):249–257, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267694.

Kowalski:1970:PSR

[Kow70] Charles J. Kowalski. The performance of some rough tests for bivariate normality before and after coordinate transformations to normality. Technometrics, 12(3):517–544, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267200.

Kowalski:1971:OAF

[Kow71] Charles J. Kowalski. The OC and ASN functions of some SPRT’s for the correlation coefficient. Technometrics, 13(4):833–841, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266958.

Koyama:1978:MSR

[Koy78] Takeshi Koyama. Modified switching rules for sampling schemes such as MIL-STD-105D. Technometrics, 20(1):95–102, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268165.

Koyama:1979:SCU

[Koy79] Takeshi Koyama. Switching characteristics under MIL-STD-105D. Technometrics, 21(1):9–19, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268576.

Kumar:1971:TCT

[KP71] S. Kumar and H. I. Patel. A test for the comparison of two exponential distributions. Technometrics, 13(1):183–189, February 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267085. See errata [Ano71g].

Kraemer:1979:CAN

[KP79] Helena Chmura Kraemer and Minja Paik. A central t approximation to the noncentral t distribution. Technometrics, 21(3):357–360, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267759.

Kraemer:1976:BRB

[Kra76] Helena Chmura Kraemer. Book review: Asymptotic Theory of Rank Tests for Independence by F. H. Ruymgaart. Technometrics, 18(1):122, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267928.

Krutchkoff:1970:LE

[Kru70] Richard G. Krutchkoff. Letters to the Editor. Technometrics, 12(2):433–434, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267253.

Krzanowski:1977:PFL

[Krz77] W. J. Krzanowski. The performance of Fisher’s linear discriminant function under non-optimal conditions. Technometrics, 19(2):191– 200, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268629.

Krzanowski:1979:SEP

[Krz79] W. J. Krzanowski. Some exact percentage points of a statistic useful in analysis of variance and principal component analysis. Technometrics, 21(2):261–263, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268526.

Kale:1971:EEL

[KS71] B. K. Kale and S. K. Sinha. Estimation of expected life in the presence of an outlier observation. Technometrics, 13(4):755–759, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266952.

Kupper:1971:BRB

[Kup71] Lawrence L. Kupper. Book review: Probability, Statistics, and Decision for Civil Engineers by Jack R. Benjamin; C. Allin Cornell. Technometrics, 13(1):211, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267091.

Kussmaul:1979:ACI

[Kus79] Keith Kussmaul. An approximate confidence interval for the mean of the dependent variable in a random one-way classification with covariates. Technometrics, 21(1):111–117, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268587.

LaBrecque:1977:GFT

[LaB77] John LaBrecque. Goodness-of-fit tests based on nonlinearity in probability plots. Technometrics, 19(3):293–306, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267699.

Lachenbruch:1971:BRBb

[Lac71a] P. A. Lachenbruch. Book review: Data Representation by R. S. Anderssen; M. R. Osborne. Technometrics, 13(4):925, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266978.

Lachenbruch:1971:BRBa

[Lac71b] P. A. Lachenbruch. Book review: Statistics: Probability Inference and Decision, Volume 1 by W. L. Hays; R. L. Winkler. Technometrics, 13(4):920, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266974.

Lachenbruch:1972:BRBb

[Lac72a] Peter A. Lachenbruch. Book review: Mathematical Statistics by John Freund. Technometrics, 14(3):807, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267312.

Lachenbruch:1972:BRBa

[Lac72b] Peter A. Lachenbruch. Book review: Statistical Methods in Medical Research by P. Armitage. Technometrics, 14(1):235, February 1972.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266938.

Lachenbruch:1974:BRB

[Lac74a] Peter A. Lachenbruch. Book review: Statistics by Example by F. Mosteller; W. H. Kruskal; R. F. Link; R. S. Pieters; G. R. Rising. Technometrics, 16(2):326, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267959.

Lachenbruch:1974:DAW

[Lac74b] Peter A. Lachenbruch. Discriminant analysis when the initial samples are misclassified II: Non-random misclassification models. Technometrics, 16(3):419–424, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267672.

Lachenbruch:1979:NIM

[Lac79] Peter A. Lachenbruch. Note on initial misclassification effects on the quadratic discriminant function. Technometrics, 21(1):129–132, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268589.

Ladany:1977:DSS

[Lad77] Shaul P. Ladany. Derivation of single-stage attribute sampling plans with one of the p’s larger than 0.5. Technometrics, 19(1):99–100, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268262.

LaMotte:1978:BLE

[LaM78] Lynn Roy LaMotte. Bayes linear estimators. Technometrics, 20(3):281–290, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268137.

Land:1972:EAC

[Lan72] Charles E. Land. An evaluation of approximate confidence interval estimation methods for lognormal means. Technometrics, 14(1):145–158, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266926.

Land:1973:BRB

[Lan73] Charles E. Land. Book review: Probability and Mathematical Statistics: An Introduction by Eugene Lukacs. Technometrics, 15 (4):941–942, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267409.

Landis:1975:BRB

[Lan75] J. Richard Landis. Book review: Introductory Statistical Analysis by T. W. Anderson; S. L. Sclove. Technometrics, 17(1):143, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268022.

Lawless:1971:PPC

[Law71] J. F. Lawless. A prediction problem concerning samples from the exponential distribution, with application in life testing. Technometrics, 13(4):725–730, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266949.

Lawless:1973:ESL

[Law73] J. F. Lawless. On the estimation of Safe life when the underlying life distribution is Weibull. Technometrics, 15(4):857–865, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267395.

Lawless:1975:CTB

[Law75a] J. F. Lawless. Construction of tolerance bounds for the extreme-value and Weibull distributions. Technometrics, 17(2):255–261, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268359.

Lawrence:1975:BRB

[Law75b] Clive J. Lawrence. Book review: Robust Estimates of Location: Survey and Advances by D. F. Andrews; P. J. Bickel; F. R. Hampel; P. J. Huber; W. H. Rogers; J. W. Tukey. Technometrics, 17(1):138–139, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268015.

Lawton:1976:NE

[Law76] William H. Lawton. New Editor. Technometrics, 18(4):367, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268650.

Lawless:1977:PIT

[Law77a] J. F. Lawless. Prediction intervals for the two parameter exponential distribution. Technometrics, 19(4):469–472, November 1977. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267887.

Lawless:1977:DAO

[Law77b] Jerald F. Lawless. Discussion of: “Asymptotically Optimum Over-Stress Tests to Estimate the Survival Probability at a Condition with a Low Expected Failure Probability”. Technometrics, 19(4):401–404, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267878.

See [MH77].

Lawless:1978:CIE

[Law78] J. F. Lawless. Confidence interval estimation for the Weibull and extreme value distributions. Technometrics, 20(4):355–364, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267633. See discus-

sion [Eas78].

Leslie:1970:DQA

[LC70] R. T. Leslie and D. Culpin. Distribution of quasimidranges and associated mixtures. Technometrics, 12(2):311–325, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267236.

Lorden:1973:DFR

[LE73] G. Lorden and I. Eisenberger. Detection of failure rate increases. Technometrics, 15(1):167–175, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266833.

Lee:1979:LE

[Lee79] Martin Lee. Letter to the Editor. Technometrics, 21(2):275, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268536.

Lehnen:1971:BRB

[Leh71] Robert G. Lehnen. Book review: Mathematics in the Behavioral and Social Sciences by John W. Bishir; Donald W. Drewes. Technometrics, 13(1):213, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267094.

Lemon:1974:PDM

[Lem74] Glen H. Lemon. A problem in the distribution of maximum flaw length after inspection. Technometrics, 16(4):577–583, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267610.

Lemon:1975:MLE

[Lem75] Glen H. Lemon. Maximum likelihood estimation for the three parameter Weibull distribution based on censored samples. Technometrics, 17(2):247–254, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268358.

Leone:1970:QTH

[Leo70] F. C. Leone. Query 1: [testing the hypothesis whether the two methods of testing used on the same set of batches are in agreement or not]. Technometrics, 12(2):417–418, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267248.

Leonard:1975:BAL

[Leo75] Tom Leonard. A Bayesian approach to the linear model with unequal variances. Technometrics, 17(1):95–102, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268006.

Levy:1970:CIE

[Lev70] Paul S. Levy. Combining independent estimators: An empirical sampling study. Technometrics, 12(1):162–165, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267362.

Lever:1979:BRB

[Lev79] W. E. Lever. Book review: Material Accountability Theory, Verification, Applications by Rudolf Avenhaus. Technometrics, 21 (1):134–135, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268592.

Lewis:1979:RAN

[LF79] T. Lewis and N. R. J. Fieller. A recursive algorithm for null distributions for outliers: I. Gamma samples. Technometrics, 21(3):371–376, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267762.

LaMotte:1970:CES

[LH70] L. R. LaMotte and R. R. Hocking. Computational efficiency in the selection of regression variables. Technometrics, 12(1):83–93, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267353.

Lauer:1974:PCT

[LH74] G. N. Lauer and C. P. Han. Power of Cochran’s test in the Behrens-Fisher problem. Technometrics, 16(4):545–549, November 1974. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267606.

Lee:1977:SML

[LH77] Austin F. S. Lee and Sylva M. Heghinian. A shift of the mean level in a sequence of independent normal random variables—a Bayesian approach. Technometrics, 19(4):503–506, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267892.

Likes:1974:POO

[Lik74] J. Likeš. Prediction of s th ordered observation for the two-parameter exponential distribution. Technometrics, 16(2):241–244, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267945.

Lin:1974:NTE

[Lin74] Chi-Yuan Lin. Numerical techniques for evaluating sample information. Technometrics, 16(3):447–454, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267676.

Links:1975:BRB

[Lin75] Graham Links. Book review: Linear and Integer Programming by S. Zionts. Technometrics, 17(3):386–387, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268086.

Lindstrom:1979:MSM

[Lin79] F. T. Lindstrom. A modified 3-spline method for evaluating the Euler digamma function. Technometrics, 21(3):307–311, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267752.

Lee:1978:SRI

[LL78] Larry Lee and S. Keith Lee. Some results on inference for the Weibull process. Technometrics, 20(1):41–45, February 1978. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268160.

Larsen:1972:UPR

[LM72] Wayne A. Larsen and Susan J. McCleary. The use of partial residual plots in regression analysis. Technometrics, 14(3):781–790, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267305.

Leone:1961:FND

[LNN61] F. C. Leone, L. S. Nelson, and R. B. Nottingham. The folded normal distribution. Technometrics, 3(4):543–550, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266560. See corrigenda [Ano80b].

Li:1979:TSS

[LO79] Loretta Li and Donald B. Owen. Two-sided screening procedures in the bivariate case. Technometrics, 21(1):79–85, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268583.

Low:1970:AMC

[Low70] Leone Low. An application of majorization to comparison of variances. Technometrics, 12(1):141–145, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267358.

Lowerre:1974:MSE

[Low74] James M. Lowerre. On the mean square error of parameter estimates for some biased estimators. Technometrics, 16(3):461–464, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267678.

Low:1976:SPV

[Low76] Leone Y. Low. Some properties of variance component estimators in the two-way classification. Technometrics, 18(1):39–46, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267914.

Levenbach:1976:FTT

[LR76] Hans Levenbach and Blake E. Reuter. Forecasting trending time series with relative growth rate models. Technometrics, 18(3):261– 272, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268735.

Littell:1978:CRL

[LR78] Ramon C. Littell and P. V. Rao. Confidence regions for location and scale parameters based on the Kolmogorov-Smirnov goodness of fit statistic. Technometrics, 20(1):23–27, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268156.

Lawton:1971:ELP

[LS71a] William H. Lawton and Edward A. Sylvestre. Elimination of linear parameters in nonlinear regression. Technometrics, 13(3):461–467, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267160.

Lawton:1971:SMC

[LS71b] William H. Lawton and Edward A. Sylvestre. Self modeling curve resolution. Technometrics, 13(3):617–633, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267173.

Lawton:1972:SMN

[LSM72] W. H. Lawton, E. A. Sylvestre, and M. S. Maggio. Self modeling nonlinear regression. Technometrics, 14(3):513–532, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267281.

Layard:1976:LE

[LSP76] M. W. J. Layard, David S. Salsburg, and P. Prescott. Letters to the Editor. Technometrics, 18(1):124–125, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267932.

Lawton:1979:SCM

[LSYF79] W. H. Lawton, E. A. Sylvestre, and B. J. Young-Ferraro. Statistical comparison of multiple analytic procedures: Application to clinical chemistry. Technometrics, 21(4):397–409, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268278. See discussion [FH79].

Lucas:1973:MVM

[Luc73] James M. Lucas. A modified “v” mask control scheme. Technometrics, 15(4):833–847, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267393.

Lucas:1974:OCD

[Luc74] James M. Lucas. Optimum composite designs. Technometrics, 16(4):561–567, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267608.

Lucas:1976:WRS

[Luc76] James M. Lucas. Which response surface design is best: A performance comparison of several types of quadratic response surface designs in symmetric regions. Technometrics, 18(4):411–417, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268656.

Lucas:1978:DOF

[Luc78] James M. Lucas. Discussion of: “D-Optimal Fractions of Three-Level Factorial Designs”. Technometrics, 20(4):381–382, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267636. See [MB78a].

Lukacs:1973:BRB

[Luk73] Eugene Lukacs. Book review: Preservation of Infinite Divisibility under Mixing by F. W. Steutel. Technometrics, 15(3):650–651, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266875.

Lund:1975:TAT

[Lun75] Richard E. Lund. Tables for an approximate test for outliers in linear models. Technometrics, 17(4):473–476, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268434.

Lutz:1973:BRB

[Lut73] Donald S. Lutz. Book review: Paradoxes of Rationality: Theory of Metagames and Political Behavior by Nigel Howard. Technometrics, 15(3):652, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266876.

Lawton:1973:FNP

[LvNN73] William H. Lawton, John van Ness, and Judith Nelson. [Formulation of a Nonlinear Predictor]: Discussion no. 1. Technometrics, 15(1):13– 17, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266820.

See [NV73b].

Luers:1971:PSL

[LW71] James K. Luers and Robert H. Wenning. Polynomial smoothing: Linear vs cubic. Technometrics, 13(3):589–600, August 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267170.

Lynch:1978:BRB

[Lyn78] John K. Lynch. Book review: Applications of Statistics to Industrial Experimentation by C. Daniel. Technometrics, 20(2):209–210, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268716.

Muse:1978:CDE

[MA78] H. D. Muse and R. L. Anderson. Comparison of designs to estimate variance components in a two-way classification model. Technometrics, 20(2):159–166, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268708. See corrigenda [Ano79e].

MacGregor:1976:OCS

[Mac76] John F. MacGregor. Optimal choice of the sampling interval for discrete process control. Technometrics, 18(2):151–160, May 1976. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267518. See erratum [Ano77h].

Mallows:1973:SC

[Mal73] C. L. Mallows. Some comments on C_{P}. Technometrics, 15 (4):661–675, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267380.

Mallows:1974:FPS

[Mal74] C. L. Mallows. [The Fitting of Power Series, Meaning Polynomials, Illustrated on Band-Spectroscopic Data]: Discussion no. 1. Technometrics, 16(2):187–188, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267937. See [BT74].

Mann:1970:EEC

[Man70a] Nancy R. Mann. Estimators and exact confidence bounds for Weibull parameters based on a few ordered observations. Technometrics, 12(2):345–361, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267239.

Mantel:1970:WSP

[Man70b] Nathan Mantel. Why stepdown procedures in variable selection. Technometrics, 12(3):621–625, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267207.

Mandel:1971:NAV

[Man71a] John Mandel. A new analysis of variance model for non-additive data. Technometrics, 13(1):1–18, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267072.

Mann:1971:BLI

[Man71b] Nancy R. Mann. Best linear invariant estimation for Weibull parameters under progressive censoring. Technometrics, 13(3):521–533, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267165.

Mann:1971:BRB

[Man71c] W. Robert Mann. Book review: Green’s Functions; Introductory Theory with Applications by G. F. Roach. Technometrics, 13(3):708, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267187.

Mantel:1971:MVS

[Man71d] Nathan Mantel. More on variable selection and an alternative approach. Technometrics, 13(2):455–457, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266815.

Mann:1972:DSL

[Man72] Nancy R. Mann. Design of over-stress life-test experiments when failure times have the two-parameter Weibull distribution. Technometrics, 14(2):437–451, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267434.

Manson:1973:BRB

[Man73] A. R. Manson. Book review: Response Surface Methodology by Raymond H. Myers. Technometrics, 15(4):936–938, November 1973. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267405.

Mantel:1975:PPE

[Man75] Nathan Mantel. Poisson parameter estimation with arbitrary factors of proportionality and arbitrary points of left truncation. Technometrics, 17(2):173–176, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268348.

Marquardt:1970:GIR

[Mar70] Donald W. Marquardt. Generalized inverses, ridge regression, biased linear estimation, and nonlinear estimation. Technometrics, 12(3):591–612, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267205.

Marcus:1973:SMM

[Mar73] Allan H. Marcus. A stochastic model of microscale air pollution from highway traffic. Technometrics, 15(2):353–363, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266994. See erratum [Ano73j].

Marquardt:1974:FPS

[Mar74] Donald W. Marquardt. [The Fitting of Power Series, Meaning Polynomials, Illustrated on Band-Spectroscopic Data]: Discussion no. 2.

Technometrics, 16(2):189–192, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267938. See [BT74].

Marcus:1975:BRB

[Mar75] Allan H. Marcus. Book review: Systems Analysis and Design Using Network Techniques by Gary E. Whitehouse. Technometrics, 17(2):274–275, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268365.

Margolin:1976:DAF

[Mar76] Barry H. Margolin. Design and analysis of factorial experiments via interactive computing in APL. Technometrics, 18(2):135–150, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267517.

Matz:1978:MLP

[Mat78] A. W. Matz. Maximum likelihood parameter estimation for the quartic exponential distribution. Technometrics, 20(4):475–484, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267648.

Marcus:1974:SSP

[MB74] Richard Marcus and Saul Blumenthal. A sequential screening procedure. Technometrics, 16(2):229–234, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267943.

Mitchell:1978:OFTa

[MB78a] T. J. Mitchell and C. K. Bayne. D-optimal fractions of three-level factorial designs. Technometrics, 20(4):369–380, November 1978. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267635. See discussion [Luc78].

Mitchell:1978:OFTb

[MB78b] T. J. Mitchell and C. K. Bayne. [D-Optimal Fractions of Three-Level Factorial Designs]: Response. Technometrics, 20(4):383, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267637.

McKay:1979:CTM

[MBC79] M. D. McKay, R. J. Beckman, and W. J. Conover. A comparison of three methods for selecting values of input variables in the analysis of output from a computer code. Technometrics, 21(2):239–245, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268522.

Maynard:1972:APT

[MC72] J. M. Maynard and B. Chow. An approximate Pitman-type “close” estimator for the negative binomial parameter p. Technometrics, 14(1):77–88, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266920.

Myers:1973:RST

[MC73] Raymond H. Myers and Walter H. Carter, Jr. Response surface techniques for dual response systems. Technometrics, 15(2):301–317, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266990.

McClave:1972:BRB

[McC72] James T. McClave. Book review: Introduction to Statistics: A Fresh Approach by G. Noether. Technometrics, 14(3):805–806, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267310.

McCabe:1975:CVS

[McC75a] George P. McCabe, Jr. Computations for variable selection in discriminant analysis. Technometrics, 17(1):103–109, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268007.

McClave:1975:SA

[McC75b] James McClave. Subset autoregression. Technometrics, 17(2):213– 220, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268353.

McCabe:1978:ERC

[McC78a] George P. McCabe, Jr. Evaluation of regression coefficient estimates using α-acceptability. Technometrics, 20(2):131–139, May 1978. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268704.

McClave:1978:BRB

[McC78b] James T. McClave. Book review: Introduction to Statistical Time Series by Wayne A. Fuller. Technometrics, 20(2):211, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268718.

McDonald:1975:DRA

[McD75] Gary C. McDonald. Discussion of: “Ridge Analysis Following a Preliminary Test of the Shrunken Hypothesis”. Technometrics, 17 (4):443–445, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268430. See [Obe75].

McDonald:1976:DVB

[McD76] Gary C. McDonald. The distribution of a variate based on independent ranges from a uniform population. Technometrics, 18(3):343– 349, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268744.

McDonald:1979:NSP

[McD79] Gary C. McDonald. Nonparametric selection procedures applied to state traffic fatality rates. Technometrics, 21(4):515–523, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268291. See corri-

gendum [Ano80c].

McKay:1976:SPD

[McK76] Ronald J. McKay. Simultaneous procedures in discriminant analysis involving two groups. Technometrics, 18(1):47–53, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267915.

McKay:1979:AVS

[McK79] R. J. McKay. The adequacy of variable subsets in multivariate regression. Technometrics, 21(4):475–479, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268286.

McLachlan:1972:ARD

[McL72] G. J. McLachlan. Asymptotic results for discriminant analysis when the initial samples are misclassified. Technometrics, 14(2):415–422, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267432.

McLachlan:1974:EEM

[McL74a] G. J. McLachlan. Estimation of the errors of misclassification on the criterion of asymptotic mean square error. Technometrics, 16(2):255– 260, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267948.

McLachlan:1974:RTA

[McL74b] G. J. McLachlan. The relationship in terms of asymptotic mean square error between the separate problems of estimating each of the three types of error rate of the linear discriminant function. Technometrics, 16(4):569–575, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267609.

Mudholkar:1976:SAN

[MCL76] Govind S. Mudholkar, Yogendra P. Chaubey, and Ching-Chuong Lin. Some approximations for the noncentral-f distribution. Technometrics, 18(3):351–358, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268745.

McMillan:1971:TOTb

[McM71] R. G. McMillan. Tests for one or two outliers in normal samples with unknown variance. Technometrics, 13(1):87–100, February 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267077. See correction [MM73].

Maxim:1975:OAT

[MCM75] L. Daniel Maxim, Daniel E. Cullen, and John G. Mardo. Optimal acceptance test plans with grouping. Technometrics, 17(3):315–320, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268067.

McMillan:1971:TOTa

[MD71] R. G. McMillan and H. A. David. Tests for one of two outliers in normal samples with known variance. Technometrics, 13(1):75–85, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267076.

McDonald:1977:NUT

[MDM77] Lyman L. McDonald, Bruce M. Davis, and George A. Milliken. A nonrandomized unconditional test for comparing two proportions in 2 × 2 contigency tables. Technometrics, 19(2):145–157, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268624.

Merten:1972:BRB

[Mer72] Alan G. Merten. Book review: Computer Simulation Experiments with Models of Economic Systems by Thomas H. Naylor. Technometrics, 14(2):504–505, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267443.

Mergenthaler:1977:BRB

[Mer77] Mitchell J. Mergenthaler. Book review: Introduction to Statistics: A Nonparametric Approach by Gottfried E. Noether. Technometrics, 19(3):342–343, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267708.

Mergenthaler:1979:BRB

[Mer79] Mitchell J. Mergenthaler. Book review: Nonparametrics: Statistical Methods Based on Ranks by E. L. Lehmann. Technometrics, 21(2):272–273, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268531.

Mann:1973:TOW

[MF73] Nancy R. Mann and Kenneth W. Fertig. Tables for obtaining Weibull confidence bounds and tolerance bounds based on best linear invariant estimates of parameters of the extreme-value distribution. Technometrics, 15(1):87–101, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266827.

Mann:1975:GFT

[MF75a] Nancy R. Mann and Kenneth W. Fertig. A goodness-of-fit test for the two parameter vs. three parameter Weibull; confidence bounds for threshold. Technometrics, 17(2):237–245, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268357.

Mann:1975:SEP

[MF75b] Nancy R. Mann and Kenneth W. Fertig. Simplified efficient point and interval estimators for Weibull parameters. Technometrics, 17(3):361–368, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268075.

Mann:1977:EUQ

[MF77] Nancy R. Mann and Kenneth W. Fertig. Efficient unbiased quantile estimators for moderate-size complete samples from extreme-value and Weibull distributions; confidence bounds and tolerance and prediction intervals. Technometrics, 19(1):87–93, February 1977. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268260.

Mann:1974:AOC

[MG74] Nancy R. Mann and Frank E. Grubbs. Approximately optimum confidence bounds for system reliability based on component test data. Technometrics, 16(3):335–347, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267663.

McGilchrist:1972:NSS

[MH72] C. A. McGilchrist and S. H. Huxham. A note on some simple time series estimates. Technometrics, 14(1):223–227, February 1972. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266934.

Meeker:1977:AOS

[MH77] William Q. Meeker, Jr. and Gerald J. Hahn. Asymptotically optimum over-stress tests to estimate the survival probability at a condition with a low expected failure probability. Technometrics, 19(4):381–399, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1267877. See discussion [Law77b].

Meeker:1978:CAT

[MH78] William Q. Meeker, Jr. and Gerald J. Hahn. A comparison of accelerated test plans to estimate the survival probability at a design stress. Technometrics, 20(3):245–247, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268132.

Maxim:1977:EDS

[MHC77] L. Daniel Maxim, Arlo D. Hendrickson, and Daniel E. Cullen. Experimental design for sensitivity testing: The Weibull model. Technometrics, 19(4):405–412, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1267879. See discussion [Ant77].

Moussa-Hamouda:1974:BEC

[MHL74] Effat Moussa-Hamouda and Fred C. Leone. The O-blue estimators for complete and censored samples in linear regression. Technometrics, 16(3):441–446, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267675.

Moussa-Hamouda:1977:EOL

[MHL77a] Effat Moussa-Hamouda and Fred C. Leone. Efficiency of ordinary least squares estimators from trimmed and Winsorized samples in linear regression. Technometrics, 19(3):265–273, August 1977. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267696.

Moussa-Hamouda:1977:REA

[MHL77b] Effat Moussa-Hamouda and Fred C. Leone. The robustness of efficiency of adjusted trimmed estimators in linear regression. Technometrics, 19(1):19–34, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268250.

Mielke:1974:SRT

[Mie74] Paul W. Mielke, Jr. Squared rank test appropriate to weather modification cross-over design. Technometrics, 16(1):13–16, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267486.

Mihram:1975:BRB

[Mih75] G. Arthur Mihram. Book review: Concepts and Methods in Discrete Event Digital Simulation by George S. Fishman. Technometrics, 17(2):272–274, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268364. See corrigenda [Mih77].

Mihram:1977:CBR

[Mih77] G. Arthur Mihram. Corrigenda: Book Review: Concepts and Methods in Discrete Event Digital Simulation by George S. Fishman. Technometrics, 19(2):222, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268644. See [Mih75].

Miller:1972:BRB

[Mil72a] M. Clinton Miller III. Book review: Introductory Statistics and Probability by David W. Blakeslee; William G. Chinn. Technometrics, 14(1):236, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266939.

Milton:1972:CEM

[Mil72b] Roy C. Milton. Computer evaluation of the multivariate normal integral. Technometrics, 14(4):881–889, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267136.

Mitchell:1974:ACO

[Mit74a] Toby J. Mitchell. An algorithm for the construction of “D-optimal” experimental designs. Technometrics, 16(2):203–210, May 1974. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267940.

Mitchell:1974:CCO

[Mit74b] Toby J. Mitchell. Computer construction of “D-optimal” firstorder designs. Technometrics, 16(2):211–220, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267941.

Martinelle:1970:CRL

[MK70] Sten Martinelle and R. G. Krutchkoff. On the choice of regression in linear calibration. Comments on a paper by R. G. Krutchkoff. Technometrics, 12(1):157–161, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267361.

Myers:1975:GRS

[ML75] Raymond H. Myers and Steve J. Lahoda. A generalization of the response surface mean square error criterion with a specific application to the scope. Technometrics, 17(4):481–486, November 1975.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268436.

Moran:1973:TOT

[MM73] M. A. Moran and R. G. McMillan. Tests for one or two outliers in normal samples with unknown variance: A correction. Technometrics, 15(3):637–640, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266870. See [McM71].

Meeker:1977:WVC

[MN77] William Q. Meeker, Jr. and Wayne Nelson. Weibull variances and confidence limits by maximum likelihood for singly censored data. Technometrics, 19(4):473–476, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267888.

Moeschberger:1972:BRB

[Moe72] M. L. Moeschberger. Book review: Basic Probability Theory by Robert B. Ash. Technometrics, 14(4):983, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267154.

Moeschberger:1974:LTU

[Moe74] M. L. Moeschberger. Life tests under dependent competing causes of failure. Technometrics, 16(1):39–47, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267490.

Miller:1972:LE

[MP72] Alan J. Miller and S. A. Patil. Letters to the Editor. Technometrics, 14(2):507–508, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267444.

Meeter:1970:CTM

[MPB70] Duane Meeter, Walter Pirie, and William Blot. A comparison of two model-discrimination criteria. Technometrics, 12(3):457–470, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267196.

McLean:1979:SMM

[MPBD79] D. D. McLean, D. J. Pritchard, D. W. Bacon, and J. Downie. Singularities in multiresponse modelling. Technometrics, 21(3):291–298, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267750.

Moberg:1978:AME

[MRR78] Thomas F. Moberg, John S. Ramberg, and Ronald H. Randles. An adaptive m-estimator and its application to a selection problem. Technometrics, 20(3):255–263, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268134.

McDonald:1973:IRE

[MS73] Gary C. McDonald and Richard C. Schwing. Instabilities of regression estimates relating air pollution to mortality. Technometrics, 15(3):463–481, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266852.

Marquardt:1974:TSM

[MS74] Donald W. Marquardt and Ronald D. Snee. Test statistics for mixture models. Technometrics, 16(4):533–537, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267604.

Michael:1979:NAT

[MS79] John R. Michael and William R. Schucany. A new approach to testing goodness of fit for censored samples. Technometrics, 21(4):435–441, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268282.

Morgan:1972:CRS

[MT72] J. A. Morgan and J. F. Tatar. Calculation of the residual sum of squares for all possible regressions. Technometrics, 14(2):317–325, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267424.

Muller:1970:CID

[Mul70] Mervin E. Muller. Computers as an instrument for data analysis. Technometrics, 12(2):259–293, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267233.

Mayer:1973:BEL

[MW73] Lawrence S. Mayer and Thomas A. Willke. On biased estimation in linear models. Technometrics, 15(3):497–508, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266855.

McGilchrist:1975:NDF

[MW75a] C. A. McGilchrist and K. D. Woodyer. Note on a distribution-free CUSUM technique. Technometrics, 17(3):321–325, August 1975. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268068.

Minder:1975:LAL

[MW75b] Ch. E. Minder and J. B. Whitney. A likelihood analysis of the linear calibration problem. Technometrics, 17(4):463–471, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268433.

Martz:1978:BMD

[MW78] H. F. Martz, Jr. and M. S. Waterman. A Bayesian model for determining the optimal test stress for a single test unit. Technometrics, 20(2):179–185, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268711.

Myers:1974:BRB

[Mye74] Raymond H. Myers. Book review: Elementary Applied Statistics by Marvin Lentner. Technometrics, 16(1):137, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267504.

Nadas:1970:EDR

[Nád70a] Arthur Nádas. On estimating the distribution of a random vector when only the smallest coordinate is observable. Technometrics, 12(4):923–924, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267338.

Nadas:1970:PHF

[Nád70b] Arthur Nádas. On proportional hazard functions. Technometrics, 12(2):413–416, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267247.

Nadas:1971:DIM

[Nád71] Arthur Nádas. The distribution of the identified minimum of a normal pair determines the distribution of the pair. Technometrics, 13(1):201–202, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267088.

Nadas:1973:BTZ

[Nád73] Arthur Nádas. Best tests for zero drift based on first passage times in Brownian motion. Technometrics, 15(1):125–132, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266830.

Naik:1974:TME

[Nai74] Umesh D. Naik. On tests of main effects and interactions in higherway layouts in the analysis of variance random effects model. Technometrics, 16(1):17–25, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267487.

Naszodi:1978:EBC

[Nas78] László J. Naszódi. Elimination of the bias in the course of calibration. Technometrics, 20(2):201–205, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268714.

Naylor:1972:BRB

[Nay72] Thomas H. Naylor. Book review: Computer Simulation of Competitive Market Response by Arnold E. Amstutz. Technometrics, 14(3):803–804, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267308.

Neave:1970:EFR

[Nea70] Henry R. Neave. Extending the frequency range of spectrum estimates by the use of two data recorders. Technometrics, 12 (4):877–890, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267332.

Nelson:1972:SLT

[Nel72a] Wayne Nelson. A short life test for comparing a sample with previous accelerated test results. Technometrics, 14(1):175–185, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266929.

Nelson:1972:TAH

[Nel72b] Wayne Nelson. Theory and applications of hazard plotting for censored failure data. Technometrics, 14(4):945–966, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267144.

Nelson:1973:ARC

[Nel73] Wayne Nelson. Analysis of residuals from censored data. Technometrics, 15(4):697–715, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267382.

Nelson:1974:BRB

[Nel74] Judith Zeh Nelson. Book review: Digital Time Series Analysis by Robert K. Otnes; Loren Enochson. Technometrics, 16(3):479–480, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267684.

Nelson:1975:SRG

[Nel75] David L. Nelson. Some remarks on generalizations of the negative binomial and Poisson distributions. Technometrics, 17(1):135–136, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268013.

Ng:1976:NBT

[Ng76] Vee Ming Ng. A note on the best test for discrimination between exponentiality and uniformity. Technometrics, 18(2):237–238, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267529.

Nalimov:1970:PUC

[NGM70] V. V. Nalimov, T. I. Golikova, and N. G. Mikeshina. On practical use of the concept of D-optimality. Technometrics, 12(4):799–812, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267326.

Nelson:1972:LER

[NH72] Wayne Nelson and Gerald J. Hahn. Linear estimation of a regression relationship from censored data. Part I: Simple methods and their application. Technometrics, 14(2):247–269, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267418. See discussion [Zel72].

Nelson:1973:LER

[NH73] Wayne Nelson and Gerald J. Hahn. Linear estimation of a regression relationship from censored data: Part II best linear unbiased estimation and theory. Technometrics, 15(1):133–150, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266831.

Nelson:1976:TOC

[NK76] Wayne Nelson and Thomas J. Kielpinski. Theory for optimum censored accelerated life tests for normal and lognormal life distributions. Technometrics, 18(1):105–114, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267923.

Narula:1975:SSC

[NL75] Subhash C. Narula and Franz S. Li. Sample size calculations in exponential life testing. Technometrics, 17(2):229–231, May 1975. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268355.

Nelson:1978:TOA

[NM78] Wayne Nelson and William Q. Meeker. Theory for optimum accelerated censored life tests for Weibull and extreme value distributions. Technometrics, 20(2):171–177, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268710.

Nelson:1979:ILN

[NS79] Wayne Nelson and Josef Schmee. Inference for (log) normal life distributions from small singly censored samples and BLUEs. Technometrics, 21(1):43–54, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268579.

Nelson:1973:CNP

[NV73a] Judith Zeh Nelson and John W. Van Ness. Choosing a nonlinear predictor. Technometrics, 15(2):219–231, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266983.

Nelson:1973:FNP

[NV73b] Judith Zeh Nelson and John W. Van Ness. Formulation of a nonlinear predictor. Technometrics, 15(1):1–12, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266819. See discussion [LvNN73].

Narula:1977:PLR

[NW77] Subhash C. Narula and John F. Wellington. Prediction, linear regression and the minimum sum of relative errors. Technometrics, 19(2):185–190, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268628. See corrigenda [Ano80b].

Narula:1979:SVL

[NW79] Subhash C. Narula and John F. Wellington. Selection of variables in linear regression using the minimum sum of weighted absolute errors criterion. Technometrics, 21(3):299–306, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267751. See corrigenda [Ano80b].

OBrien:1973:PSI

[O’B73] Peter C. O’Brien. A procedure for the sequential interval estimation of the shape parameter of the gamma distribution. Technometrics, 15(3):563–569, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266862.

Owen:1976:SMI

[OB76] D. B. Owen and J. W. Boddie. A screening method for increasing acceptable product with some parameters unknown. Technometrics, 18(2):195–199, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267522.

Obenchain:1975:RAF

[Obe75] R. L. Obenchain. Ridge analysis following a preliminary test of the shrunken hypothesis. Technometrics, 17(4):431–441, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268429. See discus-

sion [McD75].

Obenchain:1977:CTC

[Obe77] R. L. Obenchain. Classical F -tests and confidence regions for ridge regression. Technometrics, 19(4):429–439, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267882.

Odeh:1971:JST

[Ode71] Robert E. Odeh. On Jonckheere’s k-sample test against ordered alternatives. Technometrics, 13(4):912–918, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266972.

Odoroff:1977:BRB

[Odo77] Charles L. Odoroff. Book review: Discrete Multivariate Analyses: Theory and Practice by Yvonne M. M. Bishop; Stephen E. Fienberg; Paul W. Holland; Richard J. Light; Frederick Mosteller. Technometrics, 19(1):101–102, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1268263. See addendum and corrigendum [Ano77a].

Odoroff:1978:BRB

[Odo78] Charles L. Odoroff. Book review: The Analysis of Contingency Tables by B. A. Everitt. Technometrics, 20(3):319, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268142. See corrigenda [Ano79f].

OFallon:1973:BRB

[O’F73] W. M. O’Fallon. Book review: Multivariate Data Analysis by W. W. Cooley; P. R. Lohnes. Technometrics, 15(3):648–650, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266874.

Oxspring:1977:DIM

[OL77] H. H. Oxspring and L. R. LaMotte. The design of incomplete multinomial samples: The nested case. Technometrics, 19(1):59–63, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268255.

Owen:1977:SBN

[OlHS77] D. B. Owen and Yueh ling Hsiao Su. Screening based on normal variables. Technometrics, 19(1):65–68, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268256.

Ott:1972:DES

[OM72] L. Ott and W. Mendenhall. Designs for estimating the slope of a second order linear model. Technometrics, 14(2):341–353, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267426.

Olsson:1975:NMS

[ON75] Donald M. Olsson and Lloyd S. Nelson. The Nelder–Mead simplex procedure for function minimization. Technometrics, 17(1):45–51, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267998.

Pape:1972:CS

[Pap72] Elinor S. Pape. A combination of F -statistics. Technometrics, 14(1):89–99, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266921.

Park:1970:RCB

[Par70] Se-Hark Park. Regression and correlation in a bivariate distribution with different marginal densities. Technometrics, 12(3):687–691, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267212.

Parzen:1972:IAT

[Par72] Emanual Parzen. [The Inverse Autocorrelations of a Time Series and Their Applications]: Discussion no. 2. Technometrics, 14(2):295– 298, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267421.

See [Cle72b].

Park:1978:EDF

[Par78a] Sung H. Park. Experimental designs for fitting segmented polynomial regression models. Technometrics, 20(2):151–154, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268706.

Park:1978:SCA

[Par78b] Sung H. Park. Selecting contrasts among parameters in Scheffe’s mixture models: Screening components and model reduction. Technometrics, 20(3):273–279, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268136.

Passi:1979:LME

[Pas79] Ranjit M. Passi. Linear model estimation with extraneous estimates of a parameter subset, with an application to aircraft tracking. Technometrics, 21(4):511–514, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268290.

Patel:1973:QCM

[Pat73] H. I. Patel. Quality control methods for multivariate binomial and Poisson distributions. Technometrics, 15(1):103–112, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266828.

Patel:1976:CIU

[Pat76a] Jagdish K. Patel. Confidence intervals using censored data. Technometrics, 18(2):221–225, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267526.

Patel:1976:EPT

[Pat76b] Y. C. Patel. Estimation of the parameters of the triple and quadruple stuttering-Poisson distributions. Technometrics, 18(1):67–73, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267918.

Pritchard:1977:AHE

[PB77] D. J. Pritchard and D. W. Bacon. Accounting for heteroscedasticity in experimental design. Technometrics, 19(2):109–115, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268620.

Pasternack:1976:GSL

[PBT76] B. S. Pasternack, D. E. Bohning, and J. Thomas. Group-sequential leak-testing of sealed radium sources. Technometrics, 18(1):59–66, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267917.

Pritchard:1977:FCH

[PDB77] D. J. Pritchard, J. Downie, and D. W. Bacon. Further consideration of heteroscedasticity in fitting kinetic models. Technometrics, 19(3):227–236, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267692.

Pepper:1973:CIN

[Pep73] M. P. G. Pepper. A calibration of instruments with non-random errors. Technometrics, 15(3):587–599, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266865.

Pierce:1973:FFB

[Pie73] Donald A. Pierce. Fiducial, frequency, and Bayesian inference on reliability for the two parameter negative exponential distribution. Technometrics, 15(2):249–253, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266985.

Patil:1976:TPZ

[PK76] S. A. Patil and J. L. Kovner. On the test and power of zero drift on first passage times in Brownian motion. Technometrics, 18(3):341– 342, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268743.

Patil:1979:POT

[PK79] S. A. Patil and J. L. Kovner. On the power of an optimum test for the mean of the inverse Gaussian distribution. Technometrics, 21(3):379– 381, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267764.

Preece:1971:IPM

[Pre71] D. A. Preece. Iterative procedures for missing values in experiments. Technometrics, 13(4):743–753, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266951.

Prescott:1975:ATO

[Pre75] P. Prescott. An approximate test for outliers in linear models. Technometrics, 17(1):129–132, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268011.

Preston:1977:BRB

[Pre77] D. B. Preston. Book review: The Analysis of Time Series: Theory and Practice by C. Chatfield. Technometrics, 19(3):343, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267709.

Pfanzagl:1970:ESS

[PS70] J. Pfanzagl and W. Schüler. The efficiency of sequential sampling plans based on prior distributions and costs. Technometrics, 12(2):299–310, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267235.

Pettitt:1977:KSG

[PS77] A. N. Pettitt and M. A. Stephens. The Kolmogorov–Smirnov goodness-of-fit statistic with discrete and grouped data. Technometrics, 19(2):205–210, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268631.

Pope:1972:USS

[PW72] P. T. Pope and J. T. Webster. The use of an F -statistic in stepwise regression procedures. Technometrics, 14(2):327–340, May 1972. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267425.

Pandit:1977:MAC

[PW77] S. M. Pandit and S. M. Wu. Modeling and analysis of closed-loop systems from operating data. Technometrics, 19(4):477–485, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267889.

Quade:1971:BRB

[Qua71] Dana Quade. Book review: Readings in Statistics by Audrey Haber; Richard P. Runyon; Pietro Badia. Technometrics, 13(2):451, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266809.

Quade:1974:BRBb

[Qua74a] Dana Quade. Book review: The Statistics Cum Index, by James L. Dolby and John W. Tukey. Technometrics, 16(4):634, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267621.

Quade:1974:BRBa

[Qua74b] Dana Quade. Book review: The Statistics Cum Index by James L. Dolby; John W. Tukey. Technometrics, 16(4):634, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267621.

Quade:1975:BRB

[Qua75] Dana Quade. Book review: Sturdy Statistics: Nonparametrics and Order Statistics by Frederick Mosteller; Robert E. K. Rourke. Technometrics, 17(1):140–141, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268019.

Rahim:1976:PTE

[Rah76] M. A. Rahim. Pre-test estimation of simple linear regression model using prior knowledge of one of the parameters. Technometrics, 18(4):439–445, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268659.

Ramsey:1973:BRB

[Ram73] James Ramsey. Book review: Probability and Statistical Analysis by E. P. Hickman; J. G. Hilton. Technometrics, 15(2):422–423, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267005.

Randolph:1970:BRBa

[Ran70a] Paul H. Randolph. Book review: Linear Programming: Methods and Applications by Saul I. Gass. Technometrics, 12(1):174, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267367.

Randolph:1970:BRBc

[Ran70b] Paul H. Randolph. Book review: Optimal Statistical Decisions by Morris H. de Groot. Technometrics, 12(3):712–713, August 1970.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267221.

Randolph:1970:BRBb

[Ran70c] Paul H. Randolph. Book review: Spectral Analysis and Its Applications by Gwilym M. Jenkins; Donald G. Watts. Technometrics, 12(1):174–175, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267368.

Ray:1970:IFT

[Ray70] W. D. Ray. The inverse of a finite Toeplitz matrix. Technometrics, 12(1):153–156, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267360.

Rayment:1972:IPM

[Ray72] P. R. Rayment. The identification problem for a mixture of observations from two normal populations. Technometrics, 14(4):911–918, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267139.

Read:1971:RBS

[Rea71] R. R. Read. A remark on the Barlow–Scheuer reliability growth estimation scheme. Technometrics, 13(1):199–200, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267087. See [BS71].

Regier:1976:SSP

[Reg76] Mary H. Regier. Simplified selection procedures for multivariate normal populations. Technometrics, 18(4):483–489, November 1976. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268664.

Relles:1970:VRT

[Rel70] Daniel A. Relles. Variance reduction techniques for Monte Carlo sampling from Student distributions. Technometrics, 12(3):499–515, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267199.

Reynolds:1975:AAR

[Rey75] Marion R. Reynolds, Jr. Approximations to the average run length in cumulative sum control charts. Technometrics, 17(1):65–71, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268002.

Robinson:1978:ARL

[RH78] P. B. Robinson and T. Y. Ho. Average run lengths of geometric moving average charts by numerical methods. Technometrics, 20(1):85– 93, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268164.

Riffenburgh:1971:RTT

[Rif71] Robert H. Riffenburgh. Regime testimation in a time sequence. Technometrics, 13(4):795–809, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266955.

Rizvi:1971:SSP

[Riz71] M. Haseeb Rizvi. Some selection problems involving folded normal distribution. Technometrics, 13(2):355–369, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266797.

Ralston:1978:DDF

[RJ78] Mary L. Ralston and Robert I. Jennrich. Dud, a derivative-free algorithm for nonlinear least squares. Technometrics, 20(1):7–14, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268154.

Rogers:1970:BRBc

[Rog70a] G. S. Rogers. Book review: Parametric Estimation by M. T. Wasan. Technometrics, 12(3):713–714, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267223.

Rogers:1970:BRBb

[Rog70b] G. S. Rogers. Book review: Probability and Random Processes for Engineers and Scientists by A. Bruce Clarke; Ralph L. Disney. Technometrics, 12(3):712, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267220.

Rogers:1970:BRBd

[Rog70c] Gerald S. Rogers. Book review: Foundations of Probability by Alfred Renyi. Technometrics, 12(4):932, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267342.

Rogers:1970:BRBa

[Rog70d] Gerald S. Rogers. Book review: The Structure of Probability Theory with Applications by Aram J. Thomasion. Technometrics, 12(1):178–179, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267372.

Ronde:1972:BRB

[Ron72] C. A. Ronde. Book review: Review of Estimation in Linear Models by T. O. Lewis; P. L. Odell. Technometrics, 14(3):804–805, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267309.

Ronning:1977:SSG

[Ron77] Gerd Ronning. A simple scheme for generating multivariate gamma distributions with non-negative covariance matrix. Technometrics, 19(2):179–183, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268627.

Roquemore:1976:HDQ

[Roq76] K. G. Roquemore. Hybrid designs for quadratic response surfaces. Technometrics, 18(4):419–423, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268657. See errata [Ano77f].

Rosner:1975:BRB

[Ros75a] Bernard Rosner. Book review: An Introduction to Medical Statistics by H. O. Lancaster. Technometrics, 17(2):275–276, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268367.

Rosner:1975:DMO

[Ros75b] Bernard Rosner. On the detection of many outliers. Technometrics, 17(2):221–227, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268354.

Rossi:1975:AWA

[Ros75c] J. A. Dei Rossi. An application of Welch’s approximate t-solution of the Behrens–Fisher problem to confidence intervals. Technometrics, 17(1):57–60, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268000.

Rosner:1977:PPR

[Ros77] Bernard Rosner. Percentage points for the RST many outlier procedure. Technometrics, 19(3):307–312, August 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267700.

Rousu:1973:WSK

[Rou73] Dwight N. Rousu. Weibull skewness and kurtosis as a function of the shape parameter. Technometrics, 15(4):927–930, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267402.

Ran:1976:ARD

[RR76] Arne Ran and Stig I. Rosenlund. Age replacement with discounting for a continuous maintenance cost model. Technometrics, 18 (4):459–465, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268662.

Randles:1973:APS

[RRH73] Ronald H. Randles, John S. Ramberg, and Robert V. Hogg. An adaptive procedure for selecting the population with largest location parameter. Technometrics, 15(4):769–778, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267387.

Raghunandanan:1971:SEP

[RS71] K. Raghunandanan and R. Srinivasan. Simplified estimation of parameters in a double exponential distribution. Technometrics, 13(3):689–691, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267181.

Rao:1972:ENN

[RSO72] J. N. K. Rao, K. Subrahmaniam, and D. B. Owen. Effect of nonnormality on tolerance limits which control percentages in both tails of normal distribution. Technometrics, 14(3):571–575, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267285.

Ramberg:1979:PDU

[RTDM79] John S. Ramberg, Pandu R. Tadikamalla, Edward J. Dudewicz, and Edward F. Mykytka. A probability distribution and its uses in fitting data. Technometrics, 21(2):201–214, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268517.

Rubin:1976:CRW

[Rub76] Donald B. Rubin. Comparing regressions when some predictor values are missing. Technometrics, 18(2):201–205, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267523.

Rustagi:1973:BRB

[Rus73] J. S. Rustagi. Book review: Statistical Functions by Buddy L. Myers; Norbert L. Enrick. Technometrics, 15(2):423, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267006.

Schafer:1979:EWQ

[SA79] R. E. Schafer and J. E. Angus. Estimation of Weibull quantiles with minimum error in the distribution function. Technometrics, 21(3):367–370, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267761.

Sackrowitz:1972:AML

[Sac72] Harold Sackrowitz. Alternative multi-level continuous sampling plans. Technometrics, 14(3):645–652, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267292.

Sacker:1975:BRB

[Sac75] Robert J. Sacker. Book review: Ordinary Differential Equations by V. I. Arnold. Technometrics, 17(3):388–389, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268088.

Sachs:1976:IMN

[Sac76] William H. Sachs. Implicit multifunctional nonlinear regression analysis. Technometrics, 18(2):161–173, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267519.

Salsburg:1975:BRB

[Sal75a] David S. Salsburg. Book review: The Generalized Jackknife Statistic by Henry L. Gray; William R. Schucany. Technometrics, 17(3):383– 384, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268080.

Salsburg:1975:EDV

[Sal75b] David S. Salsburg. Estimating differences in variance when comparing two methods of assay. Technometrics, 17(3):381–382, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268078.

Samuels:1973:BRB

[Sam73] S. M. Samuels. Book review: An Introduction to Probability Theory and Its Applications by William Feller. Technometrics, 15(2):420– 421, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267002.

Samuels:1978:EST

[Sam78] S. M. Samuels. The exact solution to the two-stage group-testing problem. Technometrics, 20(4):497–500, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267652.

Sanathanan:1972:MEM

[San72] Lalitha Sanathanan. Models and estimation methods in visual scanning experiments. Technometrics, 14(4):813–829, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267131.

Sanathanan:1973:CSM

[San73] Lalitha Sanathanan. A comparison of some models in visual scanning experiments. Technometrics, 15(1):67–78, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266825.

Saperstein:1973:OSW

[Sap73] B. Saperstein. On the occurrence of n successes within N Bernoulli trials. Technometrics, 15(4):809–818, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267391.

Sarkar:1971:ELC

[Sar71] Tapas K. Sarkar. An exact lower confidence bound for the reliability of a series system where each component has an exponential time to failure distribution. Technometrics, 13(3):535–546, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267166.

Saunders:1973:SAS

[Sau73] Sam C. Saunders. The statistical analysis of stellar X-ray sources. Technometrics, 15(2):341–351, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266993.

Saw:1970:LE

[Saw70] John G. Saw. Letter to the Editor. Technometrics, 12(4):937, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267344.

Saxena:1972:BRB

[Sax72] K. M. Lal Saxena. Book review: Introduction to Statistical Theory by P. G. Hoel; S. C. Port; C. J. Stone. Technometrics, 14(4):980–981, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267152.

Shenton:1972:FRM

[SB72a] L. R. Shenton and K. O. Bowman. Further remarks on maximum likelihood estimators for the gamma distribution. Technometrics, 14(3):725–733, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267300.

Swindel:1972:REI

[SB72b] Benee F. Swindel and David R. Bower. Rounding errors in the independent variables in a general linear model. Technometrics, 14(1):215–218, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266932.

Schucany:1979:LE

[SBV79] William R. Schucany, John E. Boyer, Jr., and H. D. Vinod. Letters to the Editor. Technometrics, 21(1):138, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268599.

Srivastava:1971:BOF

[SC71] J. N. Srivastava and D. V. Chopra. Balanced optimal 2^{m} fractional factorial designs of resolution V, m ≤ 6. Technometrics, 13(2):257– 269, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266788.

Scheffe:1970:NSV

[Sch70] Henry Scheffé. A note on separation of variables. Technometrics, 12(2):388–393, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267243.

Scheaffer:1971:THM

[Sch71a] R. L. Scheaffer. A test for the homogeneity of a mixture. Technometrics, 13(2):325–331, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266794.

Scheaffer:1971:OAR

[Sch71b] Richard L. Scheaffer. Optimum age replacement policies with an increasing cost factor. Technometrics, 13(1):139–144, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267081.

Scheaffer:1972:SBS

[Sch72] R. L. Scheaffer. Size-biased sampling. Technometrics, 14(3):635–644, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267291.

Schafer:1974:CBM

[Sch74a] Dick Schafer. Confidence bands for the minimum fatigue life. Technometrics, 16(1):113–123, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267499.

Schafer:1974:BRB

[Sch74b] Ray E. Schafer. Book review: A First Course in Probability and Statistics by Henrick J. Malik; Kenneth Mullen. Technometrics, 16(3):481, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267686.

Schlossmacher:1975:BRB

[Sch75] Edward Schlossmacher. Book review: Modeling and Simulation in Chemical Engineering by Rodger G. E. Franks. Technometrics, 17(1):139, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268016.

Scheaffer:1976:CCM

[Sch76] Richard L. Scheaffer. On the computation of certain minimum variance unbiased estimators. Technometrics, 18(4):497–499, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268666.

Sclove:1972:VV

[Scl72] Stanley L. Sclove. (Y vs. X) or (log Y vs. X)? Technometrics, 14(2):391–403, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267430.

Smith:1976:DME

[SCSB76] M. R. Smith, S. Cohn-Sfetcu, and H. A. Buckmaster. Decomposition of multicomponent exponential decays by spectral analytic techniques. Technometrics, 18(4):467–482, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268663.

StJohn:1975:ORD

[SD75] R. C. St. John and N. R. Draper. D-optimality for regression designs: A review. Technometrics, 17(1):15–23, February 1975. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267995.

Suich:1977:REA

[SD77] Ronald Suich and George C. Derringer. Is the regression equation adequate?: One criterion. Technometrics, 19(2):213–216, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268633.

Steffens:1973:SSR

[SdV73] F. E. Steffens and Rosalie de Villiers. Sample sizes required for two-sided comparisons of two treatments with a control. Technometrics, 15(4):915–921, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267400.

Seeger:1970:MEV

[See70] Paul Seeger. A method of estimating variance components in unbalanced designs. Technometrics, 12(2):207–218, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267229.

Seegrist:1973:LSA

[See73] Donald W. Seegrist. Least squares analysis of experimental design models by augmenting the data with side conditions. Technometrics, 15(3):643–645, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266872.

Sen:1970:BRB

[Sen70] A. R. Sen. Book review: Sampling Theory by Des Raj. Technometrics, 12(2):423–425, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267250.

Sen:1974:BRB

[Sen74] P. K. Sen. Book review: Monotone Transformations and Limit Laws by A. A. Balkema. Technometrics, 16(4):635, November 1974. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267623.

Sen:1975:BRB

[Sen75] P. K. Sen. Book review: Characterization Problems in Mathematical Statistics by A. M. Kagan; Yu. V. Linnik; C. R. Rao; B. Ramachandran. Technometrics, 17(3):385, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268082.

Serfling:1973:BRB

[Ser73] R. J. Serfling. Book review: Brownian Motion and Diffusion by David Freedman. Technometrics, 15(3):653–654, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266879.

Sarhan:1971:ETE

[SG71] A. E. Sarhan and B. G. Greenberg. The effect of the two extremes in estimating the best linear unbiased estimates (BLUE) of the parameters of the exponential distribution. Technometrics, 13(1):113–125, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267079.

Schmee:1979:SMR

[SH79] Josef Schmee and Gerald J. Hahn. A simple method for regression analysis with censored data. Technometrics, 21(4):417–432, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268280. See

discussion [Hoo79].

Shah:1971:BRB

[Sha71] B. V. Shah. Book review: Integer and Nonlinear Programming by J. Abadie. Technometrics, 13(4):921–922, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266976.

Shaked:1978:ALT

[Sha78] Moshe Shaked. Accelerated life testing for a class of linear hazard rate type distributions. Technometrics, 20(4):457–466, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267646.

Shahani:1979:WWT

[Sha79] A. K. Shahani. Wald–Wolfowitz type sampling plans for continuous production. Technometrics, 21(1):21–31, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268577.

Shumway:1971:DSS

[Shu71] Robert H. Shumway. On detecting a signal in N stationarily correlated noise series. Technometrics, 13(3):499–519, August 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267164.

Shukla:1972:PC

[Shu72] G. K. Shukla. On the problem of calibration. Technometrics, 14(3):547–553, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267283.

Suich:1970:TST

[SI70] Ronald Suich and Boris Iglewicz. A truncated sequential t-test. Technometrics, 12(4):789–798, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267325.

Siddiqui:1975:BRB

[Sid75] M. M. Siddiqui. Book review: A Guide to Probability Theory and Application by Cyrus Derman; Leon J. Gleser; Ingram Olkin. Technometrics, 17(1):140, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268018.

Sielken:1975:BRB

[Sie75] Robert L. Sielken, Jr. Book review: Practical Linear Programming with Computer Applications by Michael O. Locks. Technometrics, 17(4):514, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268445.

Simons:1973:BRB

[Sim73] Gordon D. Simons. Book review: Great Expectations: Theory of Optimal Stopping by Y. S. Chow; Herbert Robbins; David Siegmund. Technometrics, 15(3):652–653, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266877.

Simon:1976:BRB

[Sim76] Gary Simon. Book review: Basic Concepts in Information Theory and Statistics: Axiomatic Foundations and Applications by A. M. Mathai; P. N. Rathie. Technometrics, 18(4):502, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268669.

Simon:1979:BRB

[Sim79] Gary A. Simon. Book review: The Analysis of Cross-Tabulated Data by Graham J. G. Upton. Technometrics, 21(4):581–582, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268299.

Singpurwalla:1971:IAL

[Sin71] N. D. Singpurwalla. Inference from accelerated life tests when observations are obtained from censored samples. Technometrics, 13(1):161–170, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267083.

Singpurwalla:1972:EVL

[Sin72] N. D. Singpurwalla. Extreme values from a lognormal law with applications to air pollution problems. Technometrics, 14(3):703–711, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267298.

Singpurwalla:1973:IAL

[Sin73] Nozer D. Singpurwalla. Inference from accelerated life tests using Arrhenius type re-parameterizations. Technometrics, 15(2):289–299, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266989.

Schmitz:1979:OCC

[SLK79] Karl H. Schmitz, Leonard R. Lamberson, and Kailash C. Kapur. Operating characteristic curves for minimum life hypothesis testing for the exponential distribution. Technometrics, 21(4):539–547, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268294.

Sylvestre:1974:CRU

[SLM74] E. A. Sylvestre, W. H. Lawton, and M. S. Maggio. Curve resolution using a postulated chemical reaction. Technometrics, 16(3):353–368, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267665.

Snee:1974:EVD

[SM74] Ronald D. Snee and Donald W. Marquardt. Extreme vertices designs for linear mixture models. Technometrics, 16(3):399–408, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267670.

Snee:1976:SCD

[SM76] Ronald D. Snee and Donald W. Marquardt. Screening concepts and designs for experiments with mixtures. Technometrics, 18(1):19–29, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267912.

Sprevak:1979:LE

[SM79] D. Sprevak and D. Meegan. Letter to the Editor. Technometrics, 21(3):395, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267780.

Smith:1970:BRB

[Smi70a] Harry Smith, Jr. Book review: Discrete Distributions by N. L. Johnson; Samuel Kotz. Technometrics, 12(3):713, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267222.

Smith:1970:NE

[Smi70b] Harry Smith, Jr. New Editor. Technometrics, 12(4):iii, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267318.

Smith:1976:LE

[Smi76] Patricia L. Smith. Letter to the Editor: “On the distribution of the studentized bivariate range” (Technometrics 17 (1975), no 4, 501– 505) by J. E. Gentle, R. L. Kodell and Smith. Technometrics, 18(3): 365, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268754.

Smith:1977:BRB

[Smi77] Donald R. Smith. Book review: Statistical Theory of Reliability and Life Testing-Probability Models by Richard E. Barlow; Frank Proschan. Technometrics, 19(2):220, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268641.

Saxena:1977:REM

[SN77] S. K. Saxena and A. K. Nigam. Restricted exploration of mixtures by symmetric-simplex design. Technometrics, 19(1):47–52, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268253.

Snell:1970:BRB

[Sne70] James L. Snell. Book review: Convergence of Probability Measures by P. Billingsley. Technometrics, 12(1):171–172, February 1970. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267364.

Snee:1971:NUR

[Sne71] Ronald D. Snee. A note on the use of residuals for examining the assumptions of covariance analysis. Technometrics, 13(2):430–437, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266805.

Snee:1972:BRB

[Sne72a] Ronald D. Snee. Book review: Fitting Equations to Data by C. Daniel; F. S. Wood. Technometrics, 14(2):501–502, May 1972. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267439.

Snee:1972:ARC

[Sne72b] Ronald D. Snee. On the analysis of response curve data. Technometrics, 14(1):47–62, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266917.

Snee:1973:TAM

[Sne73] Ronald D. Snee. Techniques for the analysis of mixture data. Technometrics, 15(3):517–528, August 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266857.

Snee:1974:BRB

[Sne74] Ronald D. Snee. Book review: Biometrika Tables for Statisticians, Volume 2 by E. S. Pearson; H. O. Hartley. Technometrics, 16(3):480– 481, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267685.

Snee:1975:DUG

[Sne75a] Ronald D. Snee. Discussion of: “The Use of Gradients to Aid in the Interpretation of Mixture Response Surfaces”. Technometrics, 17(4):425–430, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268428. See [CO75].

Snee:1975:EDQ

[Sne75b] Ronald D. Snee. Experimental designs for quadratic models in constrained mixture spaces. Technometrics, 17(2):149–159, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268345. See erra-

tum [Ano76e].

Snee:1976:BRB

[Sne76] Ronald D. Snee. Book review: Analysis of Variance in Complex Experimental Designs by Harold R. Lindman; W. H. Freeman. Technometrics, 18(3):361–362, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268747.

Snee:1977:VRM

[Sne77] Ronald D. Snee. Validation of regression models: Methods and examples. Technometrics, 19(4):415–428, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267881.

Simpson:1975:BAM

[SOE75] Joanne Simpson, Anthony Olsen, and Jane C. Eden. A Bayesian analysis of a multiplicative treatment effect in weather modification. Technometrics, 17(2):161–166, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268346.

Somerville:1970:TCP

[Som70] Paul N. Somerville. A technique for the computation of percentage points of a statistic. Technometrics, 12(2):373–382, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267241.

Sorum:1971:ECP

[Sor71] M. J. Sorum. Estimating the conditional probability of misclassification. Technometrics, 13(2):333–343, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266795.

Sorum:1972:EEO

[Sor72a] M. Sorum. Estimating the expected and the optimal probabilities of misclassification. Technometrics, 14(4):935–943, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267143.

Sorum:1972:TPM

[Sor72b] M. Sorum. Three probabilities of misclassification. Technometrics, 14(2):309–316, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267423.

Sorum:1973:EEP

[Sor73] M. Sorum. Estimating the expected probability of misclassification for a rule based on the linear discriminant function: Univariate normal case. Technometrics, 15(2):329–339, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266992.

Shah:1972:SFA

[SP72] D. K. Shah and A. G. Phatak. A simplified form of the ASN for a curtailed sampling plan. Technometrics, 14(4):925–929, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267141.

Shah:1974:MLE

[SP74] D. K. Shah and A. G. Phatak. The maximum likelihood estimate of the fraction defective under curtailed multiple sampling plans. Technometrics, 16(2):311–315, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267953.

Shah:1977:MLE

[SP77] D. K. Shah and A. G. Phatak. The maximum likelihood estimation under curtailed three class attributes plans. Technometrics, 19(2):159–166, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268625.

Specht:1971:SEP

[Spe71] Donald F. Specht. Series estimation of a probability density function. Technometrics, 13(2):409–424, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266802.

Sprott:1972:BRB

[Spr72] D. A. Sprott. Book review: Finite Mathematics with Applications by A. W. Goodman; J. S. Ratti. Technometrics, 14(2):503–504, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267442.

Sprevak:1976:SPE

[Spr76] D. Sprevak. Statistical properties of estimates of linear models. Technometrics, 18(3):283–289, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268737.

Sprott:1978:BRB

[Spr78] D. A. Sprott. Book review: Urn Models and Their Application: An Approach to Modern Discrete Probability Theory by Norman L. Johnson; Samuel Kotz. Technometrics, 20(4):501, November 1978. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267653. See corrigenda [Ano79f].

Sprott:1979:BRB

[Spr79] D. A. Sprott. Book review: Random Allocations by V. F. Kolchin; B. A. Sevast’yanov; V. P. Chistyakov. Technometrics, 21(4):584, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268303.

Schafer:1975:SCR

[SR75] Ray E. Schafer and Herbert C. Rutemiller. Some characteristics of a ranking procedure for population parameters based on chi-square statistics. Technometrics, 17(3):327–331, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268069. See note [Bis78].

Scott:1971:NSP

[SS71a] Alastair J. Scott and Michael J. Symons. A note on shortest prediction intervals for log-linear regression. Technometrics, 13(4):889–894, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266966.

Smith:1971:IMP

[SS71b] F. Beckley Smith, Jr. and David F. Shanno. An improved Marquardt procedure for nonlinear regressions. Technometrics, 13(1):63– 74, February 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267075.

Seshadri:1974:ETZ

[SS74] V. Seshadri and J. J. Shuster. Exact tests for zero drift based on first passage times in Brownian motion. Technometrics, 16(1):133–134, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267502.

Sen:1975:SOS

[SS75] Ashish Sen and Muni S. Srivastava. Some one-sided tests for change in level. Technometrics, 17(1):61–64, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268001.

Schafer:1976:PCT

[SS76] Ray E. Schafer and T. S. Sheffield. On procedures for comparing two Weibull populations. Technometrics, 18(2):231–235, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267528.

Swallow:1978:MVQ

[SS78] William H. Swallow and S. R. Searle. Minimum variance quadratic unbiased estimation (MIVQUE) of variance components. Technometrics, 20(3):265–272, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268135.

Schafer:1972:SPR

[ST72] R. E. Schafer and Roy Takenaga. A sequential probability ratio test for availability. Technometrics, 14(1):123–135, February 1972. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266924.

Schmidt:1973:DPC

[ST73] J. W. Schmidt and R. E. Taylor. A dual purpose cost based quality control system. Technometrics, 15(1):151–166, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266832.

Stapleton:1970:BRB

[Sta70] James Stapleton. Book review: Introduction to Probability Theory and Statistical Inference by Harold J. Larson. Technometrics, 12(1):175–176, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267369.

Starks:1979:IST

[Sta79] Thomas H. Starks. An improved sign test for experiments in which neutral responses are possible. Technometrics, 21(4):525–530, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268292.

Stewart:1970:NWA

[Ste70] Kirkland B. Stewart. A new weighted average. Technometrics, 12(2):247–258, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267232.

Stefansky:1972:ROF

[Ste72a] Wilhelmine Stefansky. Rejecting outliers in factorial designs. Technometrics, 14(2):469–479, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267436.

Stewart:1972:SWA

[Ste72b] Kirkland B. Stewart. Situations in which the arithmetic and geometric moving averages are inefficient estimators. Technometrics, 14(3):619–633, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267290.

Steinberg:1975:BRB

[Ste75] Joseph Steinberg. Book review: Statistical Theory of Sample Survey Design and Analysis by H. S. Konijn. Technometrics, 17(3):383, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268079.

Stephens:1978:TEO

[Ste78] M. A. Stephens. On the W test for exponentiality with origin known. Technometrics, 20(1):33–35, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268158.

Stone:1976:BRB

[Sto76] Ricardo R. Stone. Book review: Quantitative Analysis Methods for Substantive Analysts by H. F. DeFrancesco. Technometrics, 18(2): 240, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267532.

Strauss:1975:BRB

[Str75] Monty J. Strauss. Book review: A Short Course in Differential Equations by Earl D. Rainville; Phillip E. Bedient. Technometrics, 17(2): 278, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268371.

Stoline:1979:TSM

[SU79] Michael R. Stoline and Hans K. Ury. Tables of the Studentized maximum modulus distribution and an application to multiple comparisons among means. Technometrics, 21(1):87–93, February 1979. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268584.

Shapiro:1972:AVT

[SW72] S. S. Shapiro and M. B. Wilk. An analysis of variance test for the exponential distribution (complete samples). Technometrics, 14(2):355– 370, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267427.

Srinivasan:1975:CBW

[SW75] R. Srinivasan and R. M. Wharton. Confidence bands for the Weibull distribution. Technometrics, 17(3):375–380, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268077.

Stralkowski:1970:CIE

[SWD70] C. M. Stralkowski, S. M. Wu, and R. E. DeVor. Charts for the interpretation and estimation of the second order autoregressive model. Technometrics, 12(3):669–685, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267211.

Stralkowski:1974:CIE

[SWD74] C. Michael Stralkowski, S. M. Wu, and R. E. DeVor. Charts for the interpretation and estimation of the second order moving average and mixed first order autoregressive-moving average models. Technometrics, 16(2):275–285, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267950.

Symons:1971:BRB

[Sym71] M. J. Symons. Book review: An Introduction to Probability, Decision, and Inference by Irving H. LaValle. Technometrics, 13(2):450, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266808.

Syski:1975:BRB

[Sys75] R. Syski. Book review: Dynamic Programming and Markov Potential Theory by A. Hordijk. Technometrics, 17(4):514, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268444.

Syski:1976:BRB

[Sys76] R. Syski. Book review: Markov Chains by D. Revuz. Technometrics, 18(4):501, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268668.

Syski:1978:BRB

[Sys78] R. Syski. Book review: Innovation Processes by Y. A. Rozanov. Technometrics, 20(4):503–504, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267658.

Tiao:1971:ENN

[TA71] George C. Tiao and M. M. Ali. Effect of non-normality on inferences about variance components. Technometrics, 13(3):635–650, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267174.

Thompson:1975:CDE

[TA75] William O. Thompson and Richard L. Anderson. A comparison of designs and estimators for the two-stage nested random model. Technometrics, 17(1):37–44, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267997.

Takagi:1972:DUS

[Tak72] Kinji Takagi. On designing unknown-sigma sampling plans based on a wide class of non-normal distributions. Technometrics, 14(3):669– 678, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267295.

Taneja:1970:BRB

[Tan70] V. S. Taneja. Book review: Game Theory by Guillermo Owen. Technometrics, 12(1):176, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267370.

Tarone:1976:SCE

[Tar76] Robert E. Tarone. Simultaneous confidence ellipsoids in the general linear model. Technometrics, 18(1):85–87, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267920.

Tietjen:1975:DMC

[TB75] Gary L. Tietjen and Richard J. Beckman. On duplicate measurements in the chemical laboratory. Technometrics, 17(1):53–56, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267999.

Thoman:1970:MLE

[TBA70] Darrel R. Thoman, Lee J. Bain, and Charles E. Antle. Maximum likelihood estimation, exact confidence intervals for reliability, and tolerance limits in the Weibull distribution. Technometrics, 12(2):363– 371, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267240.

Trout:1972:TPP

[TC72] J. Richard Trout and Bryant Chow. Table of the percentage points of the trivariate t-distribution with an application to uniform confidence bands. Technometrics, 14(4):855–879, November 1972. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267135.

Trout:1973:UCB

[TC73] J. Richard Trout and Bryant Chow. Uniform confidence bands for a quadratic model. Technometrics, 15(3):611–624, August 1973. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266867.

Thompson:1975:BCL

[TC75a] William E. Thompson and Eugene Y. Chang. Bayes confidence limits for reliability of redundant systems. Technometrics, 17(1):89–93, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268005.

Tilquin:1975:BRP

[TC75b] C. Tilquin and R. Cléroux. Block replacement policies with general cost structures. Technometrics, 17(3):291–298, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268063.

Tenenbein:1972:DSS

[Ten72] Aaron Tenenbein. A double sampling scheme for estimating from misclassified multinomial data with applications to sampling inspection. Technometrics, 14(1):187–202, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266930.

Thompson:1973:SCS

[Tho73] William O. Thompson. Secondary criteria in the selection of minimum bias designs in two variables. Technometrics, 15(2):319–328, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266991.

Thompson:1975:NT

[Tho75] R. Thompson. A note on the W transformation. Technometrics, 17(4):511–512, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268441.

Tietjen:1979:EST

[TJ79] Gary L. Tietjen and Mark E. Johnson. Exact statistical tolerance limits for sample variances. Technometrics, 21(1):107–110, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268586.

Tarter:1972:NTC

[TK72] Michael E. Tarter and Charles J. Kowalski. A new test for and class of transformations to normality. Technometrics, 14(3):735–744, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267301.

Tietjen:1972:SGT

[TM72] Gary L. Tietjen and Roger H. Moore. Some Grubbs-type statistics for the detection of several outliers. Technometrics, 14(3):583–597, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267287.

See [Ano79g].

Tietjen:1973:TSO

[TMB73] G. L. Tietjen, R. H. Moore, and R. J. Beckman. Testing for a single outlier in simple linear regression. Technometrics, 15(4):717–721, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267383.

See corrigendum [Ano74f].

Tong:1972:BRB

[Ton72] Yung Liang Tong. Book review: Introduction to the Probability Theory by P. G. Hoel; S. C. Port; C. J. Stone. Technometrics, 14(4):980, November 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267151.

Tong:1974:NEE

[Ton74] H. Tong. A note on the estimation of PrY < X in the exponential case. Technometrics, 16(4):625, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267617. See errata [Ton75].

Tong:1975:ENE

[Ton75] H. Tong. Errata: “A Note on the Estimation of PrY < X in the Exponential Case” [(Technometrics 16 (1974), 625)]. Technometrics, 17(3):395, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268094. See [Ton74].

Trout:1979:RII

[TS79] J. Richard Trout and William H. Swallow. Regular and inverse interval estimation of individual observations using uniform confidence bands. Technometrics, 21(4):567–574, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268297.

Tsiakals:1978:BRB

[Tsi78] Joseph J. Tsiakals. Book review: Sampling Inspection and Quality Control by G. Barrie Wetherill. Technometrics, 20(2):212, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268721.

Tukey:1970:SGS

[Tuk70] John W. Tukey. Some graphic and semigraphic displays (abstract). Technometrics, 12(1):205, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267376.

Tiahrt:1970:MCR

[TW70] Kenneth J. Tiahrt and D. L. Weeks. A method of constrained randomization for 2^{n} factorials. Technometrics, 12(3):471–486, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267197.

Tee:1972:ASD

[TW72a] L. H. Tee and S. M. Wu. An application of stochastic and dynamic models for the control of a papermaking process. Technometrics, 14(2):481–496, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267437.

Thomas:1972:LOS

[TW72b] David R. Thomas and Wanda M. Wilson. Linear order statistic estimation for the two-parameter Weibull and extreme-value distributions from Type II progressively censored samples. Technometrics, 14(3):679–691, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267296.

Urquhart:1973:BRB

[Urq73] N. Scott Urquhart. Book review: Generalized Inverse Matrices (With Applications to Statistics) by R. M. Pringle; A. A. Rayner. Technometrics, 15(2):423–424, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267007.

Ury:1972:DFC

[Ury72] Hans K. Ury. On distribution-free confidence bounds for PrY ¡ X. Technometrics, 14(3):577–581, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267286.

Ury:1976:CFP

[Ury76] Hans K. Ury. A comparison of four procedures for multiple comparisons among means (pairwise contrasts) for arbitrary sample sizes. Technometrics, 18(1):89–97, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267921.

Varde:1970:ERT

[Var70] S. D. Varde. Estimation of reliability of a two exponential component series system. Technometrics, 12(4):867–875, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267331.

vanderVaart:1975:BRB

[vdV75] H. R. van der Vaart. Book review: Mathematics for the Biological Sciences by Stanley I. Grossman; James E. Turner. Technometrics, 17(3):387–388, August 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268087.

Venetis:1970:IYW

[Ven70] C. Venetis. Intra-year water storage. Technometrics, 12(4):741–749, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267321.

Vinod:1976:SEM

[Vin76] H. D. Vinod. Simulation and extension of a minimum mean squared error estimator in comparison with Stein’s. Technometrics, 18 (4):491–496, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268665.

Vinod:1976:LE

[VK76] H. D. Vinod and R. W. Kennard. Letters to the Editor. Technometrics, 18(4):504–505, November 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268673.

Vance:1979:CMR

[VM79] Lonnie C. Vance and Gary C. McDonald. A class of multiple run sampling plans. Technometrics, 21(2):141–146, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268511.

vanNess:1979:EDD

[vN79] John van Ness. On the effects of dimension in discriminant analysis for unequal covariance populations. Technometrics, 21(1):119–127, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268588.

vanNess:1976:EDD

[vNS76] John W. van Ness and Cary Simpson. On the effects of dimension in discriminant analysis. Technometrics, 18(2):175–187, May 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267520.

Wang:1971:IMT

[Wan71] Peter C. C. Wang. An inequality on the mean time of faultless operation of a parallel system I. Technometrics, 13(4):904–909, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266970.

Watts:1970:QMP

[Wat70a] D. G. Watts. Query 2: [missing Plackett and Burman design with n = 92 runs]. Technometrics, 12(2):418–419, May 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267249.

Watts:1970:PTS

[Wat70b] Donald G. Watts. A program for training statistical consultants. Technometrics, 12(4):737–740, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267320.

Watson:1973:BRB

[Wat73] G. S. Watson. Book review: Statistics of Directional Data by K. V. Mardia. Technometrics, 15(4):935–936, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267404.

Waterman:1974:RLS

[Wat74] M. S. Waterman. A restricted least squares problem. Technometrics, 16(1):135–136, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267503.

Wilson:1971:MSP

[WB71] Edwin B. Wilson and A. R. Burgess. Multiple sampling plans viewed as finite Markov chains. Technometrics, 13(2):371–383, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266798.

Watts:1974:UHT

[WB74] Donald G. Watts and David W. Bacon. Using an hyperbola as a transition model to fit two-regime straight-line data. Technometrics, 16(3):369–373, August 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1267666. See errata [Ano75f].

Weisberg:1975:AAV

[WB75] Sanford Weisberg and Christopher Bingham. An approximate analysis of variance test for non-normality suitable for machine calculation. Technometrics, 17(1):133–134, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268012.

Weiler:1970:VWM

[WC70] H. Weiler and D. Culpin. On the variance of weighted means. Technometrics, 12(4):757–773, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267323.

Wichern:1978:CRE

[WC78] Dean W. Wichern and Gilbert A. Churchill. A comparison of ridge estimators. Technometrics, 20(3):301–311, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268139.

Weinman:1973:TET

[WDFH73] David G. Weinman, Gene Dugger, W. E. Franck, and John E. Hewett. On a test for the equality of two exponential distributions. Technometrics, 15(1):177–182, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266834.

Webb:1971:SIF

[Web71] Steve R. Webb. Small incomplete factorial experiment designs for two- and three-level factors. Technometrics, 13(2):243–256, May 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266787. See corri-

genda [Ano73f].

Webster:1973:BRB

[Web73] J. T. Webster. Book review: Engineering Statistics by Albert H. Bowker; Gerald J. Lieberman. Technometrics, 15(4):939–940, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267407.

Wegman:1972:NPD

[Weg72] Edward J. Wegman. Nonparametric probability density estimation: I. A summary of available methods. Technometrics, 14(3):533–546, August 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267282.

Wegman:1974:SRN

[Weg74] Edward J. Wegman. Some results on non stationary first order autoregression. Technometrics, 16(2):321–322, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267955.

Wegman:1976:BRB

[Weg76] Edward J. Wegman. Book review: Time Series Data Analysis and Theory by David R. Brillinger. Technometrics, 18(1):121, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267927.

Weiss:1971:BRB

[Wei71] Lionel Weiss. Book review: Probability and Statistics by L. Daniel Massey. Technometrics, 13(4):921, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266975.

Westlake:1971:OSV

[Wes71] W. J. Westlake. A one-sided version of the Tukey–Duckworth test. Technometrics, 13(4):901–903, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266969.

Weinrich:1978:BSR

[WG78] Martin C. Weinrich and Alan J. Gross. The Barlow–Scheuer reliability growth model from a Bayesian viewpoint. Technometrics, 20(3):249–254, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268133.

White:1979:LRR

[WG79] J. W. White and R. F. Gunst. Latent root regression: Large sample analysis. Technometrics, 21(4):481–488, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268287. See corrigenda [Ano80b].

Webster:1974:LRR

[WGM74] J. T. Webster, R. F. Gunst, and R. L. Mason. Latent root regression analysis. Technometrics, 16(4):513–522, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267602.

Wheeler:1970:VEV

[Whe70] Donald J. Wheeler. The variance of an estimator in variables sampling. Technometrics, 12(4):751–755, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267322.

Wheeler:1974:PP

[Whe74] Robert E. Wheeler. Portable power. Technometrics, 16(2):193–201, May 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267939.

Wheeler:1975:VPP

[Whe75] Robert E. Wheeler. The validity of portable power. Technometrics, 17(2):177–179, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268349.

Whitehouse:1970:GUT

[Whi70] Gary E. Whitehouse. GERT, a useful technique for analyzing reliability problems. Technometrics, 12(1):33–48, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267349.

Wichern:1972:EEE

[Wic72] Dean W. Wichern. The effect of estimation error on a geometric moving average. Technometrics, 14(3):745–755, August 1972. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267302.

Wiggins:1976:DFS

[Wig76] Alvin D. Wiggins. Discussion of: “Fitting Systems of Linear Differential Equations Using Computer Generated Exact Derivatives”. Technometrics, 18(4):393–399, November 1976. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268654. See [JB76].

Williams:1970:BRB

[Wil70] James S. Williams. Book review: The Structure of Inference by D. A. S. Fraser. Technometrics, 12(3):709–711, August 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267219. See comment [FW72].

Wilcox:1979:BRB

[Wil79a] Rand R. Wilcox. Book review: The Statistical Implications of Pre-Test and Stein-Rule Estimators in Econometrics by George G. Judge; M. E. Bock. Technometrics, 21(3):389–390, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267769.

Wilkinson:1979:NE

[Wil79b] John W. Wilkinson. New Editor. Technometrics, 21(4):iii, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268277.

Winkler:1971:BRB

[Win71] R. L. Winkler. Book review: Statistical Analysis for Decision Making by Morris Hamburg. Technometrics, 13(3):705–706, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267185.

Wiorkowski:1970:EPV

[Wio70] John J. Wiorkowski. Estimation of the proportion of the variance explained by regression, when the number of parameters in the model may depend on the sample size. Technometrics, 12(4):915–919, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267336.

See comment [Cro71].

Wani:1971:PER

[WK71] J. K. Wani and D. G. Kabe. Point estimation of reliability of a system comprised of k elements from the same gamma model. Technometrics, 13(4):859–864, November 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266961.

Woodward:1977:MVU

[WK77] Wayne A. Woodward and Gary D. Kelley. Minimum variance unbiased estimation of P [Y < X] in the normal case. Technometrics, 19(1):95–98, February 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268261.

Wold:1974:SFD

[Wol74] Svante Wold. Spline functions in data analysis. Technometrics, 16(1):1–11, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267485.

Wolfe:1977:DFT

[Wol77a] Douglas A. Wolfe. A distribution-free test for related correlation coefficients. Technometrics, 19(4):507–509, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267893. See errata [Ano78e].

Wolfe:1977:TST

[Wol77b] Douglas A. Wolfe. Two-stage two-sample median test. Technometrics, 19(4):495–501, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267891.

Wold:1978:CVE

[Wol78] Svante Wold. Cross-validatory estimation of the number of components in factor and principal components models. Technometrics, 20(4):397–405, November 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267639.

Wood:1973:UIE

[Woo73] Fred S. Wood. The use of individual effects and residuals in fitting equations to data. Technometrics, 15(4):677–695, November 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267381.

Woolson:1974:BRBb

[Woo74a] Robert F. Woolson. Book review: Concepts of Statistical Inference by William C. Guenther. Technometrics, 16(4):635–636, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267625.

Woolson:1974:BRBa

[Woo74b] Robert F. Woolson. Book review: Introduction to Probability and Statistics by Henry L. Alder; Edward B. Roessler. Technometrics, 16(1):142, February 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267513.

Woods:1975:BRB

[Woo75] Raymond F. Woods. Book review: Quality Control by Richard C. Vaughn. Technometrics, 17(4):513, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268443.

Woodruff:1975:DEC

[WRLI75] H. B. Woodruff, G. L. Ritter, S. R. Lowry, and T. L. Isenhour. Density estimations and the characterization of binary infrared spectra. Technometrics, 17(4):455–462, November 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268432.

Wu:1973:LE

[Wu73] S. M. Wu. Letter to the Editor. Technometrics, 15(1):207–208, February 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266847.

Willan:1978:MMM

[WW78] Andrew R. Willan and Donald G. Watts. Meaningful multicollinearity measures. Technometrics, 20(4):407–412, November 1978. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267640.

Whitmore:1978:NLT

[WY78] G. A. Whitmore and M. Yalovsky. A normalizing logarithmic transformation for inverse Gaussian random variables. Technometrics, 20(2):207–208, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268715.

Yahil:1976:SCB

[YB76] Amos Yahil and Morton B. Brown. On separating clusters from background. Technometrics, 18(1):55–58, February 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267916.

Youngs:1971:SRR

[YC71] Edward A. Youngs and Elliot M. Cramer. Some results relevant to choice of sum and sum-of-product algorithms. Technometrics, 13(3):657–665, August 1971. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267176.

Yochmowitz:1978:STM

[YC78] Michael G. Yochmowitz and Richard G. Cornell. Stepwise tests for multiplicative components of interaction. Technometrics, 20(1):79– 84, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268163.

Yale:1976:WR

[YF76] Coralee Yale and Alan B. Forsythe. Winsorized regression. Technometrics, 18(3):291–300, August 1976. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1268738.

Yuen:1974:PPD

[YM74] Karen K. Yuen and V. K. Murthy. Percentage points of the distribution of the t statistic when the parent is Student’s t. Technometrics, 16(4):495–497, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267600.

Youden:1972:EV

[You72a] W. J. Youden. Enduring values. Technometrics, 14(1):1–11, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266913.

Youden:1972:RE

[You72b] W. J. Youden. Randomization and experimentation. Technometrics, 14(1):13–22, February 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266914.

Yuen:1978:QWS

[Yue78] C. K. Yuen. Quadratic windowing in the segment averaging method for power spectrum computation. Technometrics, 20(2):195–200, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268713.

Zirphile:1975:LE

[ZA75] J. Zirphile and Etudes Avancées. Letter to the Editor. Technometrics, 17(1):145, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268023.

Zacks:1970:BDS

[Zac70] S. Zacks. Bayesian design of single and double stratified sampling for estimating proportion in finite population. Technometrics, 12(1):119–130, February 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267356.

Zacks:1974:BRB

[Zac74] Shelley Zacks. Book review: Probability Distributions: An Introduction to Probability Theory with Applications by Chris P. Tsokos. Technometrics, 16(4):634–635, November 1974. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267622.

Zahn:1975:ESH

[Zah75a] Douglas A. Zahn. An empirical study of the half-normal plot. Technometrics, 17(2):201–211, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268352.

Zahn:1975:MRC

[Zah75b] Douglas A. Zahn. Modifications of and revised critical values for the half-normal plot. Technometrics, 17(2):189–200, May 1975. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268351.

Zeigler:1973:ECV

[Zei73] R. K. Zeigler. Estimators of coefficient of variation using k samples. Technometrics, 15(2):409–414, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266999.

Zelen:1972:LER

[Zel72] Marvin Zelen. [Linear Estimation of a Regression Relationship from Censored Data. Part I: Simple Methods and Their Application]: Discussion no. 1. Technometrics, 14(2):271–276, May 1972. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267419. See [NH72].

Ziegel:1978:BRB

[Zie78] Eric R. Ziegel. Book review: Introduction to Statistics and Probability by Edward J. Dudewicz. Technometrics, 20(2):213–214, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268723.

Zizes:1978:BRB

[Ziz78] P. A. Zizes. Book review: Manual of Product Assurance Films on Reliability Engineering and Management, Reliability Testing, Maintainability, and Quality Control by Dimitri Kececioglu. Technometrics, 20(1):105, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268171.

Zidek:1979:SEA

[ZNL79] James V. Zidek, Francis P. D. Navin, and Richard Lockhart. Statistics of extremes: An alternate method with application to bridge design codes. Technometrics, 21(2):185–191, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268515.

Zyskind:1970:BRB

[Zys70] George Zyskind. Book review: Introduction to Matrices with Applications in Statistics by Franklin A. Graybill. Technometrics, 12(4):929–931, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267341.